### v2

# Medical Spoken Named Entity Recognition

Abstract
Spoken Named Entity Recognition (NER) aims to extracting named entities from speech and categorizing them into types like person, location, organization, etc. In this work, we present VietMed-NER - the first spoken NER dataset in the medical domain. To our best knowledge, our real-world dataset is the largest spoken NER dataset in the world in terms of the number of entity types, featuring 18 distinct types. Secondly, we present baseline results using various state-of-the-art pre-trained models: encoder-only and sequence-to-sequence. We found that pre-trained multilingual models XLM-R outperformed all monolingual models on both reference text and ASR output. Also in general, encoders perform better than sequence-to-sequence models for the NER task. By simply translating, the transcript is applicable not just to Vietnamese but to other languages as well. All code, data and models are made publicly available here.
Medical Spoken Named Entity Recognition
Khai Le-Duc∗1,3,4,5, David Thulke1,2, Hung-Phong Tran7, Long Vo-Dang8, Khai-Nguyen Nguyen9, Truong-Son Hy5,6, Ralf Schlüter1,2 1Lehrstuhl Informatik 6 - Machine Learning and Human Language Technology Group, Computer Science Department, RWTH Aachen University, 52074 Aachen, Germany 2AppTek GmbH, 52062 Aachen, Germany 3University of Toronto, Canada 4University Health Network, Canada 5FPT Software AI Center, Vietnam 6Indiana State University, United States 7Hanoi University of Science and Technology, Vietnam 8University of Cincinnati, United States 9College of William and Mary, United States duckhai.le@mail.utoronto.ca TruongSon.Hy@indstate.edu, {thulke,schlueter}@hltpr.rwth-aachen.de
(*)(*)footnotetext: Work partially done during the bachelor thesis at Lehrstuhl Informatik 6 - Machine Learning and Human Language Technology Group, Computer Science Department, RWTH Aachen University
1 Introduction
Named Entity Recognition (NER) targets extracting named entities (NE) from text and categorizing them into types like person, location, organization, etc. Initially studied in written language, recent attention has turned to studying spoken NER Cohn et al. (2019); Shon et al. (2022), which aims to extract semantic information from speech. However, spoken NER has limited literature compared to NER on written text data Yadav et al. (2020).
Spoken NER is particularly challenging, firstly due to the impact of word segmentation on results. The Vietnamese vocabulary poses difficulties with numerous confused monosyllabic and polysyllabic words. For instance, the word "đường" alone could denote "sugar" (chemical), "street" (location), or be part of a compound word like "đường tiêu hóa" - "gastrointestinal" (anatomy). This confusion has also been reported in Chinese spoken NER by Chen et al. (2022). Secondly, data quality control and annotation consistency have been problematic, with some entities tagged in one sentence but not in others, and full NEs inconsistently tagged as multiple sub-NEs Huyen and Luong (2016); Nguyen et al. (2018, 2020); Truong et al. (2021). Thirdly, obtaining accurate medical NER from natural speech is challenging due to the lack of punctuation, speech disfluency, context, and the complexity of medical terms.
As for the medical domain, to our best knowledge, there is no dataset available for medical spoken NER. The only related work we found, Cohn et al. (2019), published a NER evaluation benchmark using an English general-domain conversational dataset, Switchboard Godfrey et al. (1992) and Fisher Cieri et al. (2004), for the task of audio de-identification specifically targeting Personal Health Identifiers.
To address this gap, we introduce VietMed-NER, a medical spoken NER dataset built on the real-world medical Automatic Speech Recognition (ASR) dataset VietMed Le-Duc (2024), featuring 18 medically-defined entity types. The dataset has the potential to serve as a foundational task in various real-world applications, such as: error correction for medical ASR output Mani et al. (2020), audio de-identification. By simply translating, the transcript is applicable not just to Vietnamese but to other languages as well, such as: privacy-preserved machine learning (e.g. for model training and public dataset creation), and virtual assistant (e.g. search engine Rüd et al. (2011), classifying content for news providers Kumaran and Allan (2004), and recommending content Koperski et al. (2017)). Our contributions are as follows:
  • We present VietMed-NER - the first publicly-available medical spoken NER dataset.
  • We present our annotation scheme, called Recursive Greedy Mapping, intending to improve annotation speed and quality in the medical domain.
  • We present baselines on several state-of-the-art pre-trained models and analyze the results.
All code, data (English-translated and Vietnamese) and models are published online [^1] https://github.com/leduckhai/MultiMed.
2 Data
2.1 Data Collection
Entity Type and Definition Train Dev Test All
Total Uni. Total Uni. Total Uni. Total Uni.
AGE 447 43 108 25 611 83 1166 151
Age of a person
GENDER 202 30 46 15 451 33 699 78
Gender of a person
OCCUPATION 543 32 133 16 562 43 1238 91
Job of a person
LOCATION 284 66 76 31 317 75 677 172
Locations and places
ORGANIZATION 19 14 2 2 58 23 79 39
Organizations
DISEASESYMTOM 2699 518 683 209 1334 357 4716 1084
Diseases and symptoms
DRUGCHEMICAL 1054 255 263 104 684 136 2001 495
Bio-chemical substances and drugs
FOODDRINK 243 77 48 26 247 43 538 146
Food and beverage
ORGAN 1827 252 444 122 1190 172 3461 546
Anatomical features, e.g. organs, cells
PERSONALCARE 353 114 82 38 95 10 530 162
Personal care, e.g. hygiene routines, skin care
DIAGNOSTICS 371 53 91 25 292 36 754 114
Diagnostic procedures, e.g. lab tests, imaging
TREATMENT 726 69 174 25 230 17 1130 111
Non-surgical treatment, e.g. rehab., injection
SURGERY 197 29 55 13 270 37 522 79
Surgical procedures, e.g. implants, neurosurgery
PREVENTIVEMED 341 53 80 25 18 6 439 84
Preventive medicine
MEDDEVICETECHNIQUE 324 84 67 30 603 144 994 258
Medical devices, instruments, and techniques
UNITCALIBRATOR 800 155 215 75 251 106 1266 336
Medical calibration, e.g. number of doses, calories
TRANSPORTATION 5 2 3 3 27 10 35 15
Means of transportation
DATETIME 674 155 159 65 657 133 1490 353
Date and time
#Entities in total 11109 2001 2729 849 7897 1464 21735 4314
#Sentences 4620 1150 3500 9270
Table 1: Entity definition and its statistics in our dataset. "Uni." means the number of unique entities, which equals to the length of gazetteer list.
We chose the VietMed dataset Le-Duc (2024), the world’s largest and most generalizable and publicly-available medical ASR dataset, for annotating NEs.
2.2 Annotation Process
Annotation of medical NEs from real-world speech is challenging because of the missing punctuation, special characters and capitalized words in ASR transcript, disfluencies and required medical knowledge. Entirely manual annotation of NEs like in VLSP dataset Huyen and Luong (2016); Nguyen et al. (2018, 2020) and PhoNER_COVID19 Truong et al. (2021) requires a large number of working hours, not to mention the difficulties in quality control and inconsistency as we found in their corpora. This inconsistency includes: i) Some entities tagged in a sentence are not also tagged in other sentences, and ii) Full NEs are inconsistently tagged as multiple sub-NEs.
To tackle these problems, we present a human-machine annotation approach, called Recursive Greedy Mapping. The method is described below:
  1. 1.
    Annotate and categorize some initial entities, then add them to a gazetteer list.
  2. 2.
    Sort entities by character length from highest to lowest, to distinguish between sub-NEs and full NEs, ensuring full NEs are mapped before sub-NEs. Time complexity = $O(k\cdot\log(k))$ where $k$ is the number of NEs.
  3. 3.
    Automatically map entities from the gazetteer list to the transcript. Time complexity = $O(m\cdot n)$, where $m$ is the number of NEs in gazeeter list, $n$ is the number of sentences. Pseudo code: {python} for NE in gazetteer_list: for sen in sentences: if NE in sen: annotate(NE, sen)
  4. 4.
    Annotators review each sentence to include correctly labeled NEs and ignore mislabeled NEs
  5. 5.
    Annotators add new NEs not in the gazetteer list during manual annotation. Steps 2 and 3 generate pre-tagged labels in the next sentences. Annotators repeat Steps 4 and 5 until the entire corpus is annotated.
In short, Recursive Greedy Mapping speeds up annotation by allowing annotators to foresee possible NEs in upcoming utterances based on previously annotated ones. Annotators can accept or reject these suggestions, saving time with correct suggestions and easily ignoring incorrect ones. Unlike training a seed model with a gazetteer, which requires initial training time and may be unreliable for low-to-mid resource languages, our method avoids these issues and eliminates the need to correct incorrect NEs.
2.3 Data Quality Control
We created initial annotation guidelines (see Appendix) and began annotating the corpus. Two developers, one with a medical background, independently annotated the corpus. Then, we held a discussion session to resolve conflicts, address complex cases, and refine the guidelines. Two other developers perform quality control using the guidelines and the annotated corpus. We consistently revisited each sentence in the entire corpus multiple times.
2.4 Data Statistics
Table 1 shows the statistics of our dataset. Our VietMed-NER contains 18 entity types across 9000 sentences, split into train-dev-test as 8-2-6 hours. To the best of our knowledge, compared to all other public spoken NER datasets, ours has the largest number of entity types.
Most NER datasets have a very small number of entities in test sets compared to train and dev set Huyen and Luong (2016); Truong et al. (2021); Chen et al. (2022). But we want to leverage the capabilities of large pre-trained models which were trained on vast amounts of unlabeled text data, resulting in good representations. Therefore, we leave the test set to be larger than train and dev sets. However, this split caused slight data imbalance.
3 Experimental Setups
We employ the 2-stage pipeline for spoken NER: A hybrid ASR model transcribes audio into text and then the transcribed text is fed into a NER model.
3.1 F1 Score
We employed seqeval [^2] https://github.com/chakki-works/seqeval framework commonly used as a default evaluation framework by HuggingFace. However, this framework only works for NER on reference text. Therefore, we also employed the F1 score calculation by Shon et al. (2023) by using the SLUE toolkit [^3] https://github.com/asappresearch/slue-toolkit. This F1 score evaluates an unordered list of named entity phrase and tag pairs predicted for each sentence. Our proposed modification of SLUE was also used and presented in the Appendix.
3.2 ASR Models
We employed 2 best baseline models fine-tuned for ASR task on VietMed published by Le-Duc (2024): monolingual w2v2-Viet and cross-lingual XLSR-53-Viet. w2v2-Viet model was pre-trained on 1204h of Vietnamese data. XLSR-53-Viet model was pre-trained on 1204h of Vietnamese with XLSR-53 Conneau et al. (2021) as initialization. Both of them have the same number of parameters (118M) and were fine-tuned on the same train set. Their WERs on test set were 29.0% and 28.8% respectively.
3.3 NER Models
Model #Params #Data
PhoBERT_base 135M 20GB
PhoBERT_large 370M
PhoBERT_base-v2 135M 140GB
ViDeBERTa_base 86M 298GB
XLM-R_base 270M 2.5TB
XLM-R_large 550M 2.5TB
mBART-50 611M 3.9TB
ViT5_base 310M  [^8] GB
BARTpho 396M 20GB
Table 2: Statistics of state-of-the-art pre-trained language models which we used for NER task.
Table 2 shows the statistics of various pre-trained monolingual and multilingual models to fine-tune on our dataset. Monolingual: PhoBERT_base, PhoBERT_large, PhoBERT_base-v2 Nguyen and Nguyen (2020); ViDeBERTa_base Tran et al. (2023), BARTpho Tran et al. (2021), ViT5 Phan et al. (2022), Multilingual: XML-R_base, XML-R_large Conneau et al. (2020). To our best knowledge, these are the best pre-trained models that achieved state-of-the-art results on various downstream tasks in the Vietnamese language, including NER.
We used HuggingFace Wolf et al. (2019) for fine-tuning pre-trained models on the NER task. All our NER experiments were done using the default hyperparameters from HuggingFace. Details of the hyperparameters are shown in the Appendix.
4 Experimental Results
Model Prec. Recall F1
BARTpho 0.64 0.73 0.68
mBART-50 0.64 0.66 0.65
PhoBERT_base 0.67 0.78 0.72
PhoBERT_base-v2 0.68 0.79 0.74
PhoBERT_large 0.69 0.77 0.73
ViDeBERTa_base 0.50 0.41 0.45
ViT5_base 0.64 0.74 0.69
XLM-R_base 0.64 0.73 0.69
XLM-R_large 0.71 0.77 0.74
Table 3: NER results (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro F1 score. Results by entity types are shown in Appendix.
NER ASR Prec Rec. F1
ViDeBERTa_base XLSR-53-Viet 0.45 0.34 0.39
w2v2-Viet 0.45 0.34 0.39
ViT5_base XLSR-53-Viet 0.52 0.46 0.48
w2v2-Viet 0.53 0.46 0.49
mBART-50 XLSR-53-Viet 0.35 0.05 0.09
w2v2-Viet 0.35 0.05 0.09
BARTpho XLSR-53-Viet 0.56 0.50 0.53
w2v2-Viet 0.55 0.50 0.52
PhoBERT_base_v2 XLSR-53-Viet 0.57 0.57 0.57
w2v2-Viet 0.58 0.56 0.57
PhoBERT_base XLSR-53-Viet 0.56 0.56 0.56
w2v2-Viet 0.56 0.56 0.56
PhoBERT_large XLSR-53-Viet 0.57 0.55 0.56
w2v2-Viet 0.58 0.55 0.56
XLM-R_base XLSR-53-Viet 0.54 0.52 0.53
w2v2-Viet 0.54 0.52 0.53
XLM-R_large XLSR-53-Viet 0.60 0.56 0.58
w2v2-Viet 0.60 0.56 0.58
Table 4: NER results (in percent) on ASR output of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro F1 score. Results by entity types are shown in Appendix.
Table 3 and 4 show results of NER using various pre-trained models. The pre-trained monolingual model PhoBERT_base-v2 outperformed other monolingual models, at 61.99% of F1 score on reference text, and 51.29% on ASR output. Even though it has fewer parameters than PhoBERT_large, it achieved equal performance probably due to the larger amount of pre-training data. All pre-trained multilingual models outperformed monolingual models. This gap can be explained by the larger amount of pre-training data (2.5TB of multilingual data for XLM-R compared to only 140GB of monolingual data for PhoBERT_base-v2). Furthermore, encoder-based models perform, on average, better than seq2seq models, which could be due to the nature of seq2seq being generative, thus less suitable for NER. From the results from Table 3 and 4, we observe that there is a performance drop in all models when evaluated on ASR transcripts. This behavior is expected due to the noisy nature of ASR data.
5 Conclusion
In this work, we present VietMed-NER - the first spoken NER dataset in the medical domain. Our dataset contains 18 entity types, including both conventional and newly defined entity types for real-world medical conversations. Furthermore, we present an incremental annotation technique based on gazeeteer list, called Recursive Greedy Mapping, to accelerate the annotation speed and remove inconsistency during annotating. Finally, we found that pre-trained multilingual models XLM-R_base and XLM-R_large outperformed all monolingual models on both reference text and ASR output, and in general, encoders perform better than seq2seqs for the NER task.
6 Limitations
Our annotation approach: Our annotation approach using the Recursive Greedy Mapping technique has some advantages over the fully manual approach. First, it allows annotators to not spend extra time tagging the entities that have been tagged in previous sentences. Second, it prevents that annotators miss entities that have been tagged in previous sentences, improving the consistency of the entire dataset. During our work, we experienced a faster annotation by using our approach compared to fully manual annotation. However, in the scope of this paper, we have not done extensive experiments to give a quantitative number of how much time has been saved and the method’s impact on annotation quality.
Appendix A Annotation Guidelines
This section describes annotation guidelines for annotators to follow in an attempt to have a unified and consistent gold-standard NER transcript.
General rules:
  • If 2 or more entities overlap, label the resulting entity as the longest, including overlapping component entities. In other words, a full NE might contain 2 or more sub-NEs. A full NE should be tagged instead of multiple sub-NEs. For example: "bác sĩ xương khớp" (orthopedic doctor) should be tagged as a whole instead of 2 distinct NEs "bác sĩ " (doctor) and "xương khớp" (orthopedic).
  • Do not assign spaces at the beginning and at the end of entities.
  • All words in the ASR transcript are lowercase, without punctuations and special characters. Treat every word as lowercase or uppercase, with or without punctuations and special characters based on the context of each utterance.
  • Each utterance should be treated as an independent utterance. The additional context given by other utterances should not influence the annotation of each utterance.
AGE:
This entity type describes the age of a person.
  • Label the word "tuổi" (age) if applicable. For example: "tuổi trưởng thành" (mature age), "hai bảy tuổi" (twenty-seven years old).
  • List a range of ages if applicable For example: "hai mươi đến ba lăm tuổi" (twenty to thirty-five years old), "dưới sáu tháng tuổi" (under six months old).
  • Include adjectives and nouns that might describe how old a person is but don’t explicitly describe gender or gender is neutral. For example: "chưa trưởng thành" (immature), "người già" (old person), "cụ" (sir, old).
GENDER:
This entity type describes the gender of a person.
  • Include typical entities that are widely understood to describe the gender of a person. For example: "nam" (male), "đàn ông" (gentleman), "phụ nữ" (woman).
  • Include the titles and pronouns that explicitly describe a gender instead of age. For example: "ông" (grandfather), "bà" (grandmother), "cô" (aunt), "chú" (uncle).
OCCUPATION:
This entity type describes the job of a person.
  • Include all jobs that might be both in medical fields and non-medical fields. For example: "khán thính giả" (audience), "kẻ bắt cóc" (kidnapper), "bệnh nhân" (patient), "người dân" (citizen), "chuyên gia" (expert).
  • Include academic titles and degrees. For example: "thạc sĩ" (master degree holder), "tiến sĩ" (doctorate), "trưởng khoa" (dean), "chủ tịch" (president).
  • Include a cluster of words that might describe the specializations of doctors. For example: "bác sĩ chuyên về rối loạn vận động" (doctor who specializes in movement disorders) instead of two distinct entities "bác sĩ" (doctor) and "rối loạn vận động" (movement disorders), "bác sĩ về parkinson" (parkinson’s doctor) instead of two distinct entities "bác sĩ" (doctor) and "parkinson", "bác sĩ chuyên khoa tim mạch" (cardiovascular specialist) instead of two distinct entities "bác sĩ" (doctor) and "chuyên khoa tim mạch" (cardiovascular).
LOCATION:
This entity type describes a location.
  • Include continents, countries, regions, cities, and geographical administrative units. For example: "châu âu" (europe), "hoa kỳ" (usa), "tây tạng" (tibet), "thành phố hồ chí minh" (ho chi minh city), "tỉnh vĩnh long" (vinh long province).
  • Label words that mean geographical administrative units if applicable. For example: "huyện" (rural district), "quận" (urban district), "đường phố" (street), "thành phố" (city).
  • Include words that might describe public and private sites. For example: "tại nhà" (at home), "đồng ruộng" (farm), "tiệm thuốc" (drugstore), "nhà máy" (factory), "cửa hiệu quần áo" (clothing store), "toilet" (toilet).
  • Include words that might describe ambient environments. For example: "tại khu phố" (in the neighborhood), "tại địa phương" (in local area), "nước ngoài" (in foreign countries), "địa bàn" (area), "ngoài trời" (outside).
  • Include words that might describe medical facilities. For example: "chuyên khoa tiêu hóa" (gastrointestinal room) , "icu" (intensive care unit), "trạm xá" (clinics), "phòng thí nghiệm" (laboratory).
  • Each level of the administrative unit is a separate entity.
  • Do not assign nationality as an entity.
  • Locations might be misrecognized as organizations. Do not label places that are not clearly identified or controversial.
DISEASESYMPTOM:
This entity type describes a symptom or disease.
  • Include the complements of the disease. For example: "biến chứng" (side-effect), "chấn thương" (damaged), "bẩm sinh" (congenital), "di chứng" (sequelae), "bị tổn thương" (damaged), "tái phát" (relapse), "dương tính" (positive), "bệnh lý mãn tính" (chronic disease), "hội chứng" (syndrome).
  • Include a cluster of words that might describe the severity of a disease. For example: "phỏng cấp độ ba" (third-degree burn), "sức đề kháng kém" (poor immune system).
  • Mental state might also describe mental diseases or their symptoms. For example: "tự ti" (self-deprecation), "tình trạng lo âu" (state of anxiety), "mệt mỏi về tinh thần" (mental fatigue).
  • Skin conditions might describe dermatosis or its symptoms. For example: "nám" (melasma), "da đổ dầu" (oily skin), "da khô" (dry skin), "sạm da" (dark skin).
  • Genital conditions might describe genital diseases or their symptoms. For example: "có kinh" (menstruation), "có thai" (pregnant), "dậy thì sớm" (early puberty).
  • Healthy conditions might help doctors diagnose. For example: "kinh nguyệt đều" (regular menstruation).
  • Words describing physical status might also speak of symptoms or diseases. For example: "buồn ngủ" (sleepy), "rụng tóc" (hair loss), "còi cọc" (stunted).
  • Words describing children’s activities might also speak of pediatric symptoms or diseases. For example: "quấy khóc" (fussy), "không thể giao tiếp" (unable to speak), "chậm đi" (delay walking).
  • Medical techniques or devices might make symptoms and diseases happen. For example: "phẫu thuật thẩm mỹ" (cosmetic surgery).
DRUGCHEMICAL:
This entity type describes a bio-chemical substance or medicament.
  • Extraction of human or animal bodies to serve medical treatment might be referred to as a biochemical substance. For example: "vắcxin" (vaccine), "huyết thanh" (blood serum)
  • Cosmetics might be referred to as chemical substances. For example: "kem chống nắng" (sunscreen), "kem dưỡng ẩm" (moisturizer).
  • Food or drink serving medical treatment purposes or as a part of a chemical compound might be referred to as chemical substances. For example: "nấm đông trùng hạ thảo" (cordyceps), "nhân sâm" (ginseng), "nhung hươu" (deer antler).
  • Substances extracted from cells or bodies not serving medical purposes might be referred to as bio-chemical substances. For example: "dịch tiêu hóa" (digestive fluids), "chất nội sinh" (endogenous substances), "mồ hôi" (sweat), "bã nhờn" (sebum).
  • Air might be referred to as chemical substances. For example: "dưỡng khí" (breath air), "oxy" (oxygen).
FOODDRINK:
This entity type describes food and beverage.
  • Include food and drink that might serve nutrient purposes. For example: "sữa" (milk), "ngũ cốc" (cereal).
  • Include food and drink that might be harmful to health. For example: "thuốc lá" (cigarette), "rượu bia" (alcohol).
  • Include words that generally describe food and beverage. For example: "thực phẩm" (aliment), "thức ăn" (food).
ORGAN:
This entity type describes an anatomical feature, e.g. human organs, biological cells, etc. Annotators should follow general rules.
PERSONALCARE:
This entity type describes a personal care procedure, e.g. hygiene routines, skin care, daily habits, etc.
  • Activities serving the improvement of physical, aesthetic and mental health instead of medical treatment purposes might be referred to as personal care procedures. For example: "ăn kiêng" (diet), "chăm sóc da" (skin care), "chăm sóc răng" (dental care).
  • Methods serving self-improvement of speech ability in speech-language pathology might be referred to as personal care. For example: "tương tác ngôn ngữ" (language interaction), "huấn luyện ngôn ngữ" (language training).
DIAGNOSTICS:
This entity type describes a diagnostic procedure, e.g. lab tests, imaging, blood measurement, etc.
  • General words describing diagnostic procedures without explicitly mentioning surgery might be referred to as diagnostic produces. For example: "chẩn đoán" (diagnosis), "xét nghiệm" (test).
  • Imaging methods might be referred to as diagnostic procedures instead of medical devices or techniques. For example: "mri" (magnetic resonance imaging), "ct" (computed tomography).
TREATMENT:
This entity type describes a non-surgical treatment method for diseases, e.g. physical rehabilitation, injection, psychology, etc.
  • Words describing methods of using bio-chemical substances as non-surgical treatment methods might be referred to as treatment methods. For example: "liệu pháp hoocmon" (hormone therapy), "điều trị hoocmon" (hormone treatment), "điều trị tế bào gốc" (stem cell treatment).
  • Words describing methods of using invasive techniques as treatment methods might be referred to as treatment methods. For example: "hóa trị" (chemotherapy), "xạ trị" (radiotherapy).
  • Words describing methods to improve skin conditions for treatment purposes rather than aesthetics might be referred to as treatment methods. For example: "phục hồi da" (skin recovery), "ức chế sự xuất sắc tố" (inhibit pigmentation).
SURGERY:
This entity type describes a surgical treatment method for diseases, e.g. implants, neurosurgery, invasion, etc.
  • Include pre-surgery procedures that might be integral parts of surgeries. For example: "gây mê" (anesthesia), "gây tê" (anesthetize).
  • Include intervention procedures that might be integral parts of dental care. For example: "nhổ răng" (tooth extraction), "implant" (dental implant).
  • Include intervention procedures that might be integral parts of pregnancy or genitals. For example: "sinh mổ" (caesarean), "cấy tránh thai" (contraceptive implant).
  • Include intervention procedures on arteries even though they might be not integral parts of surgery. For example: "truyền máu" (blood transfusion), "truyền nước biển" (seawater infusion).
  • Include neurosurgical procedures that work with brain waves even though they might be minimally invasive. For example: "kích thích não sâu" (dbs or deep brain stimulation).
MEDDEVICETECHNIQUE:
This entity type describes a medical device, instrument, bio-material and technique.
  • Medical devices and techniques might be confusing. Annotators are strongly recommended to fully annotate CHEM., FnB, ANAT., PC, DX, TX, and SX before engaging TECH.
UNITCALIBRATOR:
This entity type describes a medical calibration, e.g. number of doses, calories, length, volume, etc.
  • Include a cluster of words that both describe the quantity and its unit. Measurements including length, distance, area, weight, heat, velocity, temperature, etc., should be explicitly tagged. For example: "năm milimet" (five millimeters) instead of "năm" (five) or "milimet" (millimeter).
  • Complements to the actual quantity describing its approximation should be included. For example: "khoảng mười lăm phần trăm" (about fifteen percent) instead of "mười lăm phần trăm" (fifteen percent).
  • Include words that generally describe the quantity. For example: "gần đủ" (close enough), "cao" (high), "rất là lớn" (very large).
  • Include words that describe trends of quantity. For example: "giảm được ít nhất" (reduce at least), "mức độ gia tăng" (level increases).
TRANSPORTATION:
This entity type describes means of transportation or vehicles.
DATETIME:
This entity type describes the date and time.
  • Include words describing day, week, month, certain named period, season, year, etc.
  • Include words describing a time frame. For example: "bây giờ" (now), "về lâu về dài" (in the long run).
  • Include words describing the approximate time. For example: "nhanh nhất có thể" (as fast as possible), "càng sớm" (as soon as possible), "từ từ" (gradually).
  • Include words describing repetitions. For example: "định kỳ" (periodically).
  • Include a cluster of words that both describe time and its complements. For example: "từ tháng ba trở đi" (from march onwards) instead of 3 distinct entities "từ" (from), "tháng ba" (march), and "trở đi" (onwards).
Appendix B Discussion about NEER
B.1 Motivation of NEER
ASR system performance is typically assessed using WER, which represents the ratio of word insertion, substitution, and deletion errors in a transcript to the total number of spoken words. However, various spoken language understanding tasks, such as spoken NER, depend on identifying keywords in transcripts. Moreover, it’s essential to recognize that in medical ASR, medical terms carry much higher significance in doctor-patient conversations and should not be treated equally to regular words. KER is often used to evaluate on keywords but is not a directly comparable metric with WER.
The purpose to introduce NEER aims to bridge the gap between WER and KER. However, it is not intended to replace WER or KER as a standard metric for evaluating domain-specific ASR performance. Instead, NEER serves as a complementary metric, facilitating a more in-depth analysis of ASR errors in specific domains, such as the medical field.
B.2 Definition of WER
WER is calculated based on the Levenshtein distance Levenshtein et al. (1966), which represents the smallest count of individual edits (insertions, deletions, or substitutions) needed to transform one word into another.
$WER=\frac{S+D+I}{N}=\frac{S+D+I}{S+D+C}$ (1)
where S is the number of substitutions, D is the number of deletions, I is the number of insertions, C is the number of correct words, and N is the number of words in the reference data $(N=S+D+C)$.
In other words, S is the number of replaced words. D is the number of missed words that are not in ASR hypothesis but are in reference data. I is the number of added words that are in ASR hypothesis but are not in reference data. The alignment between ASR hypothesis and reference data goes from left to right.
B.3 Definition of KER
Like WER, KER is computed using the Levenshtein distance. Each ASR hypothesis is aligned with its corresponding reference data and KER is calculated based on the keyword set.
$WER=\frac{F+M}{N}$ (2)
where N is the number of keywords in the reference data, F is the number of falsely recognized keywords, M is the number of missed keywords.
The ASR hypothesis often exceeds the length of all keywords in the reference data, and the insertion errors caused by non-keywords may lead to a skewed result in KER. Therefore, no insertion errors are considered while calculating KER.
B.4 Definition of NEER
In KER metric, N is the number of keywords in the reference data. KER could be characterized as the average number of errors per keyword. Nevertheless, the length of keywords may range from 1 to L (where L equals 5 in certain instances such as NER), making the average number of errors per keyword obscure.
In NEER metric, we want to evaluate on keyword-only like KER metric, while also analyzing errors per word like WER metric. Therefore, we change N into the length of keywords (entities), which characterizes the average number of errors per word of keywords.
B.5 Open questions on NEER
We still leave some questions open for future work. First, the analysis of how each type of word error (substitutions, insertions, deletions) influences NER on top of ASR has not been conducted yet. Second, the empirical relationship between WER, KER, NEER, and F1 score - meaning how KER, NEER, and F1 score are affected by a varying range of WERs — has not been analyzed either.
Appendix C Modified evaluation of SLUE toolkit
Following pre-processing, we calculate the evaluation metrics for the ASR-NER SLUE task. This involves computing precision, recall, and F1-score, which provide insights into the model performance at both an individual label level (per entity) and across all labels (overall).
We introduce a "dummy" token strategy to replace the actual named entities. This approach upholds the focus on the classification of entities rather than the extraction of verbatim phrases, which is suitable for cases where ASR errors might skew the recognition of entities in spoken transcripts.
Let’s take an example:
  • Reference text: "I have a tooth pain"
  • BIO encoding of reference text: [0, 0, 0, B-DISEASESYMTOM, I-DISEASESYMTOM]
  • ASR output: "Has teeth pain"
  • BIO encoding of ASR output: [0, B-DISEASESYMTOM, I-DISEASESYMTOM]
In the SLUE toolkit, the format (NE type, NE) is used to compare reference text and ASR output, e.g. (DISEASESYMTOM, "tooth pain") and (DISEASESYMTOM, "teeth pain"). This format gives an F1 score of 0.0 although entity type is correctly recognized. In our "dummy" token strategy, we modify the format as (NE type, "dummy"), turning reference text and ASR output to (DISEASESYMTOM, "dummy") and (DISEASESYMTOM, "dummy") respectively. The modified format gives a correct F1 score of 1.0.
We compute two types of overall metrics: micro and macro averages. The micro average metrics aggregate the contributions of all classes to compute the average metric, while the macro average computes per-entity type metrics and averages them, without considering the frequency of each entity type. The micro average is therefore influenced by the class distribution and will be dominated by the performance on more frequent entity types. In contrast, macro averages treat all entity types equally, providing a measure of the system’s performance across different types of named entities, regardless of their frequency in the dataset.
Appendix D Sequence-to-sequence Training for NER Task
Following the approach proposed by Phan et al. (2021) and later adopted by ViT5 Phan et al. (2022), we formulated the token classification as a sequence-to-sequence task by training the models to generate tags of labels before and after an entity token. In cases where the models fail to follow the mentioned format for an entity token, we use an “exception” tag, which will be later ignored during metric calculation, as the label.
Appendix E Details of Training Hyperparameters
We used HuggingFace Wolf et al. (2019) for fine-tuning pre-trained models for the NER task. We only employed word-level settings for training NER models. All our NER experiments were done using the default hyperparameters by HuggingFace.
The default hyperparameters are as follows: Learning rate of 2e-5, linear learning rate scheduler, training batch size of 64, 50 training epochs, weight decay of 0.01, AdamW optimizer Loshchilov and Hutter (2019), Beta1 of 0.9, Beta2 of 0.999, and epsilon of 1e-8.
Appendix F NER Results by Entity Types
Table LABEL:appx_NER_results_byLabel shows the results of NER on reference text by entity types using various pre-trained language models. Table LABEL:appx_spk_NER_results_byLabel_ shows the results of NER on ASR output by entity types using various pre-trained language models and ASR models.
Figure 1 shows the scatter plot of NER results on reference text by entity types using various pre-trained language models, created by Table LABEL:appx_NER_results_byLabel. Figure 2 shows the scatter plot of NER results on ASR output by entity types using various pre-trained language models, created by Table LABEL:appx_spk_NER_results_byLabel_.
Table 5: NER results by entity types (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
AGE BARTpho Mod. SLUE 0,70 0,79 0,74
BARTpho seqeval 0,69 0,61 0,65
BARTpho SLUE 0,69 0,78 0,73
mBART-50 Mod. SLUE 0,64 0,81 0,71
mBART-50 seqeval 0,66 0,53 0,59
mBART-50 SLUE 0,62 0,79 0,70
PhoBERT_base Mod. SLUE 0,75 0,81 0,78
PhoBERT_base seqeval 0,76 0,62 0,68
PhoBERT_base SLUE 0,74 0,80 0,77
PhoBERT_base-v2 Mod. SLUE 0,78 0,78 0,78
PhoBERT_base-v2 seqeval 0,76 0,66 0,71
PhoBERT_base-v2 SLUE 0,77 0,77 0,77
PhoBERT_large Mod. SLUE 0,77 0,81 0,79
PhoBERT_large seqeval 0,77 0,66 0,71
PhoBERT_large SLUE 0,76 0,80 0,78
ViDeBERTa_base Mod. SLUE 0,59 0,05 0,09
ViDeBERTa_base seqeval 0,03 0,29 0,05
ViDeBERTa_base SLUE 0,50 0,04 0,07
ViT5_base Mod. SLUE 0,71 0,79 0,75
ViT5_base seqeval 0,73 0,63 0,68
ViT5_base SLUE 0,69 0,77 0,73
XLM-R_base Mod. SLUE 0,69 0,79 0,73
XLM-R_base seqeval 0,71 0,56 0,63
XLM-R_base SLUE 0,68 0,77 0,72
XLM-R_large Mod. SLUE 0,78 0,77 0,78
XLM-R_large seqeval 0,75 0,67 0,71
XLM-R_large SLUE 0,78 0,77 0,77
DATETIME BARTpho Mod. SLUE 0,76 0,76 0,76
BARTpho seqeval 0,62 0,66 0,64
BARTpho SLUE 0,74 0,75 0,75
mBART-50 Mod. SLUE 0,83 0,74 0,78
mBART-50 seqeval 0,67 0,75 0,71
mBART-50 SLUE 0,82 0,73 0,77
PhoBERT_base Mod. SLUE 0,80 0,83 0,82
PhoBERT_base seqeval 0,77 0,70 0,74
PhoBERT_base SLUE 0,80 0,82 0,81
PhoBERT_base-v2 Mod. SLUE 0,81 0,84 0,83
PhoBERT_base-v2 seqeval 0,78 0,73 0,75
PhoBERT_base-v2 SLUE 0,81 0,83 0,82
PhoBERT_large Mod. SLUE 0,83 0,82 0,83
PhoBERT_large seqeval 0,76 0,75 0,76
PhoBERT_large SLUE 0,83 0,81 0,82
ViDeBERTa_base Mod. SLUE 0,62 0,68 0,65
ViDeBERTa_base seqeval 0,58 0,43 0,50
ViDeBERTa_base SLUE 0,60 0,65 0,62
ViT5_base Mod. SLUE 0,74 0,77 0,75
ViT5_base seqeval 0,69 0,67 0,68
ViT5_base SLUE 0,72 0,75 0,74
XLM-R_base Mod. SLUE 0,81 0,78 0,80
XLM-R_base seqeval 0,72 0,70 0,71
XLM-R_base SLUE 0,80 0,77 0,78
XLM-R_large Mod. SLUE 0,85 0,82 0,83
XLM-R_large seqeval 0,76 0,77 0,77
XLM-R_large SLUE 0,84 0,81 0,82
DIAGNOSTICS BARTpho Mod. SLUE 0,66 0,82 0,73
BARTpho seqeval 0,72 0,65 0,68
BARTpho SLUE 0,65 0,82 0,73
mBART-50 Mod. SLUE 0,58 0,81 0,68
mBART-50 seqeval 0,71 0,59 0,65
mBART-50 SLUE 0,57 0,80 0,66
PhoBERT_base Mod. SLUE 0,66 0,81 0,73
PhoBERT_base seqeval 0,75 0,64 0,69
PhoBERT_base SLUE 0,66 0,80 0,72
PhoBERT_base-v2 Mod. SLUE 0,66 0,82 0,73
PhoBERT_base-v2 seqeval 0,77 0,65 0,70
PhoBERT_base-v2 SLUE 0,66 0,82 0,73
PhoBERT_large Mod. SLUE 0,70 0,83 0,76
PhoBERT_large seqeval 0,78 0,69 0,73
PhoBERT_large SLUE 0,69 0,83 0,75
ViDeBERTa_base Mod. SLUE 0,53 0,60 0,56
ViDeBERTa_base seqeval 0,55 0,39 0,45
ViDeBERTa_base SLUE 0,51 0,58 0,54
ViT5_base Mod. SLUE 0,60 0,80 0,69
ViT5_base seqeval 0,71 0,62 0,67
ViT5_base SLUE 0,59 0,78 0,67
XLM-R_base Mod. SLUE 0,61 0,82 0,70
XLM-R_base seqeval 0,75 0,58 0,65
XLM-R_base SLUE 0,60 0,82 0,69
XLM-R_large Mod. SLUE 0,69 0,78 0,73
XLM-R_large seqeval 0,78 0,69 0,73
XLM-R_large SLUE 0,69 0,78 0,73
DISEASESYMTOM BARTpho Mod. SLUE 0,75 0,81 0,78
BARTpho seqeval 0,63 0,60 0,61
BARTpho SLUE 0,73 0,78 0,75
mBART-50 Mod. SLUE 0,71 0,73 0,72
mBART-50 seqeval 0,57 0,56 0,57
mBART-50 SLUE 0,68 0,69 0,69
PhoBERT_base Mod. SLUE 0,79 0,83 0,81
PhoBERT_base seqeval 0,70 0,60 0,65
PhoBERT_base SLUE 0,77 0,80 0,78
PhoBERT_base-v2 Mod. SLUE 0,79 0,85 0,82
PhoBERT_base-v2 seqeval 0,73 0,61 0,66
PhoBERT_base-v2 SLUE 0,77 0,82 0,79
PhoBERT_large Mod. SLUE 0,82 0,81 0,82
PhoBERT_large seqeval 0,71 0,64 0,67
PhoBERT_large SLUE 0,79 0,79 0,79
ViDeBERTa_base Mod. SLUE 0,68 0,52 0,59
ViDeBERTa_base seqeval 0,35 0,38 0,37
ViDeBERTa_base SLUE 0,62 0,47 0,54
ViT5_base Mod. SLUE 0,78 0,84 0,81
ViT5_base seqeval 0,73 0,67 0,70
ViT5_base SLUE 0,77 0,82 0,79
XLM-R_base Mod. SLUE 0,75 0,83 0,79
XLM-R_base seqeval 0,69 0,56 0,61
XLM-R_base SLUE 0,72 0,79 0,75
XLM-R_large Mod. SLUE 0,83 0,81 0,82
XLM-R_large seqeval 0,70 0,66 0,68
XLM-R_large SLUE 0,81 0,79 0,80
DRUGCHEMICAL BARTpho Mod. SLUE 0,73 0,82 0,77
BARTpho seqeval 0,75 0,70 0,73
BARTpho SLUE 0,73 0,81 0,77
mBART-50 Mod. SLUE 0,79 0,69 0,74
mBART-50 seqeval 0,71 0,76 0,73
mBART-50 SLUE 0,79 0,69 0,74
PhoBERT_base Mod. SLUE 0,80 0,93 0,86
PhoBERT_base seqeval 0,91 0,77 0,83
PhoBERT_base SLUE 0,79 0,93 0,86
PhoBERT_base-v2 Mod. SLUE 0,83 0,93 0,88
PhoBERT_base-v2 seqeval 0,91 0,80 0,85
PhoBERT_base-v2 SLUE 0,83 0,93 0,88
PhoBERT_large Mod. SLUE 0,74 0,93 0,82
PhoBERT_large seqeval 0,93 0,72 0,81
PhoBERT_large SLUE 0,74 0,93 0,82
ViDeBERTa_base Mod. SLUE 0,65 0,41 0,51
ViDeBERTa_base seqeval 0,31 0,41 0,35
ViDeBERTa_base SLUE 0,60 0,38 0,47
ViT5_base Mod. SLUE 0,75 0,87 0,80
ViT5_base seqeval 0,83 0,74 0,78
ViT5_base SLUE 0,75 0,86 0,80
XLM-R_base Mod. SLUE 0,81 0,74 0,77
XLM-R_base seqeval 0,76 0,76 0,76
XLM-R_base SLUE 0,80 0,74 0,77
XLM-R_large Mod. SLUE 0,85 0,92 0,88
XLM-R_large seqeval 0,91 0,79 0,85
XLM-R_large SLUE 0,85 0,92 0,88
FOODDRINK BARTpho Mod. SLUE 0,67 0,83 0,74
BARTpho seqeval 0,69 0,59 0,64
BARTpho SLUE 0,67 0,83 0,74
mBART-50 Mod. SLUE 0,56 0,67 0,61
mBART-50 seqeval 0,57 0,52 0,54
mBART-50 SLUE 0,56 0,67 0,61
PhoBERT_base Mod. SLUE 0,78 0,84 0,81
PhoBERT_base seqeval 0,74 0,67 0,70
PhoBERT_base SLUE 0,78 0,84 0,81
PhoBERT_base-v2 Mod. SLUE 0,82 0,89 0,85
PhoBERT_base-v2 seqeval 0,83 0,76 0,79
PhoBERT_base-v2 SLUE 0,82 0,89 0,85
PhoBERT_large Mod. SLUE 0,80 0,91 0,85
PhoBERT_large seqeval 0,85 0,75 0,80
PhoBERT_large SLUE 0,80 0,91 0,85
ViDeBERTa_base Mod. SLUE 0,22 0,38 0,28
ViDeBERTa_base seqeval 0,24 0,12 0,16
ViDeBERTa_base SLUE 0,19 0,32 0,23
ViT5_base Mod. SLUE 0,70 0,86 0,77
ViT5_base seqeval 0,75 0,66 0,70
ViT5_base SLUE 0,70 0,86 0,77
XLM-R_base Mod. SLUE 0,64 0,88 0,74
XLM-R_base seqeval 0,82 0,58 0,68
XLM-R_base SLUE 0,62 0,85 0,72
XLM-R_large Mod. SLUE 0,88 0,81 0,84
XLM-R_large seqeval 0,75 0,83 0,79
XLM-R_large SLUE 0,88 0,81 0,84
GENDER BARTpho Mod. SLUE 0,83 0,90 0,86
BARTpho seqeval 0,85 0,78 0,81
BARTpho SLUE 0,82 0,90 0,86
mBART-50 Mod. SLUE 0,83 0,87 0,85
mBART-50 seqeval 0,77 0,75 0,76
mBART-50 SLUE 0,82 0,86 0,84
PhoBERT_base Mod. SLUE 0,83 0,90 0,86
PhoBERT_base seqeval 0,76 0,74 0,75
PhoBERT_base SLUE 0,83 0,90 0,86
PhoBERT_base-v2 Mod. SLUE 0,84 0,89 0,86
PhoBERT_base-v2 seqeval 0,83 0,77 0,80
PhoBERT_base-v2 SLUE 0,84 0,89 0,86
PhoBERT_large Mod. SLUE 0,83 0,90 0,87
PhoBERT_large seqeval 0,87 0,79 0,83
PhoBERT_large SLUE 0,83 0,90 0,86
ViDeBERTa_base Mod. SLUE 0,00 0,00 0,00
ViDeBERTa_base seqeval 0,00 0,00 0,00
ViDeBERTa_base SLUE 0,00 0,00 0,00
ViT5_base Mod. SLUE 0,81 0,82 0,82
ViT5_base seqeval 0,67 0,71 0,69
ViT5_base SLUE 0,81 0,82 0,81
XLM-R_base Mod. SLUE 0,79 0,72 0,75
XLM-R_base seqeval 0,56 0,49 0,52
XLM-R_base SLUE 0,78 0,71 0,75
XLM-R_large Mod. SLUE 0,91 0,90 0,90
XLM-R_large seqeval 0,82 0,79 0,80
XLM-R_large SLUE 0,91 0,89 0,90
LOCATION BARTpho Mod. SLUE 0,77 0,88 0,82
BARTpho seqeval 0,75 0,68 0,71
BARTpho SLUE 0,76 0,87 0,81
mBART-50 Mod. SLUE 0,74 0,83 0,78
mBART-50 seqeval 0,73 0,68 0,71
mBART-50 SLUE 0,74 0,82 0,78
PhoBERT_base Mod. SLUE 0,75 0,92 0,82
PhoBERT_base seqeval 0,81 0,63 0,71
PhoBERT_base SLUE 0,74 0,91 0,81
PhoBERT_base-v2 Mod. SLUE 0,78 0,95 0,86
PhoBERT_base-v2 seqeval 0,86 0,69 0,77
PhoBERT_base-v2 SLUE 0,78 0,94 0,86
PhoBERT_large Mod. SLUE 0,80 0,91 0,85
PhoBERT_large seqeval 0,82 0,69 0,75
PhoBERT_large SLUE 0,80 0,90 0,84
ViDeBERTa_base Mod. SLUE 0,78 0,33 0,47
ViDeBERTa_base seqeval 0,32 0,60 0,42
ViDeBERTa_base SLUE 0,77 0,33 0,46
ViT5_base Mod. SLUE 0,74 0,92 0,82
ViT5_base seqeval 0,83 0,65 0,73
ViT5_base SLUE 0,73 0,92 0,81
XLM-R_base Mod. SLUE 0,75 0,84 0,79
XLM-R_base seqeval 0,75 0,59 0,66
XLM-R_base SLUE 0,74 0,82 0,78
XLM-R_large Mod. SLUE 0,85 0,93 0,89
XLM-R_large seqeval 0,87 0,75 0,81
XLM-R_large SLUE 0,85 0,93 0,88
MEDDEVICETECHNIQUE BARTpho Mod. SLUE 0,54 0,21 0,30
BARTpho seqeval 0,16 0,39 0,23
BARTpho SLUE 0,48 0,18 0,26
mBART-50 Mod. SLUE 0,45 0,09 0,15
mBART-50 seqeval 0,08 0,35 0,13
mBART-50 SLUE 0,36 0,07 0,12
PhoBERT_base Mod. SLUE 0,55 0,38 0,45
PhoBERT_base seqeval 0,24 0,27 0,26
PhoBERT_base SLUE 0,49 0,34 0,40
PhoBERT_base-v2 Mod. SLUE 0,59 0,38 0,46
PhoBERT_base-v2 seqeval 0,27 0,35 0,31
PhoBERT_base-v2 SLUE 0,53 0,34 0,41
PhoBERT_large Mod. SLUE 0,61 0,30 0,40
PhoBERT_large seqeval 0,22 0,35 0,27
PhoBERT_large SLUE 0,55 0,27 0,36
ViDeBERTa_base Mod. SLUE 0,34 0,14 0,20
ViDeBERTa_base seqeval 0,06 0,09 0,07
ViDeBERTa_base SLUE 0,24 0,10 0,14
ViT5_base Mod. SLUE 0,52 0,21 0,30
ViT5_base seqeval 0,16 0,37 0,23
ViT5_base SLUE 0,47 0,19 0,27
XLM-R_base Mod. SLUE 0,48 0,27 0,34
XLM-R_base seqeval 0,12 0,18 0,14
XLM-R_base SLUE 0,41 0,23 0,29
XLM-R_large Mod. SLUE 0,58 0,28 0,38
XLM-R_large seqeval 0,20 0,32 0,25
XLM-R_large SLUE 0,51 0,25 0,33
OCCUPATION BARTpho Mod. SLUE 0,91 0,93 0,92
BARTpho seqeval 0,87 0,87 0,87
BARTpho SLUE 0,91 0,92 0,92
mBART-50 Mod. SLUE 0,97 0,93 0,95
mBART-50 seqeval 0,91 0,95 0,93
mBART-50 SLUE 0,97 0,93 0,95
PhoBERT_base Mod. SLUE 0,95 0,96 0,95
PhoBERT_base seqeval 0,94 0,92 0,93
PhoBERT_base SLUE 0,95 0,96 0,95
PhoBERT_base-v2 Mod. SLUE 0,96 0,96 0,96
PhoBERT_base-v2 seqeval 0,95 0,93 0,94
PhoBERT_base-v2 SLUE 0,96 0,96 0,96
PhoBERT_large Mod. SLUE 0,97 0,95 0,96
PhoBERT_large seqeval 0,93 0,94 0,94
PhoBERT_large SLUE 0,97 0,95 0,96
ViDeBERTa_base Mod. SLUE 0,96 0,82 0,89
ViDeBERTa_base seqeval 0,81 0,91 0,86
ViDeBERTa_base SLUE 0,96 0,81 0,88
ViT5_base Mod. SLUE 0,95 0,93 0,94
ViT5_base seqeval 0,92 0,94 0,93
ViT5_base SLUE 0,95 0,93 0,94
XLM-R_base Mod. SLUE 0,88 0,96 0,92
XLM-R_base seqeval 0,93 0,83 0,88
XLM-R_base SLUE 0,88 0,96 0,92
XLM-R_large Mod. SLUE 0,97 0,97 0,97
XLM-R_large seqeval 0,95 0,95 0,95
XLM-R_large SLUE 0,97 0,97 0,97
ORGAN BARTpho Mod. SLUE 0,72 0,80 0,76
BARTpho seqeval 0,63 0,56 0,60
BARTpho SLUE 0,70 0,77 0,73
mBART-50 Mod. SLUE 0,67 0,74 0,70
mBART-50 seqeval 0,61 0,51 0,56
mBART-50 SLUE 0,64 0,71 0,67
PhoBERT_base Mod. SLUE 0,74 0,86 0,79
PhoBERT_base seqeval 0,69 0,55 0,61
PhoBERT_base SLUE 0,71 0,83 0,77
PhoBERT_base-v2 Mod. SLUE 0,74 0,88 0,80
PhoBERT_base-v2 seqeval 0,70 0,55 0,62
PhoBERT_base-v2 SLUE 0,72 0,86 0,78
PhoBERT_large Mod. SLUE 0,73 0,87 0,80
PhoBERT_large seqeval 0,71 0,55 0,62
PhoBERT_large SLUE 0,71 0,85 0,77
ViDeBERTa_base Mod. SLUE 0,53 0,51 0,52
ViDeBERTa_base seqeval 0,32 0,27 0,30
ViDeBERTa_base SLUE 0,45 0,44 0,44
ViT5_base Mod. SLUE 0,71 0,85 0,78
ViT5_base seqeval 0,70 0,58 0,64
ViT5_base SLUE 0,70 0,84 0,76
XLM-R_base Mod. SLUE 0,71 0,87 0,78
XLM-R_base seqeval 0,70 0,55 0,61
XLM-R_base SLUE 0,69 0,85 0,76
XLM-R_large Mod. SLUE 0,77 0,87 0,82
XLM-R_large seqeval 0,75 0,60 0,66
XLM-R_large SLUE 0,75 0,85 0,80
ORGANIZATION BARTpho Mod. SLUE 0,73 0,54 0,62
BARTpho seqeval 0,47 0,68 0,56
BARTpho SLUE 0,71 0,53 0,61
mBART-50 Mod. SLUE 0,77 0,31 0,44
mBART-50 seqeval 0,28 0,70 0,40
mBART-50 SLUE 0,77 0,31 0,44
PhoBERT_base Mod. SLUE 0,71 0,57 0,63
PhoBERT_base seqeval 0,51 0,56 0,53
PhoBERT_base SLUE 0,70 0,56 0,62
PhoBERT_base-v2 Mod. SLUE 0,76 0,71 0,73
PhoBERT_base-v2 seqeval 0,59 0,60 0,60
PhoBERT_base-v2 SLUE 0,74 0,70 0,72
PhoBERT_large Mod. SLUE 0,71 0,65 0,68
PhoBERT_large seqeval 0,59 0,50 0,54
PhoBERT_large SLUE 0,71 0,65 0,68
ViDeBERTa_base Mod. SLUE 0,00 0,00 0,00
ViDeBERTa_base seqeval 0,00 0,00 0,00
ViDeBERTa_base SLUE 0,00 0,00 0,00
ViT5_base Mod. SLUE 0,94 0,36 0,52
ViT5_base seqeval 0,34 0,91 0,50
ViT5_base SLUE 0,94 0,36 0,52
XLM-R_base Mod. SLUE 0,59 0,31 0,41
XLM-R_base seqeval 0,08 0,12 0,10
XLM-R_base SLUE 0,56 0,29 0,39
XLM-R_large Mod. SLUE 0,82 0,66 0,73
XLM-R_large seqeval 0,61 0,70 0,65
XLM-R_large SLUE 0,80 0,64 0,71
PERSONALCARE BARTpho Mod. SLUE 0,17 0,77 0,28
BARTpho seqeval 0,75 0,21 0,32
BARTpho SLUE 0,17 0,77 0,28
mBART-50 Mod. SLUE 0,23 0,80 0,36
mBART-50 seqeval 0,76 0,26 0,39
mBART-50 SLUE 0,23 0,80 0,36
PhoBERT_base Mod. SLUE 0,18 0,87 0,30
PhoBERT_base seqeval 0,85 0,20 0,33
PhoBERT_base SLUE 0,18 0,85 0,30
PhoBERT_base-v2 Mod. SLUE 0,16 0,85 0,27
PhoBERT_base-v2 seqeval 0,85 0,18 0,30
PhoBERT_base-v2 SLUE 0,16 0,84 0,27
PhoBERT_large Mod. SLUE 0,19 0,83 0,31
PhoBERT_large seqeval 0,80 0,21 0,33
PhoBERT_large SLUE 0,19 0,82 0,31
ViDeBERTa_base Mod. SLUE 0,17 0,76 0,28
ViDeBERTa_base seqeval 0,68 0,14 0,23
ViDeBERTa_base SLUE 0,17 0,75 0,28
ViT5_base Mod. SLUE 0,17 0,79 0,28
ViT5_base seqeval 0,75 0,18 0,29
ViT5_base SLUE 0,17 0,78 0,27
XLM-R_base Mod. SLUE 0,21 0,83 0,33
XLM-R_base seqeval 0,77 0,21 0,33
XLM-R_base SLUE 0,20 0,81 0,32
XLM-R_large Mod. SLUE 0,22 0,87 0,36
XLM-R_large seqeval 0,84 0,24 0,38
XLM-R_large SLUE 0,22 0,86 0,35
PREVENTIVEMED BARTpho Mod. SLUE 0,01 0,15 0,01
BARTpho seqeval 0,21 0,01 0,02
BARTpho SLUE 0,01 0,15 0,01
mBART-50 Mod. SLUE 0,02 0,45 0,04
mBART-50 seqeval 0,42 0,02 0,04
mBART-50 SLUE 0,02 0,45 0,04
PhoBERT_base Mod. SLUE 0,03 0,75 0,06
PhoBERT_base seqeval 0,56 0,02 0,05
PhoBERT_base SLUE 0,03 0,69 0,06
PhoBERT_base-v2 Mod. SLUE 0,02 0,44 0,04
PhoBERT_base-v2 seqeval 0,33 0,02 0,03
PhoBERT_base-v2 SLUE 0,02 0,39 0,03
PhoBERT_large Mod. SLUE 0,03 0,78 0,06
PhoBERT_large seqeval 0,72 0,03 0,06
PhoBERT_large SLUE 0,03 0,75 0,06
ViDeBERTa_base Mod. SLUE 0,00 0,06 0,01
ViDeBERTa_base seqeval 0,00 0,00 0,00
ViDeBERTa_base SLUE 0,00 0,00 0,00
ViT5_base Mod. SLUE 0,02 0,39 0,04
ViT5_base seqeval 0,39 0,03 0,05
ViT5_base SLUE 0,02 0,35 0,04
XLM-R_base Mod. SLUE 0,01 0,14 0,02
XLM-R_base seqeval 0,00 0,00 0,00
XLM-R_base SLUE 0,01 0,08 0,01
XLM-R_large Mod. SLUE 0,03 0,78 0,06
XLM-R_large seqeval 0,72 0,03 0,05
XLM-R_large SLUE 0,03 0,78 0,06
SURGERY BARTpho Mod. SLUE 0,57 0,57 0,57
BARTpho seqeval 0,48 0,55 0,51
BARTpho SLUE 0,57 0,56 0,57
mBART-50 Mod. SLUE 0,67 0,38 0,48
mBART-50 seqeval 0,29 0,55 0,38
mBART-50 SLUE 0,64 0,37 0,47
PhoBERT_base Mod. SLUE 0,70 0,48 0,56
PhoBERT_base seqeval 0,36 0,50 0,42
PhoBERT_base SLUE 0,68 0,46 0,55
PhoBERT_base-v2 Mod. SLUE 0,79 0,66 0,72
PhoBERT_base-v2 seqeval 0,62 0,59 0,61
PhoBERT_base-v2 SLUE 0,78 0,65 0,71
PhoBERT_large Mod. SLUE 0,85 0,46 0,59
PhoBERT_large seqeval 0,38 0,66 0,49
PhoBERT_large SLUE 0,85 0,45 0,59
ViDeBERTa_base Mod. SLUE 0,78 0,32 0,46
ViDeBERTa_base seqeval 0,24 0,45 0,32
ViDeBERTa_base SLUE 0,76 0,31 0,44
ViT5_base Mod. SLUE 0,61 0,69 0,65
ViT5_base seqeval 0,55 0,51 0,52
ViT5_base SLUE 0,60 0,69 0,64
XLM-R_base Mod. SLUE 0,66 0,67 0,66
XLM-R_base seqeval 0,59 0,46 0,52
XLM-R_base SLUE 0,66 0,66 0,66
XLM-R_large Mod. SLUE 0,80 0,46 0,59
XLM-R_large seqeval 0,39 0,65 0,49
XLM-R_large SLUE 0,80 0,46 0,58
TRANSPORTATION BARTpho Mod. SLUE 0,98 0,96 0,97
BARTpho seqeval 0,93 0,93 0,93
BARTpho SLUE 0,98 0,96 0,97
mBART-50 Mod. SLUE 1,00 0,70 0,82
mBART-50 seqeval 0,67 0,95 0,78
mBART-50 SLUE 1,00 0,70 0,82
PhoBERT_base Mod. SLUE 1,00 0,90 0,95
PhoBERT_base seqeval 0,85 0,92 0,88
PhoBERT_base SLUE 1,00 0,90 0,95
PhoBERT_base-v2 Mod. SLUE 1,00 0,69 0,82
PhoBERT_base-v2 seqeval 0,59 0,80 0,68
PhoBERT_base-v2 SLUE 1,00 0,69 0,82
PhoBERT_large Mod. SLUE 0,95 0,91 0,93
PhoBERT_large seqeval 0,89 0,92 0,91
PhoBERT_large SLUE 0,95 0,91 0,93
ViDeBERTa_base Mod. SLUE 0,00 0,00 0,00
ViDeBERTa_base seqeval 0,00 0,00 0,00
ViDeBERTa_base SLUE 0,00 0,00 0,00
ViT5_base Mod. SLUE 0,00 0,00 0,00
ViT5_base seqeval 0,00 0,00 0,00
ViT5_base SLUE 0,00 0,00 0,00
XLM-R_base Mod. SLUE 0,00 0,00 0,00
XLM-R_base seqeval 0,00 0,00 0,00
XLM-R_base SLUE 0,00 0,00 0,00
XLM-R_large Mod. SLUE 1,00 0,72 0,84
XLM-R_large seqeval 0,63 0,81 0,71
XLM-R_large SLUE 1,00 0,72 0,84
TREATMENT BARTpho Mod. SLUE 0,63 0,82 0,71
BARTpho seqeval 0,75 0,57 0,65
BARTpho SLUE 0,63 0,82 0,71
mBART-50 Mod. SLUE 0,66 0,83 0,74
mBART-50 seqeval 0,81 0,61 0,70
mBART-50 SLUE 0,66 0,83 0,74
PhoBERT_base Mod. SLUE 0,66 0,88 0,75
PhoBERT_base seqeval 0,86 0,60 0,71
PhoBERT_base SLUE 0,66 0,88 0,75
PhoBERT_base-v2 Mod. SLUE 0,62 0,88 0,73
PhoBERT_base-v2 seqeval 0,87 0,56 0,68
PhoBERT_base-v2 SLUE 0,62 0,88 0,72
PhoBERT_large Mod. SLUE 0,59 0,87 0,70
PhoBERT_large seqeval 0,86 0,53 0,65
PhoBERT_large SLUE 0,59 0,87 0,70
ViDeBERTa_base Mod. SLUE 0,62 0,86 0,72
ViDeBERTa_base seqeval 0,84 0,52 0,64
ViDeBERTa_base SLUE 0,62 0,86 0,72
ViT5_base Mod. SLUE 0,46 0,84 0,60
ViT5_base seqeval 0,81 0,41 0,55
ViT5_base SLUE 0,46 0,84 0,60
XLM-R_base Mod. SLUE 0,48 0,86 0,62
XLM-R_base seqeval 0,84 0,40 0,54
XLM-R_base SLUE 0,48 0,86 0,62
XLM-R_large Mod. SLUE 0,59 0,89 0,71
XLM-R_large seqeval 0,89 0,54 0,67
XLM-R_large SLUE 0,59 0,89 0,71
UNITCALIBRATOR BARTpho Mod. SLUE 0,42 0,65 0,51
BARTpho seqeval 0,52 0,31 0,39
BARTpho SLUE 0,41 0,63 0,50
mBART-50 Mod. SLUE 0,38 0,58 0,46
mBART-50 seqeval 0,45 0,26 0,33
mBART-50 SLUE 0,37 0,56 0,44
PhoBERT_base Mod. SLUE 0,44 0,73 0,55
PhoBERT_base seqeval 0,61 0,30 0,41
PhoBERT_base SLUE 0,43 0,71 0,53
PhoBERT_base-v2 Mod. SLUE 0,46 0,74 0,57
PhoBERT_base-v2 seqeval 0,61 0,33 0,43
PhoBERT_base-v2 SLUE 0,45 0,72 0,55
PhoBERT_large Mod. SLUE 0,48 0,73 0,58
PhoBERT_large seqeval 0,60 0,34 0,43
PhoBERT_large SLUE 0,47 0,71 0,56
ViDeBERTa_base Mod. SLUE 0,32 0,44 0,37
ViDeBERTa_base seqeval 0,20 0,10 0,14
ViDeBERTa_base SLUE 0,28 0,39 0,33
ViT5_base Mod. SLUE 0,34 0,63 0,44
ViT5_base seqeval 0,50 0,24 0,32
ViT5_base SLUE 0,33 0,62 0,43
XLM-R_base Mod. SLUE 0,44 0,72 0,54
XLM-R_base seqeval 0,54 0,27 0,36
XLM-R_base SLUE 0,42 0,69 0,52
XLM-R_large Mod. SLUE 0,50 0,75 0,60
XLM-R_large seqeval 0,65 0,37 0,47
XLM-R_large SLUE 0,49 0,74 0,59
Overall Macro BARTpho Mod. SLUE 0,64 0,72 0,66
BARTpho seqeval 0,64 0,58 0,59
BARTpho SLUE 0,63 0,71 0,65
mBART-50 Mod. SLUE 0,64 0,66 0,61
mBART-50 seqeval 0,59 0,57 0,55
mBART-50 SLUE 0,63 0,65 0,60
PhoBERT_base Mod. SLUE 0,67 0,79 0,69
PhoBERT_base seqeval 0,70 0,57 0,60
PhoBERT_base SLUE 0,66 0,78 0,68
PhoBERT_base-v2 Mod. SLUE 0,69 0,79 0,71
PhoBERT_base-v2 seqeval 0,71 0,59 0,62
PhoBERT_base-v2 SLUE 0,68 0,77 0,70
PhoBERT_large Mod. SLUE 0,69 0,79 0,70
PhoBERT_large seqeval 0,73 0,60 0,63
PhoBERT_large SLUE 0,68 0,78 0,69
ViDeBERTa_base Mod. SLUE 0,43 0,38 0,37
ViDeBERTa_base seqeval 0,31 0,28 0,27
ViDeBERTa_base SLUE 0,40 0,36 0,34
ViT5_base Mod. SLUE 0,59 0,69 0,60
ViT5_base seqeval 0,61 0,53 0,54
ViT5_base SLUE 0,58 0,68 0,59
XLM-R_base Mod. SLUE 0,57 0,67 0,59
XLM-R_base seqeval 0,57 0,44 0,48
XLM-R_base SLUE 0,56 0,65 0,58
XLM-R_large Mod. SLUE 0,72 0,78 0,71
XLM-R_large seqeval 0,72 0,62 0,63
XLM-R_large SLUE 0,71 0,77 0,70
Overral Micro BARTpho Mod. SLUE 0,66 0,74 0,70
BARTpho seqeval 0,64 0,58 0,61
BARTpho SLUE 0,64 0,73 0,68
mBART-50 Mod. SLUE 0,65 0,68 0,67
mBART-50 seqeval 0,60 0,57 0,59
mBART-50 SLUE 0,64 0,66 0,65
PhoBERT_base Mod. SLUE 0,69 0,79 0,74
PhoBERT_base seqeval 0,71 0,57 0,63
PhoBERT_base SLUE 0,67 0,78 0,72
PhoBERT_base-v2 Mod. SLUE 0,70 0,81 0,75
PhoBERT_base-v2 seqeval 0,74 0,59 0,66
PhoBERT_base-v2 SLUE 0,68 0,79 0,74
PhoBERT_large Mod. SLUE 0,70 0,79 0,74
PhoBERT_large seqeval 0,73 0,60 0,66
PhoBERT_large SLUE 0,69 0,77 0,73
ViDeBERTa_base Mod. SLUE 0,55 0,45 0,49
ViDeBERTa_base seqeval 0,33 0,34 0,34
ViDeBERTa_base SLUE 0,50 0,41 0,45
ViT5_base Mod. SLUE 0,65 0,76 0,70
ViT5_base seqeval 0,68 0,59 0,63
ViT5_base SLUE 0,64 0,74 0,69
XLM-R_base Mod. SLUE 0,66 0,75 0,70
XLM-R_base seqeval 0,66 0,53 0,59
XLM-R_base SLUE 0,64 0,73 0,69
XLM-R_large Mod. SLUE 0,72 0,78 0,75
XLM-R_large seqeval 0,72 0,63 0,67
XLM-R_large SLUE 0,71 0,77 0,74
Table 6: NER results by entity types (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score.
Entity Type NER ASR Eval toolkit Prec. Rec. F1
AGE ViDeBERTa_base XLSR-53-Viet SLUE 0,24 0,03 0,05
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,32 0,04 0,07
ViDeBERTa_base w2v2-Viet SLUE 0,24 0,03 0,05
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,33 0,04 0,07
ViT5_base XLSR-53-Viet SLUE 0,63 0,49 0,55
ViT5_base XLSR-53-Viet Mod. SLUE 0,68 0,53 0,59
ViT5_base w2v2-Viet SLUE 0,63 0,52 0,57
ViT5_base w2v2-Viet Mod. SLUE 0,68 0,56 0,61
mBART-50 XLSR-53-Viet SLUE 0,44 0,06 0,11
mBART-50 XLSR-53-Viet Mod. SLUE 0,55 0,08 0,14
mBART-50 w2v2-Viet SLUE 0,44 0,08 0,13
mBART-50 w2v2-Viet Mod. SLUE 0,58 0,10 0,17
BARTpho XLSR-53-Viet SLUE 0,62 0,57 0,59
BARTpho XLSR-53-Viet Mod. SLUE 0,68 0,62 0,65
BARTpho w2v2-Viet SLUE 0,59 0,58 0,58
BARTpho w2v2-Viet Mod. SLUE 0,64 0,63 0,63
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,72 0,58 0,64
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,75 0,61 0,67
PhoBERT_base-v2 w2v2-Viet SLUE 0,70 0,58 0,64
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,74 0,61 0,67
PhoBERT_base XLSR-53-Viet SLUE 0,69 0,60 0,64
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,74 0,64 0,68
PhoBERT_base w2v2-Viet SLUE 0,67 0,61 0,64
PhoBERT_base w2v2-Viet Mod. SLUE 0,70 0,64 0,67
PhoBERT_large XLSR-53-Viet SLUE 0,71 0,59 0,65
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,75 0,62 0,68
PhoBERT_large w2v2-Viet SLUE 0,69 0,60 0,64
PhoBERT_large w2v2-Viet Mod. SLUE 0,72 0,63 0,67
XLM-R_base XLSR-53-Viet SLUE 0,61 0,58 0,60
XLM-R_base XLSR-53-Viet Mod. SLUE 0,65 0,62 0,63
XLM-R_base w2v2-Viet SLUE 0,59 0,59 0,59
XLM-R_base w2v2-Viet Mod. SLUE 0,63 0,63 0,63
XLM-R_large XLSR-53-Viet SLUE 0,72 0,61 0,66
XLM-R_large XLSR-53-Viet Mod. SLUE 0,76 0,64 0,69
XLM-R_large w2v2-Viet SLUE 0,70 0,62 0,66
XLM-R_large w2v2-Viet Mod. SLUE 0,73 0,65 0,69
DATETIME ViDeBERTa_base XLSR-53-Viet SLUE 0,52 0,56 0,54
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,56 0,60 0,58
ViDeBERTa_base w2v2-Viet SLUE 0,53 0,57 0,55
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,57 0,60 0,59
ViT5_base XLSR-53-Viet SLUE 0,57 0,57 0,57
ViT5_base XLSR-53-Viet Mod. SLUE 0,63 0,62 0,62
ViT5_base w2v2-Viet SLUE 0,60 0,56 0,58
ViT5_base w2v2-Viet Mod. SLUE 0,63 0,59 0,61
mBART-50 XLSR-53-Viet SLUE 0,42 0,06 0,11
mBART-50 XLSR-53-Viet Mod. SLUE 0,57 0,09 0,15
mBART-50 w2v2-Viet SLUE 0,40 0,06 0,10
mBART-50 w2v2-Viet Mod. SLUE 0,61 0,09 0,15
BARTpho XLSR-53-Viet SLUE 0,64 0,62 0,63
BARTpho XLSR-53-Viet Mod. SLUE 0,68 0,66 0,67
BARTpho w2v2-Viet SLUE 0,65 0,60 0,62
BARTpho w2v2-Viet Mod. SLUE 0,69 0,64 0,66
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,67 0,69 0,68
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,70 0,72 0,71
PhoBERT_base-v2 w2v2-Viet SLUE 0,69 0,69 0,69
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,71 0,71 0,71
PhoBERT_base XLSR-53-Viet SLUE 0,66 0,68 0,67
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,69 0,72 0,70
PhoBERT_base w2v2-Viet SLUE 0,68 0,67 0,68
PhoBERT_base w2v2-Viet Mod. SLUE 0,71 0,70 0,70
PhoBERT_large XLSR-53-Viet SLUE 0,67 0,68 0,68
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,70 0,71 0,71
PhoBERT_large w2v2-Viet SLUE 0,70 0,66 0,68
PhoBERT_large w2v2-Viet Mod. SLUE 0,72 0,69 0,70
XLM-R_base XLSR-53-Viet SLUE 0,64 0,65 0,64
XLM-R_base XLSR-53-Viet Mod. SLUE 0,67 0,68 0,68
XLM-R_base w2v2-Viet SLUE 0,66 0,64 0,65
XLM-R_base w2v2-Viet Mod. SLUE 0,69 0,66 0,67
XLM-R_large XLSR-53-Viet SLUE 0,69 0,67 0,68
XLM-R_large XLSR-53-Viet Mod. SLUE 0,72 0,70 0,71
XLM-R_large w2v2-Viet SLUE 0,72 0,66 0,69
XLM-R_large w2v2-Viet Mod. SLUE 0,74 0,69 0,71
DIAGNOSTICS ViDeBERTa_base XLSR-53-Viet SLUE 0,53 0,47 0,50
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,55 0,49 0,52
ViDeBERTa_base w2v2-Viet SLUE 0,51 0,44 0,47
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,53 0,46 0,49
ViT5_base XLSR-53-Viet SLUE 0,52 0,47 0,50
ViT5_base XLSR-53-Viet Mod. SLUE 0,54 0,49 0,51
ViT5_base w2v2-Viet SLUE 0,53 0,46 0,49
ViT5_base w2v2-Viet Mod. SLUE 0,56 0,49 0,52
mBART-50 XLSR-53-Viet SLUE 0,30 0,08 0,13
mBART-50 XLSR-53-Viet Mod. SLUE 0,34 0,10 0,15
mBART-50 w2v2-Viet SLUE 0,41 0,06 0,11
mBART-50 w2v2-Viet Mod. SLUE 0,48 0,08 0,13
BARTpho XLSR-53-Viet SLUE 0,59 0,56 0,58
BARTpho XLSR-53-Viet Mod. SLUE 0,62 0,59 0,60
BARTpho w2v2-Viet SLUE 0,60 0,53 0,56
BARTpho w2v2-Viet Mod. SLUE 0,63 0,56 0,59
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,56 0,57 0,57
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,58 0,59 0,58
PhoBERT_base-v2 w2v2-Viet SLUE 0,56 0,56 0,56
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,59 0,59 0,59
PhoBERT_base XLSR-53-Viet SLUE 0,57 0,57 0,57
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,59 0,59 0,59
PhoBERT_base w2v2-Viet SLUE 0,55 0,56 0,55
PhoBERT_base w2v2-Viet Mod. SLUE 0,57 0,58 0,57
PhoBERT_large XLSR-53-Viet SLUE 0,61 0,58 0,59
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,62 0,59 0,61
PhoBERT_large w2v2-Viet SLUE 0,58 0,57 0,57
PhoBERT_large w2v2-Viet Mod. SLUE 0,61 0,59 0,60
XLM-R_base XLSR-53-Viet SLUE 0,51 0,58 0,54
XLM-R_base XLSR-53-Viet Mod. SLUE 0,52 0,59 0,55
XLM-R_base w2v2-Viet SLUE 0,50 0,57 0,53
XLM-R_base w2v2-Viet Mod. SLUE 0,52 0,59 0,55
XLM-R_large XLSR-53-Viet SLUE 0,59 0,58 0,59
XLM-R_large XLSR-53-Viet Mod. SLUE 0,61 0,59 0,60
XLM-R_large w2v2-Viet SLUE 0,58 0,57 0,57
XLM-R_large w2v2-Viet Mod. SLUE 0,60 0,59 0,60
DISEASESYMTOM ViDeBERTa_base XLSR-53-Viet SLUE 0,53 0,39 0,45
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,65 0,48 0,55
ViDeBERTa_base w2v2-Viet SLUE 0,53 0,39 0,45
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,64 0,48 0,55
ViT5_base XLSR-53-Viet SLUE 0,59 0,49 0,54
ViT5_base XLSR-53-Viet Mod. SLUE 0,69 0,58 0,63
ViT5_base w2v2-Viet SLUE 0,59 0,49 0,54
ViT5_base w2v2-Viet Mod. SLUE 0,70 0,58 0,64
mBART-50 XLSR-53-Viet SLUE 0,51 0,09 0,15
mBART-50 XLSR-53-Viet Mod. SLUE 0,64 0,11 0,19
mBART-50 w2v2-Viet SLUE 0,48 0,07 0,12
mBART-50 w2v2-Viet Mod. SLUE 0,65 0,09 0,16
BARTpho XLSR-53-Viet SLUE 0,61 0,53 0,57
BARTpho XLSR-53-Viet Mod. SLUE 0,71 0,62 0,66
BARTpho w2v2-Viet SLUE 0,58 0,53 0,55
BARTpho w2v2-Viet Mod. SLUE 0,69 0,62 0,65
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,63 0,57 0,60
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,71 0,65 0,68
PhoBERT_base-v2 w2v2-Viet SLUE 0,61 0,57 0,59
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,71 0,66 0,68
PhoBERT_base XLSR-53-Viet SLUE 0,62 0,56 0,59
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,71 0,64 0,67
PhoBERT_base w2v2-Viet SLUE 0,61 0,55 0,58
PhoBERT_base w2v2-Viet Mod. SLUE 0,71 0,64 0,67
PhoBERT_large XLSR-53-Viet SLUE 0,63 0,56 0,59
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,71 0,63 0,67
PhoBERT_large w2v2-Viet SLUE 0,62 0,54 0,58
PhoBERT_large w2v2-Viet Mod. SLUE 0,71 0,62 0,66
XLM-R_base XLSR-53-Viet SLUE 0,58 0,55 0,57
XLM-R_base XLSR-53-Viet Mod. SLUE 0,67 0,64 0,65
XLM-R_base w2v2-Viet SLUE 0,57 0,54 0,56
XLM-R_base w2v2-Viet Mod. SLUE 0,67 0,64 0,65
XLM-R_large XLSR-53-Viet SLUE 0,64 0,57 0,60
XLM-R_large XLSR-53-Viet Mod. SLUE 0,72 0,64 0,68
XLM-R_large w2v2-Viet SLUE 0,64 0,56 0,60
XLM-R_large w2v2-Viet Mod. SLUE 0,72 0,63 0,67
DRUGCHEMICAL ViDeBERTa_base XLSR-53-Viet SLUE 0,51 0,34 0,41
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,59 0,40 0,47
ViDeBERTa_base w2v2-Viet SLUE 0,49 0,34 0,40
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,56 0,39 0,46
ViT5_base XLSR-53-Viet SLUE 0,58 0,52 0,55
ViT5_base XLSR-53-Viet Mod. SLUE 0,66 0,59 0,62
ViT5_base w2v2-Viet SLUE 0,60 0,51 0,55
ViT5_base w2v2-Viet Mod. SLUE 0,68 0,58 0,62
mBART-50 XLSR-53-Viet SLUE 0,32 0,06 0,11
mBART-50 XLSR-53-Viet Mod. SLUE 0,46 0,09 0,15
mBART-50 w2v2-Viet SLUE 0,35 0,07 0,11
mBART-50 w2v2-Viet Mod. SLUE 0,54 0,10 0,17
BARTpho XLSR-53-Viet SLUE 0,60 0,55 0,57
BARTpho XLSR-53-Viet Mod. SLUE 0,67 0,62 0,64
BARTpho w2v2-Viet SLUE 0,60 0,55 0,57
BARTpho w2v2-Viet Mod. SLUE 0,66 0,61 0,64
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,67 0,66 0,67
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,73 0,72 0,72
PhoBERT_base-v2 w2v2-Viet SLUE 0,66 0,66 0,66
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,72 0,72 0,72
PhoBERT_base XLSR-53-Viet SLUE 0,66 0,66 0,66
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,72 0,72 0,72
PhoBERT_base w2v2-Viet SLUE 0,64 0,66 0,65
PhoBERT_base w2v2-Viet Mod. SLUE 0,71 0,72 0,72
PhoBERT_large XLSR-53-Viet SLUE 0,63 0,67 0,65
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,69 0,72 0,71
PhoBERT_large w2v2-Viet SLUE 0,64 0,66 0,65
PhoBERT_large w2v2-Viet Mod. SLUE 0,70 0,71 0,70
XLM-R_base XLSR-53-Viet SLUE 0,63 0,52 0,57
XLM-R_base XLSR-53-Viet Mod. SLUE 0,71 0,57 0,63
XLM-R_base w2v2-Viet SLUE 0,64 0,52 0,57
XLM-R_base w2v2-Viet Mod. SLUE 0,71 0,57 0,64
XLM-R_large XLSR-53-Viet SLUE 0,70 0,68 0,69
XLM-R_large XLSR-53-Viet Mod. SLUE 0,76 0,73 0,74
XLM-R_large w2v2-Viet SLUE 0,69 0,66 0,67
XLM-R_large w2v2-Viet Mod. SLUE 0,75 0,71 0,73
FOODDRINK ViDeBERTa_base XLSR-53-Viet SLUE 0,25 0,29 0,27
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,28 0,33 0,30
ViDeBERTa_base w2v2-Viet SLUE 0,22 0,27 0,24
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,25 0,31 0,28
ViT5_base XLSR-53-Viet SLUE 0,54 0,53 0,54
ViT5_base XLSR-53-Viet Mod. SLUE 0,58 0,56 0,57
ViT5_base w2v2-Viet SLUE 0,55 0,53 0,54
ViT5_base w2v2-Viet Mod. SLUE 0,59 0,58 0,59
mBART-50 XLSR-53-Viet SLUE 0,59 0,06 0,12
mBART-50 XLSR-53-Viet Mod. SLUE 0,61 0,07 0,12
mBART-50 w2v2-Viet SLUE 0,54 0,05 0,09
mBART-50 w2v2-Viet Mod. SLUE 0,63 0,06 0,11
BARTpho XLSR-53-Viet SLUE 0,65 0,49 0,56
BARTpho XLSR-53-Viet Mod. SLUE 0,68 0,51 0,58
BARTpho w2v2-Viet SLUE 0,68 0,52 0,59
BARTpho w2v2-Viet Mod. SLUE 0,71 0,54 0,61
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,68 0,68 0,68
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,69 0,68 0,68
PhoBERT_base-v2 w2v2-Viet SLUE 0,69 0,66 0,67
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,69 0,67 0,68
PhoBERT_base XLSR-53-Viet SLUE 0,63 0,65 0,64
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,64 0,66 0,65
PhoBERT_base w2v2-Viet SLUE 0,64 0,62 0,63
PhoBERT_base w2v2-Viet Mod. SLUE 0,66 0,64 0,65
PhoBERT_large XLSR-53-Viet SLUE 0,68 0,69 0,68
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,69 0,70 0,69
PhoBERT_large w2v2-Viet SLUE 0,67 0,67 0,67
PhoBERT_large w2v2-Viet Mod. SLUE 0,68 0,68 0,68
XLM-R_base XLSR-53-Viet SLUE 0,52 0,63 0,57
XLM-R_base XLSR-53-Viet Mod. SLUE 0,55 0,67 0,61
XLM-R_base w2v2-Viet SLUE 0,53 0,63 0,57
XLM-R_base w2v2-Viet Mod. SLUE 0,56 0,67 0,61
XLM-R_large XLSR-53-Viet SLUE 0,75 0,69 0,72
XLM-R_large XLSR-53-Viet Mod. SLUE 0,76 0,70 0,73
XLM-R_large w2v2-Viet SLUE 0,76 0,68 0,72
XLM-R_large w2v2-Viet Mod. SLUE 0,77 0,69 0,73
GENDER ViDeBERTa_base XLSR-53-Viet SLUE 0,00 0,00 0,00
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,00 0,00 0,00
ViDeBERTa_base w2v2-Viet SLUE 1,00 0,00 0,00
ViDeBERTa_base w2v2-Viet Mod. SLUE 1,00 0,00 0,00
ViT5_base XLSR-53-Viet SLUE 0,74 0,51 0,60
ViT5_base XLSR-53-Viet Mod. SLUE 0,79 0,54 0,65
ViT5_base w2v2-Viet SLUE 0,74 0,51 0,60
ViT5_base w2v2-Viet Mod. SLUE 0,76 0,53 0,62
mBART-50 XLSR-53-Viet SLUE 0,39 0,03 0,06
mBART-50 XLSR-53-Viet Mod. SLUE 0,50 0,04 0,08
mBART-50 w2v2-Viet SLUE 0,40 0,05 0,09
mBART-50 w2v2-Viet Mod. SLUE 0,53 0,07 0,12
BARTpho XLSR-53-Viet SLUE 0,78 0,59 0,67
BARTpho XLSR-53-Viet Mod. SLUE 0,82 0,61 0,70
BARTpho w2v2-Viet SLUE 0,78 0,58 0,66
BARTpho w2v2-Viet Mod. SLUE 0,80 0,59 0,68
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,78 0,60 0,68
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,82 0,63 0,71
PhoBERT_base-v2 w2v2-Viet SLUE 0,78 0,60 0,68
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,80 0,62 0,70
PhoBERT_base XLSR-53-Viet SLUE 0,78 0,61 0,68
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,81 0,63 0,71
PhoBERT_base w2v2-Viet SLUE 0,77 0,61 0,68
PhoBERT_base w2v2-Viet Mod. SLUE 0,80 0,62 0,70
PhoBERT_large XLSR-53-Viet SLUE 0,75 0,60 0,67
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,78 0,63 0,70
PhoBERT_large w2v2-Viet SLUE 0,76 0,60 0,67
PhoBERT_large w2v2-Viet Mod. SLUE 0,78 0,62 0,69
XLM-R_base XLSR-53-Viet SLUE 0,72 0,42 0,53
XLM-R_base XLSR-53-Viet Mod. SLUE 0,78 0,45 0,57
XLM-R_base w2v2-Viet SLUE 0,70 0,44 0,54
XLM-R_base w2v2-Viet Mod. SLUE 0,74 0,47 0,57
XLM-R_large XLSR-53-Viet SLUE 0,79 0,63 0,70
XLM-R_large XLSR-53-Viet Mod. SLUE 0,82 0,66 0,73
XLM-R_large w2v2-Viet SLUE 0,79 0,64 0,71
XLM-R_large w2v2-Viet Mod. SLUE 0,81 0,65 0,72
LOCATION ViDeBERTa_base XLSR-53-Viet SLUE 0,63 0,28 0,39
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,67 0,29 0,41
ViDeBERTa_base w2v2-Viet SLUE 0,59 0,26 0,36
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,60 0,26 0,37
ViT5_base XLSR-53-Viet SLUE 0,62 0,57 0,59
ViT5_base XLSR-53-Viet Mod. SLUE 0,65 0,60 0,63
ViT5_base w2v2-Viet SLUE 0,61 0,53 0,57
ViT5_base w2v2-Viet Mod. SLUE 0,64 0,56 0,60
mBART-50 XLSR-53-Viet SLUE 0,18 0,01 0,02
mBART-50 XLSR-53-Viet Mod. SLUE 0,41 0,02 0,04
mBART-50 w2v2-Viet SLUE 0,26 0,01 0,03
mBART-50 w2v2-Viet Mod. SLUE 0,33 0,02 0,04
BARTpho XLSR-53-Viet SLUE 0,65 0,66 0,65
BARTpho XLSR-53-Viet Mod. SLUE 0,68 0,69 0,69
BARTpho w2v2-Viet SLUE 0,65 0,65 0,65
BARTpho w2v2-Viet Mod. SLUE 0,69 0,68 0,68
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,66 0,71 0,69
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,69 0,74 0,71
PhoBERT_base-v2 w2v2-Viet SLUE 0,67 0,69 0,68
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,70 0,72 0,71
PhoBERT_base XLSR-53-Viet SLUE 0,62 0,68 0,65
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,66 0,72 0,69
PhoBERT_base w2v2-Viet SLUE 0,64 0,67 0,65
PhoBERT_base w2v2-Viet Mod. SLUE 0,67 0,71 0,69
PhoBERT_large XLSR-53-Viet SLUE 0,67 0,68 0,67
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,69 0,70 0,70
PhoBERT_large w2v2-Viet SLUE 0,70 0,66 0,68
PhoBERT_large w2v2-Viet Mod. SLUE 0,72 0,68 0,70
XLM-R_base XLSR-53-Viet SLUE 0,64 0,61 0,62
XLM-R_base XLSR-53-Viet Mod. SLUE 0,67 0,64 0,65
XLM-R_base w2v2-Viet SLUE 0,65 0,59 0,62
XLM-R_base w2v2-Viet Mod. SLUE 0,68 0,62 0,65
XLM-R_large XLSR-53-Viet SLUE 0,71 0,69 0,70
XLM-R_large XLSR-53-Viet Mod. SLUE 0,74 0,72 0,73
XLM-R_large w2v2-Viet SLUE 0,72 0,67 0,69
XLM-R_large w2v2-Viet Mod. SLUE 0,74 0,70 0,72
MEDDEVICETECHNIQUE ViDeBERTa_base XLSR-53-Viet SLUE 0,22 0,07 0,10
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,33 0,10 0,16
ViDeBERTa_base w2v2-Viet SLUE 0,20 0,06 0,10
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,33 0,11 0,16
ViT5_base XLSR-53-Viet SLUE 0,39 0,09 0,14
ViT5_base XLSR-53-Viet Mod. SLUE 0,48 0,11 0,18
ViT5_base w2v2-Viet SLUE 0,44 0,11 0,17
ViT5_base w2v2-Viet Mod. SLUE 0,53 0,13 0,21
mBART-50 XLSR-53-Viet SLUE 0,31 0,01 0,01
mBART-50 XLSR-53-Viet Mod. SLUE 0,38 0,01 0,01
mBART-50 w2v2-Viet SLUE 0,26 0,01 0,01
mBART-50 w2v2-Viet Mod. SLUE 0,32 0,01 0,01
BARTpho XLSR-53-Viet SLUE 0,42 0,09 0,15
BARTpho XLSR-53-Viet Mod. SLUE 0,51 0,11 0,18
BARTpho w2v2-Viet SLUE 0,38 0,09 0,14
BARTpho w2v2-Viet Mod. SLUE 0,49 0,11 0,18
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,46 0,20 0,28
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,56 0,24 0,33
PhoBERT_base-v2 w2v2-Viet SLUE 0,43 0,18 0,25
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,52 0,22 0,30
PhoBERT_base XLSR-53-Viet SLUE 0,48 0,22 0,30
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,56 0,26 0,35
PhoBERT_base w2v2-Viet SLUE 0,44 0,20 0,28
PhoBERT_base w2v2-Viet Mod. SLUE 0,52 0,24 0,33
PhoBERT_large XLSR-53-Viet SLUE 0,46 0,15 0,23
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,55 0,18 0,27
PhoBERT_large w2v2-Viet SLUE 0,45 0,15 0,22
PhoBERT_large w2v2-Viet Mod. SLUE 0,55 0,18 0,27
XLM-R_base XLSR-53-Viet SLUE 0,38 0,13 0,19
XLM-R_base XLSR-53-Viet Mod. SLUE 0,47 0,16 0,24
XLM-R_base w2v2-Viet SLUE 0,34 0,12 0,17
XLM-R_base w2v2-Viet Mod. SLUE 0,44 0,15 0,22
XLM-R_large XLSR-53-Viet SLUE 0,47 0,15 0,23
XLM-R_large XLSR-53-Viet Mod. SLUE 0,55 0,18 0,27
XLM-R_large w2v2-Viet SLUE 0,43 0,13 0,20
XLM-R_large w2v2-Viet Mod. SLUE 0,53 0,16 0,25
OCCUPATION ViDeBERTa_base XLSR-53-Viet SLUE 0,76 0,75 0,76
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,78 0,76 0,77
ViDeBERTa_base w2v2-Viet SLUE 0,77 0,74 0,76
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,79 0,75 0,77
ViT5_base XLSR-53-Viet SLUE 0,78 0,68 0,73
ViT5_base XLSR-53-Viet Mod. SLUE 0,79 0,70 0,74
ViT5_base w2v2-Viet SLUE 0,79 0,69 0,74
ViT5_base w2v2-Viet Mod. SLUE 0,80 0,70 0,75
mBART-50 XLSR-53-Viet SLUE 0,52 0,02 0,04
mBART-50 XLSR-53-Viet Mod. SLUE 0,54 0,02 0,05
mBART-50 w2v2-Viet SLUE 0,57 0,03 0,05
mBART-50 w2v2-Viet Mod. SLUE 0,60 0,03 0,06
BARTpho XLSR-53-Viet SLUE 0,79 0,78 0,79
BARTpho XLSR-53-Viet Mod. SLUE 0,80 0,79 0,80
BARTpho w2v2-Viet SLUE 0,79 0,77 0,78
BARTpho w2v2-Viet Mod. SLUE 0,81 0,79 0,80
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,79 0,84 0,81
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,80 0,85 0,82
PhoBERT_base-v2 w2v2-Viet SLUE 0,81 0,84 0,82
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,82 0,84 0,83
PhoBERT_base XLSR-53-Viet SLUE 0,79 0,84 0,81
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,79 0,85 0,82
PhoBERT_base w2v2-Viet SLUE 0,80 0,84 0,82
PhoBERT_base w2v2-Viet Mod. SLUE 0,81 0,85 0,83
PhoBERT_large XLSR-53-Viet SLUE 0,80 0,84 0,82
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,81 0,85 0,83
PhoBERT_large w2v2-Viet SLUE 0,82 0,84 0,83
PhoBERT_large w2v2-Viet Mod. SLUE 0,82 0,85 0,83
XLM-R_base XLSR-53-Viet SLUE 0,74 0,83 0,79
XLM-R_base XLSR-53-Viet Mod. SLUE 0,75 0,85 0,80
XLM-R_base w2v2-Viet SLUE 0,76 0,83 0,79
XLM-R_base w2v2-Viet Mod. SLUE 0,77 0,83 0,80
XLM-R_large XLSR-53-Viet SLUE 0,81 0,84 0,82
XLM-R_large XLSR-53-Viet Mod. SLUE 0,81 0,85 0,83
XLM-R_large w2v2-Viet SLUE 0,82 0,84 0,83
XLM-R_large w2v2-Viet Mod. SLUE 0,82 0,84 0,83
ORGAN ViDeBERTa_base XLSR-53-Viet SLUE 0,36 0,31 0,33
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,49 0,41 0,45
ViDeBERTa_base w2v2-Viet SLUE 0,37 0,31 0,34
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,50 0,42 0,46
ViT5_base XLSR-53-Viet SLUE 0,50 0,39 0,44
ViT5_base XLSR-53-Viet Mod. SLUE 0,59 0,46 0,52
ViT5_base w2v2-Viet SLUE 0,53 0,40 0,46
ViT5_base w2v2-Viet Mod. SLUE 0,64 0,48 0,55
mBART-50 XLSR-53-Viet SLUE 0,29 0,06 0,09
mBART-50 XLSR-53-Viet Mod. SLUE 0,46 0,09 0,15
mBART-50 w2v2-Viet SLUE 0,30 0,06 0,10
mBART-50 w2v2-Viet Mod. SLUE 0,42 0,08 0,14
BARTpho XLSR-53-Viet SLUE 0,52 0,42 0,46
BARTpho XLSR-53-Viet Mod. SLUE 0,62 0,50 0,55
BARTpho w2v2-Viet SLUE 0,53 0,42 0,47
BARTpho w2v2-Viet Mod. SLUE 0,63 0,50 0,56
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,53 0,48 0,50
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,62 0,56 0,59
PhoBERT_base-v2 w2v2-Viet SLUE 0,55 0,49 0,52
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,64 0,57 0,60
PhoBERT_base XLSR-53-Viet SLUE 0,53 0,47 0,50
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,63 0,56 0,59
PhoBERT_base w2v2-Viet SLUE 0,54 0,48 0,51
PhoBERT_base w2v2-Viet Mod. SLUE 0,63 0,56 0,59
PhoBERT_large XLSR-53-Viet SLUE 0,52 0,46 0,49
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,62 0,55 0,58
PhoBERT_large w2v2-Viet SLUE 0,53 0,48 0,50
PhoBERT_large w2v2-Viet Mod. SLUE 0,63 0,57 0,59
XLM-R_base XLSR-53-Viet SLUE 0,52 0,48 0,50
XLM-R_base XLSR-53-Viet Mod. SLUE 0,61 0,56 0,59
XLM-R_base w2v2-Viet SLUE 0,53 0,50 0,51
XLM-R_base w2v2-Viet Mod. SLUE 0,62 0,58 0,60
XLM-R_large XLSR-53-Viet SLUE 0,55 0,47 0,50
XLM-R_large XLSR-53-Viet Mod. SLUE 0,64 0,55 0,59
XLM-R_large w2v2-Viet SLUE 0,56 0,48 0,52
XLM-R_large w2v2-Viet Mod. SLUE 0,65 0,56 0,60
ORGANIZATION ViDeBERTa_base XLSR-53-Viet SLUE 0,00 0,00 0,00
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,00 0,00 0,00
ViDeBERTa_base w2v2-Viet SLUE 0,00 0,00 0,00
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,00 0,00 0,00
ViT5_base XLSR-53-Viet SLUE 0,88 0,30 0,45
ViT5_base XLSR-53-Viet Mod. SLUE 0,88 0,30 0,45
ViT5_base w2v2-Viet SLUE 0,81 0,27 0,41
ViT5_base w2v2-Viet Mod. SLUE 0,81 0,27 0,41
mBART-50 XLSR-53-Viet SLUE 0,00 0,00 0,00
mBART-50 XLSR-53-Viet Mod. SLUE 0,00 0,00 0,00
mBART-50 w2v2-Viet SLUE 0,00 0,00 0,00
mBART-50 w2v2-Viet Mod. SLUE 0,00 0,00 0,00
BARTpho XLSR-53-Viet SLUE 0,78 0,34 0,47
BARTpho XLSR-53-Viet Mod. SLUE 0,82 0,36 0,50
BARTpho w2v2-Viet SLUE 0,76 0,36 0,48
BARTpho w2v2-Viet Mod. SLUE 0,82 0,39 0,53
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,75 0,56 0,64
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,78 0,59 0,67
PhoBERT_base-v2 w2v2-Viet SLUE 0,76 0,55 0,64
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,79 0,57 0,66
PhoBERT_base XLSR-53-Viet SLUE 0,65 0,49 0,56
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,66 0,49 0,56
PhoBERT_base w2v2-Viet SLUE 0,62 0,45 0,52
PhoBERT_base w2v2-Viet Mod. SLUE 0,63 0,46 0,53
PhoBERT_large XLSR-53-Viet SLUE 0,73 0,50 0,59
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,73 0,50 0,60
PhoBERT_large w2v2-Viet SLUE 0,69 0,49 0,57
PhoBERT_large w2v2-Viet Mod. SLUE 0,71 0,50 0,59
XLM-R_base XLSR-53-Viet SLUE 0,62 0,26 0,37
XLM-R_base XLSR-53-Viet Mod. SLUE 0,64 0,27 0,38
XLM-R_base w2v2-Viet SLUE 0,57 0,26 0,35
XLM-R_base w2v2-Viet Mod. SLUE 0,59 0,27 0,37
XLM-R_large XLSR-53-Viet SLUE 0,66 0,56 0,61
XLM-R_large XLSR-53-Viet Mod. SLUE 0,68 0,58 0,63
XLM-R_large w2v2-Viet SLUE 0,65 0,55 0,60
XLM-R_large w2v2-Viet Mod. SLUE 0,69 0,58 0,63
PERSONALCARE ViDeBERTa_base XLSR-53-Viet SLUE 0,20 0,67 0,30
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,20 0,70 0,32
ViDeBERTa_base w2v2-Viet SLUE 0,20 0,67 0,31
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,20 0,69 0,31
ViT5_base XLSR-53-Viet SLUE 0,17 0,54 0,26
ViT5_base XLSR-53-Viet Mod. SLUE 0,17 0,57 0,27
ViT5_base w2v2-Viet SLUE 0,17 0,53 0,26
ViT5_base w2v2-Viet Mod. SLUE 0,18 0,57 0,27
mBART-50 XLSR-53-Viet SLUE 0,03 0,02 0,02
mBART-50 XLSR-53-Viet Mod. SLUE 0,06 0,05 0,06
mBART-50 w2v2-Viet SLUE 0,05 0,04 0,04
mBART-50 w2v2-Viet Mod. SLUE 0,05 0,05 0,05
BARTpho XLSR-53-Viet SLUE 0,19 0,63 0,29
BARTpho XLSR-53-Viet Mod. SLUE 0,19 0,65 0,30
BARTpho w2v2-Viet SLUE 0,18 0,63 0,28
BARTpho w2v2-Viet Mod. SLUE 0,19 0,64 0,29
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,18 0,71 0,29
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,19 0,73 0,30
PhoBERT_base-v2 w2v2-Viet SLUE 0,18 0,70 0,29
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,18 0,71 0,29
PhoBERT_base XLSR-53-Viet SLUE 0,19 0,69 0,30
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,20 0,71 0,31
PhoBERT_base w2v2-Viet SLUE 0,20 0,69 0,31
PhoBERT_base w2v2-Viet Mod. SLUE 0,20 0,70 0,32
PhoBERT_large XLSR-53-Viet SLUE 0,20 0,70 0,32
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,21 0,71 0,32
PhoBERT_large w2v2-Viet SLUE 0,20 0,68 0,31
PhoBERT_large w2v2-Viet Mod. SLUE 0,21 0,69 0,32
XLM-R_base XLSR-53-Viet SLUE 0,22 0,70 0,34
XLM-R_base XLSR-53-Viet Mod. SLUE 0,23 0,73 0,35
XLM-R_base w2v2-Viet SLUE 0,23 0,70 0,34
XLM-R_base w2v2-Viet Mod. SLUE 0,23 0,72 0,35
XLM-R_large XLSR-53-Viet SLUE 0,24 0,70 0,36
XLM-R_large XLSR-53-Viet Mod. SLUE 0,25 0,71 0,37
XLM-R_large w2v2-Viet SLUE 0,25 0,69 0,37
XLM-R_large w2v2-Viet Mod. SLUE 0,26 0,70 0,38
PREVENTIVEMED ViDeBERTa_base XLSR-53-Viet SLUE 0,00 0,00 0,00
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,00 0,06 0,01
ViDeBERTa_base w2v2-Viet SLUE 0,00 0,00 0,00
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,00 0,06 0,01
ViT5_base XLSR-53-Viet SLUE 0,01 0,11 0,01
ViT5_base XLSR-53-Viet Mod. SLUE 0,01 0,14 0,02
ViT5_base w2v2-Viet SLUE 0,01 0,19 0,03
ViT5_base w2v2-Viet Mod. SLUE 0,02 0,22 0,03
mBART-50 XLSR-53-Viet SLUE 0,00 0,00 0,00
mBART-50 XLSR-53-Viet Mod. SLUE 0,00 0,00 0,00
mBART-50 w2v2-Viet SLUE 0,00 0,00 0,00
mBART-50 w2v2-Viet Mod. SLUE 0,00 0,00 0,00
BARTpho XLSR-53-Viet SLUE 0,00 0,06 0,01
BARTpho XLSR-53-Viet Mod. SLUE 0,01 0,19 0,02
BARTpho w2v2-Viet SLUE 0,00 0,06 0,01
BARTpho w2v2-Viet Mod. SLUE 0,01 0,19 0,02
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,01 0,14 0,01
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,01 0,22 0,02
PhoBERT_base-v2 w2v2-Viet SLUE 0,01 0,19 0,02
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,01 0,28 0,03
PhoBERT_base XLSR-53-Viet SLUE 0,01 0,22 0,02
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,01 0,31 0,03
PhoBERT_base w2v2-Viet SLUE 0,01 0,28 0,03
PhoBERT_base w2v2-Viet Mod. SLUE 0,02 0,39 0,04
PhoBERT_large XLSR-53-Viet SLUE 0,01 0,22 0,02
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,02 0,31 0,03
PhoBERT_large w2v2-Viet SLUE 0,01 0,28 0,03
PhoBERT_large w2v2-Viet Mod. SLUE 0,02 0,36 0,04
XLM-R_base XLSR-53-Viet SLUE 0,00 0,00 0,00
XLM-R_base XLSR-53-Viet Mod. SLUE 0,00 0,06 0,01
XLM-R_base w2v2-Viet SLUE 0,00 0,00 0,00
XLM-R_base w2v2-Viet Mod. SLUE 0,00 0,06 0,01
XLM-R_large XLSR-53-Viet SLUE 0,01 0,11 0,01
XLM-R_large XLSR-53-Viet Mod. SLUE 0,01 0,19 0,02
XLM-R_large w2v2-Viet SLUE 0,01 0,14 0,01
XLM-R_large w2v2-Viet Mod. SLUE 0,01 0,22 0,02
SURGERY ViDeBERTa_base XLSR-53-Viet SLUE 0,56 0,24 0,33
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,61 0,26 0,36
ViDeBERTa_base w2v2-Viet SLUE 0,57 0,24 0,34
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,62 0,26 0,36
ViT5_base XLSR-53-Viet SLUE 0,46 0,29 0,36
ViT5_base XLSR-53-Viet Mod. SLUE 0,51 0,32 0,39
ViT5_base w2v2-Viet SLUE 0,51 0,28 0,36
ViT5_base w2v2-Viet Mod. SLUE 0,57 0,32 0,41
mBART-50 XLSR-53-Viet SLUE 0,22 0,05 0,08
mBART-50 XLSR-53-Viet Mod. SLUE 0,32 0,07 0,11
mBART-50 w2v2-Viet SLUE 0,22 0,06 0,09
mBART-50 w2v2-Viet Mod. SLUE 0,31 0,08 0,13
BARTpho XLSR-53-Viet SLUE 0,52 0,29 0,37
BARTpho XLSR-53-Viet Mod. SLUE 0,59 0,33 0,42
BARTpho w2v2-Viet SLUE 0,54 0,29 0,38
BARTpho w2v2-Viet Mod. SLUE 0,61 0,33 0,43
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,57 0,29 0,39
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,62 0,32 0,42
PhoBERT_base-v2 w2v2-Viet SLUE 0,60 0,32 0,42
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,65 0,35 0,46
PhoBERT_base XLSR-53-Viet SLUE 0,53 0,26 0,35
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,56 0,27 0,37
PhoBERT_base w2v2-Viet SLUE 0,56 0,27 0,37
PhoBERT_base w2v2-Viet Mod. SLUE 0,61 0,30 0,40
PhoBERT_large XLSR-53-Viet SLUE 0,57 0,28 0,38
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,61 0,31 0,41
PhoBERT_large w2v2-Viet SLUE 0,61 0,28 0,39
PhoBERT_large w2v2-Viet Mod. SLUE 0,66 0,31 0,42
XLM-R_base XLSR-53-Viet SLUE 0,50 0,32 0,39
XLM-R_base XLSR-53-Viet Mod. SLUE 0,54 0,35 0,42
XLM-R_base w2v2-Viet SLUE 0,54 0,33 0,41
XLM-R_base w2v2-Viet Mod. SLUE 0,58 0,35 0,44
XLM-R_large XLSR-53-Viet SLUE 0,58 0,28 0,38
XLM-R_large XLSR-53-Viet Mod. SLUE 0,62 0,30 0,41
XLM-R_large w2v2-Viet SLUE 0,62 0,30 0,40
XLM-R_large w2v2-Viet Mod. SLUE 0,65 0,32 0,43
TRANSPORTATION ViDeBERTa_base XLSR-53-Viet SLUE 0,00 0,00 0,00
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,00 0,00 0,00
ViDeBERTa_base w2v2-Viet SLUE 0,00 0,00 0,00
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,00 0,00 0,00
ViT5_base XLSR-53-Viet SLUE 0,00 0,00 0,00
ViT5_base XLSR-53-Viet Mod. SLUE 0,00 0,00 0,00
ViT5_base w2v2-Viet SLUE 0,00 0,00 0,00
ViT5_base w2v2-Viet Mod. SLUE 0,00 0,00 0,00
mBART-50 XLSR-53-Viet SLUE 0,00 0,00 0,00
mBART-50 XLSR-53-Viet Mod. SLUE 0,00 0,00 0,00
mBART-50 w2v2-Viet SLUE 0,00 0,00 0,00
mBART-50 w2v2-Viet Mod. SLUE 0,00 0,00 0,00
BARTpho XLSR-53-Viet SLUE 0,17 0,50 0,25
BARTpho XLSR-53-Viet Mod. SLUE 0,17 0,50 0,25
BARTpho w2v2-Viet SLUE 0,00 0,00 0,00
BARTpho w2v2-Viet Mod. SLUE 0,00 0,00 0,00
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,50 1,00 0,67
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,50 1,00 0,67
PhoBERT_base-v2 w2v2-Viet SLUE 0,00 0,00 0,00
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,00 0,00 0,00
PhoBERT_base XLSR-53-Viet SLUE 0,40 1,00 0,57
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,40 1,00 0,57
PhoBERT_base w2v2-Viet SLUE 0,00 0,00 0,00
PhoBERT_base w2v2-Viet Mod. SLUE 0,00 0,00 0,00
PhoBERT_large XLSR-53-Viet SLUE 0,50 1,00 0,67
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,50 1,00 0,67
PhoBERT_large w2v2-Viet SLUE 0,00 0,00 0,00
PhoBERT_large w2v2-Viet Mod. SLUE 0,00 0,00 0,00
XLM-R_base XLSR-53-Viet SLUE 0,00 0,00 0,00
XLM-R_base XLSR-53-Viet Mod. SLUE 0,00 0,00 0,00
XLM-R_base w2v2-Viet SLUE 0,00 0,00 0,00
XLM-R_base w2v2-Viet Mod. SLUE 0,00 0,00 0,00
XLM-R_large XLSR-53-Viet SLUE 0,50 1,00 0,67
XLM-R_large XLSR-53-Viet Mod. SLUE 0,50 1,00 0,67
XLM-R_large w2v2-Viet SLUE 0,00 0,00 0,00
XLM-R_large w2v2-Viet Mod. SLUE 0,00 0,00 0,00
TREATMENT ViDeBERTa_base XLSR-53-Viet SLUE 0,56 0,77 0,65
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,57 0,78 0,66
ViDeBERTa_base w2v2-Viet SLUE 0,58 0,75 0,65
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,58 0,76 0,66
ViT5_base XLSR-53-Viet SLUE 0,45 0,62 0,52
ViT5_base XLSR-53-Viet Mod. SLUE 0,46 0,63 0,53
ViT5_base w2v2-Viet SLUE 0,44 0,62 0,52
ViT5_base w2v2-Viet Mod. SLUE 0,45 0,64 0,53
mBART-50 XLSR-53-Viet SLUE 0,19 0,08 0,11
mBART-50 XLSR-53-Viet Mod. SLUE 0,23 0,09 0,13
mBART-50 w2v2-Viet SLUE 0,30 0,12 0,17
mBART-50 w2v2-Viet Mod. SLUE 0,34 0,13 0,19
BARTpho XLSR-53-Viet SLUE 0,56 0,71 0,62
BARTpho XLSR-53-Viet Mod. SLUE 0,57 0,72 0,64
BARTpho w2v2-Viet SLUE 0,58 0,71 0,64
BARTpho w2v2-Viet Mod. SLUE 0,60 0,73 0,66
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,54 0,77 0,64
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,55 0,78 0,65
PhoBERT_base-v2 w2v2-Viet SLUE 0,55 0,76 0,64
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,56 0,77 0,65
PhoBERT_base XLSR-53-Viet SLUE 0,56 0,77 0,65
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,58 0,79 0,67
PhoBERT_base w2v2-Viet SLUE 0,57 0,77 0,65
PhoBERT_base w2v2-Viet Mod. SLUE 0,58 0,78 0,66
PhoBERT_large XLSR-53-Viet SLUE 0,54 0,77 0,64
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,55 0,79 0,65
PhoBERT_large w2v2-Viet SLUE 0,54 0,76 0,63
PhoBERT_large w2v2-Viet Mod. SLUE 0,55 0,77 0,64
XLM-R_base XLSR-53-Viet SLUE 0,46 0,76 0,57
XLM-R_base XLSR-53-Viet Mod. SLUE 0,47 0,78 0,58
XLM-R_base w2v2-Viet SLUE 0,45 0,76 0,56
XLM-R_base w2v2-Viet Mod. SLUE 0,46 0,77 0,57
XLM-R_large XLSR-53-Viet SLUE 0,60 0,77 0,68
XLM-R_large XLSR-53-Viet Mod. SLUE 0,61 0,78 0,69
XLM-R_large w2v2-Viet SLUE 0,60 0,76 0,67
XLM-R_large w2v2-Viet Mod. SLUE 0,61 0,77 0,68
UNITCALIBRATOR ViDeBERTa_base XLSR-53-Viet SLUE 0,24 0,34 0,28
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,29 0,42 0,34
ViDeBERTa_base w2v2-Viet SLUE 0,25 0,34 0,29
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,29 0,40 0,34
ViT5_base XLSR-53-Viet SLUE 0,25 0,38 0,30
ViT5_base XLSR-53-Viet Mod. SLUE 0,28 0,42 0,33
ViT5_base w2v2-Viet SLUE 0,26 0,38 0,31
ViT5_base w2v2-Viet Mod. SLUE 0,28 0,42 0,34
mBART-50 XLSR-53-Viet SLUE 0,26 0,05 0,09
mBART-50 XLSR-53-Viet Mod. SLUE 0,41 0,08 0,14
mBART-50 w2v2-Viet SLUE 0,28 0,06 0,10
mBART-50 w2v2-Viet Mod. SLUE 0,36 0,08 0,13
BARTpho XLSR-53-Viet SLUE 0,34 0,41 0,37
BARTpho XLSR-53-Viet Mod. SLUE 0,37 0,44 0,40
BARTpho w2v2-Viet SLUE 0,35 0,42 0,39
BARTpho w2v2-Viet Mod. SLUE 0,38 0,45 0,41
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,35 0,55 0,43
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,40 0,63 0,49
PhoBERT_base-v2 w2v2-Viet SLUE 0,39 0,55 0,45
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,42 0,60 0,50
PhoBERT_base XLSR-53-Viet SLUE 0,34 0,52 0,41
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,39 0,59 0,47
PhoBERT_base w2v2-Viet SLUE 0,36 0,53 0,43
PhoBERT_base w2v2-Viet Mod. SLUE 0,40 0,59 0,47
PhoBERT_large XLSR-53-Viet SLUE 0,36 0,54 0,43
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,40 0,60 0,48
PhoBERT_large w2v2-Viet SLUE 0,38 0,55 0,45
PhoBERT_large w2v2-Viet Mod. SLUE 0,41 0,59 0,49
XLM-R_base XLSR-53-Viet SLUE 0,32 0,51 0,40
XLM-R_base XLSR-53-Viet Mod. SLUE 0,36 0,57 0,44
XLM-R_base w2v2-Viet SLUE 0,34 0,51 0,41
XLM-R_base w2v2-Viet Mod. SLUE 0,37 0,55 0,44
XLM-R_large XLSR-53-Viet SLUE 0,36 0,52 0,42
XLM-R_large XLSR-53-Viet Mod. SLUE 0,40 0,58 0,47
XLM-R_large w2v2-Viet SLUE 0,40 0,55 0,46
XLM-R_large w2v2-Viet Mod. SLUE 0,42 0,58 0,49
Overral Macro ViDeBERTa_base XLSR-53-Viet SLUE 0,34 0,31 0,30
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,38 0,34 0,33
ViDeBERTa_base w2v2-Viet SLUE 0,39 0,30 0,29
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,43 0,33 0,33
ViT5_base XLSR-53-Viet SLUE 0,48 0,42 0,42
ViT5_base XLSR-53-Viet Mod. SLUE 0,52 0,45 0,46
ViT5_base w2v2-Viet SLUE 0,49 0,42 0,43
ViT5_base w2v2-Viet Mod. SLUE 0,53 0,46 0,46
mBART-50 XLSR-53-Viet SLUE 0,28 0,04 0,07
mBART-50 XLSR-53-Viet Mod. SLUE 0,36 0,06 0,09
mBART-50 w2v2-Viet SLUE 0,29 0,05 0,08
mBART-50 w2v2-Viet Mod. SLUE 0,38 0,06 0,10
BARTpho XLSR-53-Viet SLUE 0,52 0,49 0,48
BARTpho XLSR-53-Viet Mod. SLUE 0,57 0,53 0,51
BARTpho w2v2-Viet SLUE 0,51 0,46 0,47
BARTpho w2v2-Viet Mod. SLUE 0,56 0,50 0,50
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,56 0,59 0,55
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,59 0,63 0,58
PhoBERT_base-v2 w2v2-Viet SLUE 0,54 0,53 0,51
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,57 0,57 0,54
PhoBERT_base XLSR-53-Viet SLUE 0,54 0,58 0,53
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,57 0,62 0,56
PhoBERT_base w2v2-Viet SLUE 0,52 0,52 0,50
PhoBERT_base w2v2-Viet Mod. SLUE 0,55 0,56 0,53
PhoBERT_large XLSR-53-Viet SLUE 0,56 0,58 0,54
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,59 0,62 0,57
PhoBERT_large w2v2-Viet SLUE 0,53 0,53 0,50
PhoBERT_large w2v2-Viet Mod. SLUE 0,57 0,56 0,53
XLM-R_base XLSR-53-Viet SLUE 0,48 0,47 0,45
XLM-R_base XLSR-53-Viet Mod. SLUE 0,52 0,51 0,49
XLM-R_base w2v2-Viet SLUE 0,48 0,47 0,45
XLM-R_base w2v2-Viet Mod. SLUE 0,51 0,51 0,49
XLM-R_large XLSR-53-Viet SLUE 0,58 0,59 0,56
XLM-R_large XLSR-53-Viet Mod. SLUE 0,61 0,62 0,59
XLM-R_large w2v2-Viet SLUE 0,55 0,53 0,52
XLM-R_large w2v2-Viet Mod. SLUE 0,58 0,56 0,55
Overral Micro ViDeBERTa_base XLSR-53-Viet SLUE 0,45 0,34 0,39
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0,51 0,39 0,45
ViDeBERTa_base w2v2-Viet SLUE 0,45 0,34 0,39
ViDeBERTa_base w2v2-Viet Mod. SLUE 0,51 0,39 0,44
ViT5_base XLSR-53-Viet SLUE 0,52 0,46 0,48
ViT5_base XLSR-53-Viet Mod. SLUE 0,58 0,50 0,54
ViT5_base w2v2-Viet SLUE 0,53 0,46 0,49
ViT5_base w2v2-Viet Mod. SLUE 0,59 0,51 0,54
mBART-50 XLSR-53-Viet SLUE 0,35 0,05 0,09
mBART-50 XLSR-53-Viet Mod. SLUE 0,46 0,07 0,12
mBART-50 w2v2-Viet SLUE 0,35 0,05 0,09
mBART-50 w2v2-Viet Mod. SLUE 0,47 0,07 0,12
BARTpho XLSR-53-Viet SLUE 0,56 0,50 0,53
BARTpho XLSR-53-Viet Mod. SLUE 0,61 0,55 0,58
BARTpho w2v2-Viet SLUE 0,55 0,50 0,52
BARTpho w2v2-Viet Mod. SLUE 0,61 0,55 0,58
PhoBERT_base-v2 XLSR-53-Viet SLUE 0,57 0,57 0,57
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0,62 0,61 0,62
PhoBERT_base-v2 w2v2-Viet SLUE 0,58 0,56 0,57
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0,62 0,61 0,62
PhoBERT_base XLSR-53-Viet SLUE 0,56 0,56 0,56
PhoBERT_base XLSR-53-Viet Mod. SLUE 0,61 0,61 0,61
PhoBERT_base w2v2-Viet SLUE 0,56 0,56 0,56
PhoBERT_base w2v2-Viet Mod. SLUE 0,61 0,60 0,61
PhoBERT_large XLSR-53-Viet SLUE 0,57 0,55 0,56
PhoBERT_large XLSR-53-Viet Mod. SLUE 0,62 0,60 0,61
PhoBERT_large w2v2-Viet SLUE 0,58 0,55 0,56
PhoBERT_large w2v2-Viet Mod. SLUE 0,62 0,59 0,61
XLM-R_base XLSR-53-Viet SLUE 0,54 0,52 0,53
XLM-R_base XLSR-53-Viet Mod. SLUE 0,59 0,57 0,58
XLM-R_base w2v2-Viet SLUE 0,54 0,52 0,53
XLM-R_base w2v2-Viet Mod. SLUE 0,59 0,57 0,58
XLM-R_large XLSR-53-Viet SLUE 0,60 0,56 0,58
XLM-R_large XLSR-53-Viet Mod. SLUE 0,64 0,60 0,62
XLM-R_large w2v2-Viet SLUE 0,60 0,56 0,58
XLM-R_large w2v2-Viet Mod. SLUE 0,64 0,60 0,62
Figure 1: Scatter plot of NER results on reference text by entity types using various pre-trained language models, created by Table LABEL:appx_NER_results_byLabel.
Figure 2: Scatter plot of NER results on ASR output by entity types using various pre-trained language models, created by Table LABEL:appx_spk_NER_results_byLabel_.