### v3

# Medical Spoken Named Entity Recognition

Abstract
Spoken Named Entity Recognition (NER) aims to extract named entities from speech and categorise them into types like person, location, organization, etc. In this work, we present VietMed-NER - the first spoken NER dataset in the medical domain. To our knowledge, our Vietnamese real-world dataset is the largest spoken NER dataset in the world regarding the number of entity types, featuring 18 distinct types. Furthermore, we present baseline results using various state-of-the-art pre-trained models: encoder-only and sequence-to-sequence; and conduct quantitative and qualitative error analysis. We found that pre-trained multilingual models generally outperform monolingual models on reference text and ASR output and encoders outperform sequence-to-sequence models in NER tasks. By translating the transcripts, the dataset can also be utilised for text NER in the medical domain in other languages than Vietnamese. All code, data and models are publicly available [^1] https://github.com/leduckhai/MultiMed/tree/master/VietMed-NER.
Medical Spoken Named Entity Recognition
Khai Le-Duc1,2, David Thulke4,5, Hung-Phong Tran3, Long Vo-Dang7, Khai-Nguyen Nguyen8, Truong-Son Hy6, Ralf Schlüter4,5 1University of Toronto, Canada 2University Health Network, Canada 3Hanoi University of Science and Technology, Vietnam 4Machine Learning and Human Language Technology Group, RWTH Aachen University, Germany 5AppTek GmbH, Germany 6University of Alabama at Birmingham, United States 7University of Cincinnati, United States 8College of William and Mary, United States duckhai.le@mail.utoronto.ca thy@uab.edu, {thulke,schlueter}@hltpr.rwth-aachen.de
1 Introduction
Named Entity Recognition (NER) targets extracting named entities (NE) from text and categorizing them into types like person, location, organization, etc. Initially studied in written language, recent attention has turned to study spoken NER Cohn et al. (2019); Shon et al. (2022), which aims to extract semantic information from speech. However, spoken NER has limited literature compared to NER on written text data Yadav et al. (2020).
Spoken NER is particularly challenging, firstly due to the impact of word segmentation on results. The medical vocabulary poses difficulties with numerous confused monosyllabic and polysyllabic words. For instance, the word "đường" alone could denote "sugar" (chemical), "street" (location), or be part of a compound word like "đường tiêu hóa" - "gastrointestinal" (anatomy). This confusion has also been reported in Chinese spoken NER by Chen et al. (2022). Further, data quality control and annotation consistency have been problematic, with some entities tagged in one sentence but not in others, and full NEs inconsistently tagged as multiple sub-NEs Huyen and Luong (2016); Nguyen et al. (2018, 2020); Truong et al. (2021). Finally, obtaining accurate medical NER from natural speech is challenging due to the lack of punctuation Ertopçu et al. (2017), speech disfluencies Kim and Woodland (2000), context, and the complexity of medical terms.
As for the medical domain, to the best of our knowledge, there is no dataset available for medical spoken NER. The only related work we found, Cohn et al. (2019), published a NER evaluation benchmark using an English general-domain conversational dataset, Switchboard Godfrey et al. (1992) and Fisher Cieri et al. (2004), for the task of audio de-identification specifically targeting Personal Health Identifiers.
To address this gap, we introduce VietMed-NER, a medical spoken NER dataset built on the real-world medical Automatic Speech Recognition (ASR) dataset VietMed Le-Duc (2024), featuring 18 medically-defined entity types. In the era of the advanced in-context learning capabilities of Large Language Models (LLMs) and human-level text-to-speech technologies, the dataset, with entity positional labels maintained during translation, is applicable not only to Vietnamese but also to other languages (see Appendix C). This enables various real-world applications, including: search engines Rüd et al. (2011), content classification for news providers Kumaran and Allan (2004), medical ASR error correction Mani et al. (2020), audio de-identification Cohn et al. (2019) and content recommendation systems Koperski et al. (2017).
Our contributions are as follows:
  • We present VietMed-NER - the first publicly-available medical spoken NER dataset.
  • We present baselines on several state-of-the-art pre-trained models
  • We conduct quantitative and qualitative error analysis for medical spoken NER in Vietnamese
All code, data and models are published online11footnotemark: 1.
Definition Train Dev Test All
Total Uni. Total Uni. Total Uni. Total Uni.
AGE Age of a person 447 43 108 25 611 83 1166 151
GENDER Gender of a person 202 30 46 15 451 33 699 78
JOB Job of a person 543 32 133 16 562 43 1238 91
LOCATION Locations and places 284 66 76 31 317 75 677 172
ORGANIZATION Organizations 19 14 2 2 58 23 79 39
DISEASESYMTOM Symptoms and diseases 2699 518 683 209 1334 357 4716 1084
DRUGCHEMICAL Bio-chemical substances and drugs 1054 255 263 104 684 136 2001 495
FOODDRINK Food and beverage 243 77 48 26 247 43 538 146
ORGAN Anatomical features, e.g. organs, cells 1827 252 444 122 1190 172 3461 546
PERSONALCARE Personal care, e.g. hygiene routines, skin care 353 114 82 38 95 10 530 162
DIAGNOSTICS Diagnostic procedures, e.g. lab tests, imaging 371 53 91 25 292 36 754 114
TREATMENT Non-surgical treatment, e.g. rehab., injection 726 69 174 25 230 17 1130 111
SURGERY Surgical procedures, e.g. implants, neurosurgery 197 29 55 13 270 37 522 79
PREVENTIVEMED Preventive medicine 341 53 80 25 18 6 439 84
MEDDEVICETECHNIQUE Medical devices, instruments, and techniques 324 84 67 30 603 144 994 258
UNITCALIBRATOR Medical calibration, e.g. number of doses, calories 800 155 215 75 251 106 1266 336
TRANSPORTATION Means of transportation 5 2 3 3 27 10 35 15
DATETIME Date and time 674 155 159 65 657 133 1490 353
#Entities in total 11109 2001 2729 849 7897 1464 21735 4314
#Sentences 4620 1150 3500 9270
Table 1: Entity definition and its statistics in our dataset. "Uni." means the number of unique entities.
2 Related Works
Traditionally, spoken NER has been done using a pipeline methodology, also known as cascaded approach, starting with an ASR stage, followed by NER applied to the generated transcriptions Jannet et al. (2017); Benaicha et al. (2024). Another variant of the cascaded approach involves embedding specific entity expressions into the lexicon, thereby improving the language model’s ability to accurately recognize these expressions Hatmi et al. (2013).
Besides, end-to-end NER has recently garnered some attention within the research community. This approach seeks to optimize ASR and NER processes simultaneously, offering a potentially more efficient alternative to traditional pipeline methods by harnessing the ability of trainable acoustic features. However, its accuracy advantage over the cascaded approach remains a subject of debate, and the end-to-end training setup introduces additional complexity Tomashenko et al. (2019); Yadav et al. (2020).
3 Data
3.1 Data Collection
We chose the VietMed dataset Le-Duc (2024), the world’s largest publicly available medical ASR dataset, for annotating NEs.
The original dataset is in Vietnamese. We annotate the Vietnamese version with the methodology described in Section 3.2 and automatically translate the transcripts to English together with transferring the NE annotation.
3.2 Annotation Process
The annotation of medical NEs from real-world speech is challenging because of the missing punctuation, special characters and capitalized words in ASR transcripts, disfluencies and required medical knowledge. Entirely manual annotation of NEs like in VLSP dataset Huyen and Luong (2016); Nguyen et al. (2018, 2020) and PhoNER_COVID19 Truong et al. (2021) requires a large number of working hours, not to mention the difficulties in quality control and inconsistency as we found in their corpora. These inconsistencies include: i) Some entities tagged in one sentence are not tagged in another sentence, and ii) Full NEs are inconsistently tagged as multiple sub-NEs. The best approach to tag nested NEs is the subject of ongoing debate Muis and Lu (2017); Li et al. (2021a). For simplicity, higher consistency and to reduce the annotation effort, we only annotate the largest and outermost full entity span.
Moreover, using fine-tuned models for pre-tagging doesn’t apply to specific medical entity types. Similarly, using prompt engineering with large language models like GPT-4 for pre-tagging did not achieve acceptable accuracy. Training a seed model with a gazetteer list requires initial training time, subsequent repetitive training schedules, and may prove unreliable due to its statistical reliance on a small amount of low-resource data Kozareva (2006).
To tackle these problems, we conduct a human-machine annotation approach, as described below:
  1. 1.
    Annotate and categorize a set of initial entities, then add them to a gazetteer list.
  2. 2.
    Sort entities by character length from highest to lowest, to distinguish between sub-NEs and full NEs, ensuring full NEs are mapped before sub-NEs. Time complexity = $O(k\cdot\log(k))$ where $k$ is the number of NEs. For example, "tooth pain" should be mapped before "pain".
  3. 3.
    Automatically map entities from the gazetteer list to the transcript. Time complexity = $O(m\cdot n)$, where $m$ is the number of NEs in gazeeter list, $n$ is the number of sentences. Pseudo code: {python} for NE in gazetteer_list: for sen in sentences: if NE in sen: annotate(NE, sen)
  4. 4.
    Annotators review each sentence to include correctly labeled NEs and ignore mislabeled NEs
  5. 5.
    Annotators add new NEs not in the gazetteer list during manual annotation. Steps 2 and 3 generate pre-tagged labels in the next sentences. Annotators repeat Steps 4 and 5 until the entire corpus is annotated.
We experience faster annotation by allowing annotators to foresee possible NEs in upcoming utterances based on previously annotated ones. Annotators can accept or reject these suggestions, saving time with correct suggestions and easily ignoring incorrect ones. Unlike training a seed model with a gazetteer, which requires initial training time and may be unreliable for low-to-mid resource languages, our method avoids these issues and eliminates the need to correct incorrect NEs.
3.3 Data Quality Control
We created initial annotation guidelines (see Appendix A) and began annotating the corpus. Two developers, one with a medical background, independently annotated the corpus. Then, we held a discussion session to resolve conflicts, address complex cases, and refine the guidelines. Two other developers perform quality control using the guidelines and the annotated corpus. We consistently revisited each sentence in the entire corpus multiple times. This data quality control process is inspired by Tran et al. (2022).
3.4 Data Splitting
Most NER datasets have a very small number of entities in their test sets compared to train and dev set Huyen and Luong (2016); Truong et al. (2021); Chen et al. (2022). However, we want to leverage the capabilities of large pre-trained models which are trained on vast amounts of unlabeled text data, resulting in good representations. Therefore, we focus on creating a large test set to obtain more statistically significant evaluation results and keep the training set relatively small in comparison.
3.5 Data Statistics
Table 1 shows the statistics of our dataset. Our VietMed-NER contains 18 entity types across 9000 sentences, split into train-dev-test as 8-2-6 hours. To the best of our knowledge, compared to all other public spoken NER datasets, ours has the largest number of entity types.
4 Experimental Setups
We employ the cascaded (two-stage) pipeline for spoken NER: A hybrid ASR model transcribes audio into text and then the transcribed text is fed into a text NER model.
4.1 Evaluation Metrics
We employed the F1 score metric as it is commonly used for spoken NER Shon et al. (2022); Benaicha et al. (2024), which evaluates an unordered list of NE phrases and tag pairs predicted for each sentence. We used 3 toolkits for a more comprehensive comparison, as described in Appendix D.
4.2 ASR Models
We employed two baseline models fine-tuned for ASR on VietMed published by Le-Duc (2024): an acoustic monolingual pre-trained w2v2-Viet and an acoustic multilingual pre-trained XLSR-53-Viet model. w2v2-Viet model was pre-trained from scratch on 1204h of Vietnamese data. For the XLSR-53-Viet model, continued pre-training on 1204h of Vietnamese starting with XLSR-53 Conneau et al. (2021) was performed. Both have the same number of parameters (118M) and were fine-tuned on the same training set. Their WERs on the test set are 29.0% and 28.8% respectively.
4.3 NER Models
Model #Params #Data
PhoBERT_base 135M 20GB
PhoBERT_large 370M
PhoBERT_base-v2 135M 140GB
ViDeBERTa_base 86M 298GB
XLM-R_base 270M 2.5TB
XLM-R_large 550M 2.5TB
mBART-50 611M 3.9TB
ViT5_base 310M  [^8] GB
BARTpho 396M 20GB
Table 2: Statistics of state-of-the-art pre-trained language models which we used for NER task.
Table 2 shows the statistics of various pre-trained monolingual and multilingual models we consider to fine-tune on our dataset. To our knowledge, these are the best pre-trained models that achieved state-of-the-art results on various downstream tasks in the Vietnamese language, including NER.
Monolingual encoder models: PhoBERT_base, PhoBERT_large, PhoBERT_base-v2 Nguyen and Tuan Nguyen (2020), ViDeBERTa_base Tran et al. (2023).
Monolingual sequence-to-sequence (seq2seq) models: BARTpho Tran et al. (2022), ViT5 Phan et al. (2022),
Multilingual encoder models: XML-R_base, XML-R_large Conneau et al. (2020).
Multilingual seq2seq models: mBART-50 Tang et al. (2020).
4.3.1 Seq2seq Training for NER Task
Following the approach proposed by Phan et al. (2021) and later adopted by ViT5 Phan et al. (2022), we formulated the sequence tagging task as a sequence-to-sequence task by training the models to generate tags of labels before and after an entity token. In cases where the models fail to follow the mentioned format for an entity token, we use an “exception” tag, which will be later ignored during metric calculation, as the label.
4.3.2 Training Hyperparameters
We used HuggingFace Transformers Wolf et al. (2019) for fine-tuning pre-trained models for the NER task. Vietnamese input sentences can be represented in either syllable or word level as described by Truong et al. (2021). However, we only employed word-level settings to train NER models. All our NER experiments were done by using the default hyperparameters by HuggingFace.
The default hyperparameters are as follows: Learning rate of 2e-5, linear learning rate scheduler, training batch size of 64, 50 training epochs, weight decay of 0.01, AdamW optimizer Loshchilov and Hutter (2019), Beta1 of 0.9, Beta2 of 0.999, and epsilon of 1e-8.
5 Experimental Results
NER Model Prec. Rec. F1
BARTpho 0.64 0.73 0.68
mBART-50 0.64 0.66 0.65
PhoBERT_base 0.67 0.78 0.72
PhoBERT_base-v2 0.68 0.79 0.74
PhoBERT_large 0.69 0.77 0.73
ViDeBERTa_base 0.50 0.41 0.45
ViT5_base 0.64 0.74 0.69
XLM-R_base 0.64 0.73 0.69
XLM-R_large 0.71 0.77 0.74
Table 3: NER results on reference text of test set. The metrics shown are Precision, Recall, and overall micro F1 score. Results by entity types are shown in Tables 5-24 in the Appendix.
NER ASR Prec. Rec. F1
ViDeBERTa_base XLSR-53-Viet 0.45 0.34 0.39
w2v2-Viet 0.45 0.34 0.39
ViT5_base XLSR-53-Viet 0.52 0.46 0.48
w2v2-Viet 0.53 0.46 0.49
mBART-50 XLSR-53-Viet 0.35 0.05 0.09
w2v2-Viet 0.35 0.05 0.09
BARTpho XLSR-53-Viet 0.56 0.50 0.53
w2v2-Viet 0.55 0.50 0.52
PhoBERT_base_v2 XLSR-53-Viet 0.57 0.57 0.57
w2v2-Viet 0.58 0.56 0.57
PhoBERT_base XLSR-53-Viet 0.56 0.56 0.56
w2v2-Viet 0.56 0.56 0.56
PhoBERT_large XLSR-53-Viet 0.57 0.55 0.56
w2v2-Viet 0.58 0.55 0.56
XLM-R_base XLSR-53-Viet 0.54 0.52 0.53
w2v2-Viet 0.54 0.52 0.53
XLM-R_large XLSR-53-Viet 0.60 0.56 0.58
w2v2-Viet 0.60 0.56 0.58
Table 4: NER results on ASR output of test set for different NER and ASR models. Metrics shown are Precision, Recall, and overall micro F1 score. Results by entity types are shown in Tables 25-44 in the Appendix.
Table 3 and 4 show results of NER using various pre-trained models. We observe that there was a performance drop in all models when evaluated on ASR transcripts, as expected due to the noisy nature of ASR output.
1. Pre-trained multilingual models outperformed monolingual models, if multilingual models overcome the capacity dilution: The pre-trained monolingual model PhoBERT_base-v2 outperformed other monolingual models, at 0.74 of F1 score on reference text, and 0.57 on ASR output. Despite having fewer parameters than PhoBERT_large, it performed similarly, likely due to more pre-training data. The pre-trained multilingual model XLM-R_large achieved the best performance with an F1 score of 0.74 on reference text and 0.58 on ASR output, while XLM-R_base performed worse than PhoBERT_base-v2. This gap is explained by the larger pre-training data (2.5TB multilingual data for XLM-R vs. 140GB monolingual data for PhoBERT_base-v2). Our results with PhoBERT_base-v2 and XLM-R_large confirmed that pre-trained multilingual representations improve performance on medical spoken NER tasks, similar to other language-specific downstream tasks by Conneau et al. (2020); Liu et al. (2020). However, multilingual models may face a Transfer-dilution Trade-off Conneau et al. (2020), where they lack the capacity to learn effective multilingual representations. In other words, for a fixed sized model, the per-language performance decreases as we increase the number of languages Gurgurov et al. (2024). To address this trade-off, multilingual models should possess sufficient capacity, necessitating an adequately large model size Chen and Chen (2024). This is evident in the performance comparison between PhoBERT_base-v2 and XLM-R_base, as seen in other language-specific downstream tasks by Conneau et al. (2020); Arivazhagan et al. (2019).
2. Encoder-based models outperform seq2seq models: The best seq2seq model, BARTpho, achieved F1 scores of 0.68 on reference text and 0.53 on ASR output. Encoders generally performed better than seq2seq models, possibly because seq2seq’s generative nature is less suited for classification tasks like NER.
3. Multi-lingual pre-training of the acoustic model does not affect cascaded NER performance
As expected by the similar WERs for the acoustic pre-trained monolingual model w2v2-Viet and the multilingual model XLSR-53-Viet, all NER models show comparable F1 scores, precision, and recall. This indicates that in addition to overall WER the models do not differ significantly in the recognition accuracy of medical NEs. Non-cascaded models might have advantages in utilising the additional pre-training data for the downstream task.
6 Error Analysis
We performed an error analysis using the best-performing models.
Figure 1: Scatter plot of NER results on reference text by entity types using various pre-trained language models and evaluation variants, created by Tables 5-24.
Figure 2: Scatter plot of NER results on ASR output by entity types using various pre-trained language models and evaluation variants, created by Tables 25-44.
6.1 Quantitative
We provide a detailed error analysis for each entity type across best models, utilizing three evaluation toolkits. The results are summarized in Tables 5-24 and Tables 25-44, with corresponding visual representations in the scatter plots shown in Figure 1 and Figure 2.
The top NER models showed high accuracy in recognizing OCCUPATION and TRANSPORTATION entities in both reference text and ASR output. Despite TRANSPORTATION having only 35 total and 15 unique samples, the best models performed well, likely due to the semantic clarity and predictable spans of these entities.
In contrast, PREVENTIVEMED showed higher misrecognition rates, despite sufficient sample size mitigating class imbalance. This may stem from two factors. First, preventive medicine terms often overlap with general medical terminology, making it difficult for the model to distinguish them from DRUGCHEMICAL or TREATMENT concepts. For example, "vaccination" is frequently misclassified as a therapeutic intervention (TREATMENT) in sentences that describe its role in disease prevention (PREVENTIVEMED). Also, models frequently struggle to differentiate between "vaccination" and "vaccine" (DRUGCHEMICAL). Second, preventive medicine involves long-term health strategies often expressed in non-clinical or non-standardized language, complicating entity recognition.
6.2 Qualitative
A common error was confusion between LOCATION and ORGANIZATION, due to the inherent ambiguity where the same entity can function as either depending on context. An organization-related entity may be labelled as LOCATION if it implies a patient visited there, but this inference requires external knowledge about the entity. Another confusion involved DRUGCHEMICAL and FOODDRINK. Both categories share similar names, descriptors, and consumption contexts (e.g., caffeine, alcohol, sugar). Insufficient context length often causes errors, especially with ambiguous terms like "vitamin," "cordyceps," or "sea daffodils," which can refer to both supplements and nutrients depending on context. Another case is DIAGNOSTICS, TREATMENT, and SURGERY. For example, a "biopsy" can be both a diagnostic and treatment, while "radiation therapy" may be linked to surgery.
A common error in NER involves incorrect entity spans, which fall into two types: (1) correct label but wrong span, and (2) wrong label but correct span. The first type often occurs with multi-word entities in the medical domain, like ORGANIZATION, LOCATION, or DISEASESYMPTOM. For example, "high blood pressure" (B-DISEASESYMPTOM, I-DISEASESYMPTOM, I-DISEASESYMPTOM) may be misrecognized as "high blood" (B-DISEASESYMPTOM, I-DISEASESYMPTOM, O), keeping the meaning but shortening the span. The second type occurs when compound-word entities are split, such as "vagina cells" (B-ORGAN, I-ORGAN) being misrecognized as "vagina" and "cells" (B-ORGAN, B-ORGAN).
7 Conclusion
In this work, we present VietMed-NER - the first spoken NER dataset in the medical domain. Our dataset contains 18 entity types, including both conventional and newly defined entity types for real-world medical conversations. Our results show that pre-trained multilingual models typically outperform monolingual models on both reference text and ASR output if the multilingual models are sufficiently large to learn multilingual representations. Additionally, encoders generally demonstrate better performance than seq2seq models in the NER task. Finally, while pre-trained audio data impacts ASR output, it does not significantly impact NER performance in the cascaded setting.
8 Limitations
Our annotation approach: Our annotation approach has some advantages over the fully manual approach. First, it allows annotators to not spend extra time tagging the entities that have been tagged in previous sentences. Second, it prevents that annotators miss entities that have been tagged in previous sentences, improving the consistency of the entire dataset. During our work, we experienced a faster annotation by using our approach compared to fully manual annotation. However, in the scope of this paper, we have not done extensive experiments to give a quantitative number of how much time has been saved and the method’s impact on annotation quality.
Evaluation metrics for medical terms: ASR system performance is commonly evaluated using WER, which quantifies the ratio of word insertion, substitution, and deletion errors in a transcript relative to the total number of spoken words. However, various spoken language understanding tasks, such as spoken NER, rely on accurately identifying key terms within transcripts. In medical ASR, it is critical to account for the disproportionate importance of medical terms in doctor-patient interactions, as they hold significantly more weight than general vocabulary, as discussed in Section B in the Appendix. We believe that other domain-specific spoken NER tasks follow a similar pattern. Consequently, future comprehensive investigations into evaluation metrics are needed to determine the most appropriate metric for spoken NER in the medical and other domains.
Appendix A Annotation Guidelines
This section describes annotation guidelines for annotators to follow in an attempt to have a unified and consistent gold-standard NER transcript.
General rules:
  • If 2 or more entities overlap, label the resulting entity as the longest, including overlapping component entities. In other words, a full NE might contain 2 or more sub-NEs. A full NE should be tagged instead of multiple sub-NEs. For example: "bác sĩ xương khớp" (orthopedic doctor) should be tagged as a whole instead of 2 distinct NEs "bác sĩ " (doctor) and "xương khớp" (orthopedic).
  • We adhere to the conventional approach of annotating overlapping NE components as a whole, utilizing the BIO encoding scheme. In recent years, research on overlapping and discontinuous NER has introduced alternative annotation frameworks to solve NE overlapping, such as BILOU encoding, which represents "Beginning, Inside, and Last tokens of multi-token chunks, Unit-length chunks, and Outside" Straková et al. (2019); Duan et al. (2017). Another approach by Li et al. (2021b) introduces a novel span-based model capable of jointly recognizing both overlapping and discontinuous entities. The model operates in two primary stages. First, entity fragments are identified by systematically traversing all possible text spans, enabling the detection of overlapping entities. Second, a relation classification step determines whether a given pair of entity fragments exhibits an overlapping or successive relationship. This approach facilitates the recognition of discontinuous entities while simultaneously verifying overlapping entities.
  • Do not assign spaces at the beginning and at the end of entities.
  • All words in the ASR transcript are lowercase, without punctuations and special characters. Treat every word as lowercase or uppercase, with or without punctuations and special characters based on the context of each utterance.
  • Each utterance should be treated as an independent utterance. The additional context given by other utterances should not influence the annotation of each utterance.
AGE:
This entity type describes the age of a person.
  • Label the word "tuổi" (age) if applicable. For example: "tuổi trưởng thành" (mature age), "hai bảy tuổi" (twenty-seven years old).
  • List a range of ages if applicable For example: "hai mươi đến ba lăm tuổi" (twenty to thirty-five years old), "dưới sáu tháng tuổi" (under six months old).
  • Include adjectives and nouns that might describe how old a person is but don’t explicitly describe gender or gender is neutral. For example: "chưa trưởng thành" (immature), "người già" (old person), "cụ" (sir, old).
GENDER:
This entity type describes the gender of a person.
  • Include typical entities that are widely understood to describe the gender of a person. For example: "nam" (male), "đàn ông" (gentleman), "phụ nữ" (woman).
  • Include the titles and pronouns that explicitly describe a gender instead of age. For example: "ông" (grandfather), "bà" (grandmother), "cô" (aunt), "chú" (uncle).
OCCUPATION:
This entity type describes the job of a person.
  • Include all jobs that might be both in medical fields and non-medical fields. For example: "khán thính giả" (audience), "bệnh nhân" (patient), "người dân" (citizen), "chuyên gia" (expert).
  • Include academic titles and degrees. For example: "thạc sĩ" (master degree holder), "tiến sĩ" (doctorate), "trưởng khoa" (dean), "chủ tịch" (president).
  • Include a cluster of words that might describe the specializations of doctors. For example: "bác sĩ chuyên về rối loạn vận động" (doctor who specializes in movement disorders) instead of two distinct entities "bác sĩ" (doctor) and "rối loạn vận động" (movement disorders), "bác sĩ về parkinson" (parkinson’s doctor) instead of two distinct entities "bác sĩ" (doctor) and "parkinson", "bác sĩ chuyên khoa tim mạch" (cardiovascular specialist) instead of two distinct entities "bác sĩ" (doctor) and "chuyên khoa tim mạch" (cardiovascular).
LOCATION:
This entity type describes a location.
  • Include continents, countries, regions, cities, and geographical administrative units. For example: "châu âu" (europe), "hoa kỳ" (usa), "tây tạng" (tibet), "thành phố hồ chí minh" (ho chi minh city), "tỉnh vĩnh long" (vinh long province).
  • Label words that mean geographical administrative units if applicable. For example: "huyện" (rural district), "quận" (urban district), "đường phố" (street), "thành phố" (city).
  • Include words that might describe public and private sites. For example: "tại nhà" (at home), "đồng ruộng" (farm), "tiệm thuốc" (drugstore), "nhà máy" (factory), "cửa hiệu quần áo" (clothing store), "toilet" (toilet).
  • Include words that might describe ambient environments. For example: "tại khu phố" (in the neighborhood), "tại địa phương" (in local area), "nước ngoài" (in foreign countries), "địa bàn" (area), "ngoài trời" (outside).
  • Include words that might describe medical facilities. For example: "chuyên khoa tiêu hóa" (gastrointestinal room) , "icu" (intensive care unit), "trạm xá" (clinics), "phòng thí nghiệm" (laboratory).
  • Each level of the administrative unit is a separate entity.
  • Do not assign nationality as an entity.
  • Locations might be misrecognized as organizations. Do not label places that are not clearly identified or controversial.
DISEASESYMPTOM:
This entity type describes a symptom or disease.
  • Include the complements of the disease. For example: "biến chứng" (side-effect), "chấn thương" (damaged), "bẩm sinh" (congenital), "di chứng" (sequelae), "bị tổn thương" (damaged), "tái phát" (relapse), "dương tính" (positive), "bệnh lý mãn tính" (chronic disease), "hội chứng" (syndrome).
  • Include a cluster of words that might describe the severity of a disease. For example: "phỏng cấp độ ba" (third-degree burn), "sức đề kháng kém" (poor immune system).
  • Mental state might also describe mental diseases or their symptoms. For example: "tự ti" (self-deprecation), "tình trạng lo âu" (state of anxiety), "mệt mỏi về tinh thần" (mental fatigue).
  • Skin conditions might describe dermatosis or its symptoms. For example: "nám" (melasma), "da đổ dầu" (oily skin), "da khô" (dry skin), "sạm da" (dark skin).
  • Genital conditions might describe genital diseases or their symptoms. For example: "có kinh" (menstruation), "có thai" (pregnant), "dậy thì sớm" (early puberty).
  • Healthy conditions might help doctors diagnose. For example: "kinh nguyệt đều" (regular menstruation).
  • Words describing physical status might also speak of symptoms or diseases. For example: "buồn ngủ" (sleepy), "rụng tóc" (hair loss), "còi cọc" (stunted).
  • Words describing children’s activities might also speak of pediatric symptoms or diseases. For example: "quấy khóc" (fussy), "không thể giao tiếp" (unable to speak), "chậm đi" (delay walking).
  • Medical techniques or devices might make symptoms and diseases happen. For example: "phẫu thuật thẩm mỹ" (cosmetic surgery).
DRUGCHEMICAL:
This entity type describes a bio-chemical substance or medicament.
  • Extraction of human or animal bodies to serve medical treatment might be referred to as a biochemical substance. For example: "vắcxin" (vaccine), "huyết thanh" (blood serum)
  • Cosmetics might be referred to as chemical substances. For example: "kem chống nắng" (sunscreen), "kem dưỡng ẩm" (moisturizer).
  • Food or drink serving medical treatment purposes or as a part of a chemical compound might be referred to as chemical substances. For example: "nấm đông trùng hạ thảo" (cordyceps), "nhân sâm" (ginseng), "nhung hươu" (deer antler).
  • Substances extracted from cells or bodies not serving medical purposes might be referred to as bio-chemical substances. For example: "dịch tiêu hóa" (digestive fluids), "chất nội sinh" (endogenous substances), "mồ hôi" (sweat), "bã nhờn" (sebum).
  • Air might be referred to as chemical substances. For example: "dưỡng khí" (breath air), "oxy" (oxygen).
FOODDRINK:
This entity type describes food and beverage.
  • Include food and drink that might serve nutrient purposes. For example: "sữa" (milk), "ngũ cốc" (cereal).
  • Include food and drink that might be harmful to health. For example: "thuốc lá" (cigarette), "rượu bia" (alcohol).
  • Include words that generally describe food and beverage. For example: "thực phẩm" (aliment), "thức ăn" (food).
ORGAN:
This entity type describes an anatomical feature, e.g. human organs, biological cells, etc. Annotators should follow general rules.
PERSONALCARE:
This entity type describes a personal care procedure, e.g. hygiene routines, skin care, daily habits, etc.
  • Activities serving the improvement of physical, aesthetic and mental health instead of medical treatment purposes might be referred to as personal care procedures. For example: "ăn kiêng" (diet), "chăm sóc da" (skin care), "chăm sóc răng" (dental care).
  • Methods serving self-improvement of speech ability in speech-language pathology might be referred to as personal care. For example: "tương tác ngôn ngữ" (language interaction), "huấn luyện ngôn ngữ" (language training).
DIAGNOSTICS:
This entity type describes a diagnostic procedure, e.g. lab tests, imaging, blood measurement, etc.
  • General words describing diagnostic procedures without explicitly mentioning surgery might be referred to as diagnostic produces. For example: "chẩn đoán" (diagnosis), "xét nghiệm" (test).
  • Imaging methods might be referred to as diagnostic procedures instead of medical devices or techniques. For example: "mri" (magnetic resonance imaging), "ct" (computed tomography).
TREATMENT:
This entity type describes a non-surgical treatment method for diseases, e.g. physical rehabilitation, injection, psychology, etc.
  • Words describing methods of using bio-chemical substances as non-surgical treatment methods might be referred to as treatment methods. For example: "liệu pháp hoocmon" (hormone therapy), "điều trị hoocmon" (hormone treatment), "điều trị tế bào gốc" (stem cell treatment).
  • Words describing methods of using invasive techniques as treatment methods might be referred to as treatment methods. For example: "hóa trị" (chemotherapy), "xạ trị" (radiotherapy).
  • Words describing methods to improve skin conditions for treatment purposes rather than aesthetics might be referred to as treatment methods. For example: "phục hồi da" (skin recovery), "ức chế sự xuất sắc tố" (inhibit pigmentation).
SURGERY:
This entity type describes a surgical treatment method for diseases, e.g. implants, neurosurgery, invasion, etc.
  • Include pre-surgery procedures that might be integral parts of surgeries. For example: "gây mê" (anesthesia), "gây tê" (anesthetize).
  • Include intervention procedures that might be integral parts of dental care. For example: "nhổ răng" (tooth extraction), "implant" (dental implant).
  • Include intervention procedures that might be integral parts of pregnancy or genitals. For example: "sinh mổ" (caesarean), "cấy tránh thai" (contraceptive implant).
  • Include intervention procedures on arteries even though they might be not integral parts of surgery. For example: "truyền máu" (blood transfusion), "truyền nước biển" (seawater infusion).
  • Include neurosurgical procedures that work with brain waves even though they might be minimally invasive. For example: "kích thích não sâu" (dbs or deep brain stimulation).
MEDDEVICETECHNIQUE:
This entity type describes a medical device, instrument, bio-material and technique.
  • Medical devices and techniques might be confusing. Annotators are strongly recommended to fully annotate CHEM., FnB, ANAT., PC, DX, TX, and SX before engaging TECH.
UNITCALIBRATOR:
This entity type describes a medical calibration, e.g. number of doses, calories, length, volume, etc.
  • Include a cluster of words that both describe the quantity and its unit. Measurements including length, distance, area, weight, heat, velocity, temperature, etc., should be explicitly tagged. For example: "năm milimet" (five millimeters) instead of "năm" (five) or "milimet" (millimeter).
  • Complements to the actual quantity describing its approximation should be included. For example: "khoảng mười lăm phần trăm" (about fifteen percent) instead of "mười lăm phần trăm" (fifteen percent).
  • Include words that generally describe the quantity. For example: "gần đủ" (close enough), "cao" (high), "rất là lớn" (very large).
  • Include words that describe trends of quantity. For example: "giảm được ít nhất" (reduce at least), "mức độ gia tăng" (level increases).
TRANSPORTATION:
This entity type describes means of transportation or vehicles.
DATETIME:
This entity type describes the date and time.
  • Include words describing day, week, month, certain named period, season, year, etc.
  • Include words describing a time frame. For example: "bây giờ" (now), "về lâu về dài" (in the long run).
  • Include words describing the approximate time. For example: "nhanh nhất có thể" (as fast as possible), "càng sớm" (as soon as possible), "từ từ" (gradually).
  • Include words describing repetitions. For example: "định kỳ" (periodically).
  • Include a cluster of words that both describe time and its complements. For example: "từ tháng ba trở đi" (from march onwards) instead of 3 distinct entities "từ" (from), "tháng ba" (march), and "trở đi" (onwards).
Appendix B Discussion about Named-Entity-Error-Rate (NEER)
B.1 Motivation of NEER
ASR system performance is typically assessed using WER, which represents the ratio of word insertion, substitution, and deletion errors in a transcript to the total number of spoken words. However, various spoken language understanding tasks, such as spoken NER, depend on identifying keywords in transcripts. Moreover, it’s essential to recognize that in medical ASR, medical terms carry much higher significance in doctor-patient conversations and should not be treated equally to regular words. KER is often used to evaluate on keywords but is not a directly comparable metric with WER.
The purpose to introduce NEER aims to bridge the gap between WER and KER. However, it is not intended to replace WER or KER as a standard metric for evaluating domain-specific ASR performance. Instead, NEER serves as a complementary metric, facilitating a more in-depth analysis of ASR errors in specific domains, such as the medical field.
B.2 Definition of WER
WER is calculated based on the Levenshtein distance Levenshtein et al. (1966), which represents the smallest count of individual edits (insertions, deletions, or substitutions) needed to transform one word into another.
$WER=\frac{S+D+I}{N}=\frac{S+D+I}{S+D+C}$ (1)
where S is the number of substitutions, D is the number of deletions, I is the number of insertions, C is the number of correct words, and N is the number of words in the reference data $(N=S+D+C)$.
In other words, S is the number of replaced words. D is the number of missed words that are not in ASR hypothesis but are in reference data. I is the number of added words that are in ASR hypothesis but are not in reference data. The alignment between ASR hypothesis and reference data goes from left to right.
B.3 Definition of KER
Like WER, KER is computed using the Levenshtein distance. Each ASR hypothesis is aligned with its corresponding reference data and KER is calculated based on the keyword set.
$KER=\frac{F+M}{N}$ (2)
where N is the number of keywords in the reference data, F is the number of falsely recognized keywords, M is the number of missed keywords.
The ASR hypothesis often exceeds the length of all keywords in the reference data, and the insertion errors caused by non-keywords may lead to a skewed result in KER. Therefore, no insertion errors are considered while calculating KER.
B.4 Definition of NEER
In KER metric, N is the number of keywords in the reference data. KER could be characterized as the average number of errors per keyword. Nevertheless, the length of keywords may range from 1 to L (where L equals 5 in certain instances such as NER), making the average number of errors per keyword obscure.
In NEER metric, we want to evaluate on keyword-only like KER metric, while also analyzing errors per word like WER metric. Therefore, we change N into the length of keywords (entities), which characterizes the average number of errors per word of keywords.
B.5 Open questions on NEER
We still leave some questions open for future work. First, the analysis of how each type of word error (substitutions, insertions, deletions) influences NER on top of ASR has not been conducted yet. Second, the empirical relationship between WER, KER, NEER, and F1 score - meaning how KER, NEER, and F1 score are affected by a varying range of WERs — has not been analyzed either.
Appendix C Possible Applications
In the context of advanced in-context learning capabilities of LLMs and human-level text-to-speech technologies, the dataset, which preserves entity positional labels throughout the translation process, is applicable not only to Vietnamese but also to other languages.
To facilitate dataset translation using LLMs, we employed a concise, brief contextual instruction via GPT-3.5 Turbo, as illustrated below.
In this example, we utilized a straightforward one-shot in-context learning approach, which successfully produced a high-quality translation while preserving all entity positional labels. We hypothesize that employing many-shot learning, extending the context length of instructions, and more advanced LLMs (e.g. GPT-4, GPT-4o) could yield even more enhanced results.
Recent research in text-to-speech (TTS) technology has achieved human-level performance in aspects such as naturalness and emotional expressiveness Tan et al. (2024); Li et al. (2023); Chen et al. (2023). Consequently, the application of TTS to translated NER transcripts could generate realistic audio data, facilitating further research in additional languages.
gpt-3.5-turbo Instruction: Given the text: "thì cũng giống như ba má mình đã từng bị béo phì rồi bị bệnh này bệnh kia những người thân quen của mình mình biết bị" Its coressponding named entity position is: [ "0", "0", "0", "0", "B-GENDER", "B-GENDER", "0", "0", "0", "0", "B-DISEASESYMTOM", "I-DISEASESYMTOM", "0", "B-DISEASESYMTOM", "I-DISEASESYMTOM", "0", "0", "0", "0", "0", "0", "0", "0", "0", "0", "0", "0" ], where "0" is none-entity. Translate to English and give the corresponding named entities. LLM response: The given Vietnamese text translates to: "[it is just like my mom and dad who were once obese and got this disease or that disease, people close to me, I know, have been]". The corresponding named entities, based on the positions provided, are: - "B-GENDER", "B-GENDER" → **"mom" and "dad"** (Gender entities) - "B-DISEASESYMTOM", "I-DISEASESYMTOM" → **"obese"** (Disease/Symptom) - "B-DISEASESYMTOM", "I-DISEASESYMTOM" → **"disease"** (Disease/Symptom) So, the translation with the named entities marked would be: "[it is just like **mom** and **dad** who were once **obese** and got this **disease** or that **disease**, people close to me, I know, have been]".
Appendix D Details about Experimental Setups
D.1 Evaluation Toolkit
We employed seqeval [^2] https://github.com/chakki-works/seqeval framework commonly used as a default evaluation framework by HuggingFace. However, this framework only works for NER on reference text. Therefore, we also employed the F1 score calculation by Shon et al. (2023) by using the SLUE toolkit [^3] https://github.com/asappresearch/slue-toolkit. This F1 score evaluates an unordered list of NE phrase and tag pairs predicted for each sentence. Our proposed modification of SLUE toolkit was also used and presented below.
D.2 Modified Evaluation of SLUE toolkit
Following pre-processing, we calculate the evaluation metrics for the ASR-NER SLUE task. This involves computing precision, recall, and F1-score, which provide insights into the model performance at both an individual label level (per entity) and across all labels (overall).
We introduce a "dummy" token strategy to replace the actual NEs. This approach upholds the focus on the classification of entities rather than the extraction of verbatim phrases, which is suitable for cases where ASR errors might skew the recognition of entities in spoken transcripts.
Let’s take an example:
  • Reference text: "I have a tooth pain"
  • BIO encoding of reference text: [0, 0, 0, B-DISEASESYMTOM, I-DISEASESYMTOM]
  • ASR output: "Has teeth pain"
  • BIO encoding of ASR output: [0, B-DISEASESYMTOM, I-DISEASESYMTOM]
In the SLUE toolkit, the format (NE type, NE) is used to compare reference text and ASR output, e.g. (DISEASESYMTOM, "tooth pain") and (DISEASESYMTOM, "teeth pain"). This format gives an F1 score of 0.0 although entity type is correctly recognized. In our "dummy" token strategy, we modify the format as (NE type, "dummy"), turning reference text and ASR output to (DISEASESYMTOM, "dummy") and (DISEASESYMTOM, "dummy") respectively. The modified format gives a correct F1 score of 1.0.
We compute two types of overall metrics: micro and macro averages. The micro average metrics aggregate the contributions of all classes to compute the average metric, while the macro average computes per-entity type metrics and averages them, without considering the frequency of each entity type. The micro average is therefore influenced by the class distribution and will be dominated by the performance on more frequent entity types. In contrast, macro averages treat all entity types equally, providing a measure of the system’s performance across different types of NEs, regardless of their frequency in the dataset.
Appendix E NER Results by Entity Types
Tables 5-24 show the results of NER on reference text by entity types using various pre-trained language models. Tables 25-44 show the results of NER on ASR output by entity types using various pre-trained language models and ASR models.
Figure 1 shows the scatter plot of NER results on reference text by entity types using various pre-trained language models, created by Tables 5-24. Figure 2 shows the scatter plot of NER results on ASR output by entity types using various pre-trained language models, created by Tables 25-44.
Entity Type NER Eval. Toolkit Prec. Rec. F1
AGE BARTpho Mod. SLUE 0.70 0.79 0.74
BARTpho seqeval 0.69 0.61 0.65
BARTpho SLUE 0.69 0.78 0.73
mBART-50 Mod. SLUE 0.64 0.81 0.71
mBART-50 seqeval 0.66 0.53 0.59
mBART-50 SLUE 0.62 0.79 0.70
PhoBERT_base Mod. SLUE 0.75 0.81 0.78
PhoBERT_base seqeval 0.76 0.62 0.68
PhoBERT_base SLUE 0.74 0.80 0.77
PhoBERT_base-v2 Mod. SLUE 0.78 0.78 0.78
PhoBERT_base-v2 seqeval 0.76 0.66 0.71
PhoBERT_base-v2 SLUE 0.77 0.77 0.77
PhoBERT_large Mod. SLUE 0.77 0.81 0.79
PhoBERT_large seqeval 0.77 0.66 0.71
PhoBERT_large SLUE 0.76 0.80 0.78
ViDeBERTa_base Mod. SLUE 0.59 0.05 0.09
ViDeBERTa_base seqeval 0.03 0.29 0.05
ViDeBERTa_base SLUE 0.50 0.04 0.07
ViT5_base Mod. SLUE 0.71 0.79 0.75
ViT5_base seqeval 0.73 0.63 0.68
ViT5_base SLUE 0.69 0.77 0.73
XLM-R_base Mod. SLUE 0.69 0.79 0.73
XLM-R_base seqeval 0.71 0.56 0.63
XLM-R_base SLUE 0.68 0.77 0.72
XLM-R_large Mod. SLUE 0.78 0.77 0.78
XLM-R_large seqeval 0.75 0.67 0.71
XLM-R_large SLUE 0.78 0.77 0.77
Table 5: NER results of AGE entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
DATETIME BARTpho Mod. SLUE 0.76 0.76 0.76
BARTpho seqeval 0.62 0.66 0.64
BARTpho SLUE 0.74 0.75 0.75
mBART-50 Mod. SLUE 0.83 0.74 0.78
mBART-50 seqeval 0.67 0.75 0.71
mBART-50 SLUE 0.82 0.73 0.77
PhoBERT_base Mod. SLUE 0.80 0.83 0.82
PhoBERT_base seqeval 0.77 0.70 0.74
PhoBERT_base SLUE 0.80 0.82 0.81
PhoBERT_base-v2 Mod. SLUE 0.81 0.84 0.83
PhoBERT_base-v2 seqeval 0.78 0.73 0.75
PhoBERT_base-v2 SLUE 0.81 0.83 0.82
PhoBERT_large Mod. SLUE 0.83 0.82 0.83
PhoBERT_large seqeval 0.76 0.75 0.76
PhoBERT_large SLUE 0.83 0.81 0.82
ViDeBERTa_base Mod. SLUE 0.62 0.68 0.65
ViDeBERTa_base seqeval 0.58 0.43 0.50
ViDeBERTa_base SLUE 0.60 0.65 0.62
ViT5_base Mod. SLUE 0.74 0.77 0.75
ViT5_base seqeval 0.69 0.67 0.68
ViT5_base SLUE 0.72 0.75 0.74
XLM-R_base Mod. SLUE 0.81 0.78 0.80
XLM-R_base seqeval 0.72 0.70 0.71
XLM-R_base SLUE 0.80 0.77 0.78
XLM-R_large Mod. SLUE 0.85 0.82 0.83
XLM-R_large seqeval 0.76 0.77 0.77
XLM-R_large SLUE 0.84 0.81 0.82
Table 6: NER results of DATETIME entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
DIAGNOSTICS BARTpho Mod. SLUE 0.66 0.82 0.73
BARTpho seqeval 0.72 0.65 0.68
BARTpho SLUE 0.65 0.82 0.73
mBART-50 Mod. SLUE 0.58 0.81 0.68
mBART-50 seqeval 0.71 0.59 0.65
mBART-50 SLUE 0.57 0.80 0.66
PhoBERT_base Mod. SLUE 0.66 0.81 0.73
PhoBERT_base seqeval 0.75 0.64 0.69
PhoBERT_base SLUE 0.66 0.80 0.72
PhoBERT_base-v2 Mod. SLUE 0.66 0.82 0.73
PhoBERT_base-v2 seqeval 0.77 0.65 0.70
PhoBERT_base-v2 SLUE 0.66 0.82 0.73
PhoBERT_large Mod. SLUE 0.70 0.83 0.76
PhoBERT_large seqeval 0.78 0.69 0.73
PhoBERT_large SLUE 0.69 0.83 0.75
ViDeBERTa_base Mod. SLUE 0.53 0.60 0.56
ViDeBERTa_base seqeval 0.55 0.39 0.45
ViDeBERTa_base SLUE 0.51 0.58 0.54
ViT5_base Mod. SLUE 0.60 0.80 0.69
ViT5_base seqeval 0.71 0.62 0.67
ViT5_base SLUE 0.59 0.78 0.67
XLM-R_base Mod. SLUE 0.61 0.82 0.70
XLM-R_base seqeval 0.75 0.58 0.65
XLM-R_base SLUE 0.60 0.82 0.69
XLM-R_large Mod. SLUE 0.69 0.78 0.73
XLM-R_large seqeval 0.78 0.69 0.73
XLM-R_large SLUE 0.69 0.78 0.73
Table 7: NER results of DIAGNOSTICS entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
DISEASESYMTOM BARTpho Mod. SLUE 0.75 0.81 0.78
BARTpho seqeval 0.63 0.60 0.61
BARTpho SLUE 0.73 0.78 0.75
mBART-50 Mod. SLUE 0.71 0.73 0.72
mBART-50 seqeval 0.57 0.56 0.57
mBART-50 SLUE 0.68 0.69 0.69
PhoBERT_base Mod. SLUE 0.79 0.83 0.81
PhoBERT_base seqeval 0.70 0.60 0.65
PhoBERT_base SLUE 0.77 0.80 0.78
PhoBERT_base-v2 Mod. SLUE 0.79 0.85 0.82
PhoBERT_base-v2 seqeval 0.73 0.61 0.66
PhoBERT_base-v2 SLUE 0.77 0.82 0.79
PhoBERT_large Mod. SLUE 0.82 0.81 0.82
PhoBERT_large seqeval 0.71 0.64 0.67
PhoBERT_large SLUE 0.79 0.79 0.79
ViDeBERTa_base Mod. SLUE 0.68 0.52 0.59
ViDeBERTa_base seqeval 0.35 0.38 0.37
ViDeBERTa_base SLUE 0.62 0.47 0.54
ViT5_base Mod. SLUE 0.78 0.84 0.81
ViT5_base seqeval 0.73 0.67 0.70
ViT5_base SLUE 0.77 0.82 0.79
XLM-R_base Mod. SLUE 0.75 0.83 0.79
XLM-R_base seqeval 0.69 0.56 0.61
XLM-R_base SLUE 0.72 0.79 0.75
XLM-R_large Mod. SLUE 0.83 0.81 0.82
XLM-R_large seqeval 0.70 0.66 0.68
XLM-R_large SLUE 0.81 0.79 0.80
Table 8: NER results of DISEASESYMTOM entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
DRUGCHEMICAL BARTpho Mod. SLUE 0.73 0.82 0.77
BARTpho seqeval 0.75 0.70 0.73
BARTpho SLUE 0.73 0.81 0.77
mBART-50 Mod. SLUE 0.79 0.69 0.74
mBART-50 seqeval 0.71 0.76 0.73
mBART-50 SLUE 0.79 0.69 0.74
PhoBERT_base Mod. SLUE 0.80 0.93 0.86
PhoBERT_base seqeval 0.91 0.77 0.83
PhoBERT_base SLUE 0.79 0.93 0.86
PhoBERT_base-v2 Mod. SLUE 0.83 0.93 0.88
PhoBERT_base-v2 seqeval 0.91 0.80 0.85
PhoBERT_base-v2 SLUE 0.83 0.93 0.88
PhoBERT_large Mod. SLUE 0.74 0.93 0.82
PhoBERT_large seqeval 0.93 0.72 0.81
PhoBERT_large SLUE 0.74 0.93 0.82
ViDeBERTa_base Mod. SLUE 0.65 0.41 0.51
ViDeBERTa_base seqeval 0.31 0.41 0.35
ViDeBERTa_base SLUE 0.60 0.38 0.47
ViT5_base Mod. SLUE 0.75 0.87 0.80
ViT5_base seqeval 0.83 0.74 0.78
ViT5_base SLUE 0.75 0.86 0.80
XLM-R_base Mod. SLUE 0.81 0.74 0.77
XLM-R_base seqeval 0.76 0.76 0.76
XLM-R_base SLUE 0.80 0.74 0.77
XLM-R_large Mod. SLUE 0.85 0.92 0.88
XLM-R_large seqeval 0.91 0.79 0.85
XLM-R_large SLUE 0.85 0.92 0.88
Table 9: NER results of DRUGCHEMICAL entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
FOODDRINK BARTpho Mod. SLUE 0.67 0.83 0.74
BARTpho seqeval 0.69 0.59 0.64
BARTpho SLUE 0.67 0.83 0.74
mBART-50 Mod. SLUE 0.56 0.67 0.61
mBART-50 seqeval 0.57 0.52 0.54
mBART-50 SLUE 0.56 0.67 0.61
PhoBERT_base Mod. SLUE 0.78 0.84 0.81
PhoBERT_base seqeval 0.74 0.67 0.70
PhoBERT_base SLUE 0.78 0.84 0.81
PhoBERT_base-v2 Mod. SLUE 0.82 0.89 0.85
PhoBERT_base-v2 seqeval 0.83 0.76 0.79
PhoBERT_base-v2 SLUE 0.82 0.89 0.85
PhoBERT_large Mod. SLUE 0.80 0.91 0.85
PhoBERT_large seqeval 0.85 0.75 0.80
PhoBERT_large SLUE 0.80 0.91 0.85
ViDeBERTa_base Mod. SLUE 0.22 0.38 0.28
ViDeBERTa_base seqeval 0.24 0.12 0.16
ViDeBERTa_base SLUE 0.19 0.32 0.23
ViT5_base Mod. SLUE 0.70 0.86 0.77
ViT5_base seqeval 0.75 0.66 0.70
ViT5_base SLUE 0.70 0.86 0.77
XLM-R_base Mod. SLUE 0.64 0.88 0.74
XLM-R_base seqeval 0.82 0.58 0.68
XLM-R_base SLUE 0.62 0.85 0.72
XLM-R_large Mod. SLUE 0.88 0.81 0.84
XLM-R_large seqeval 0.75 0.83 0.79
XLM-R_large SLUE 0.88 0.81 0.84
Table 10: NER results of FOODDRINK entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
GENDER BARTpho Mod. SLUE 0.83 0.90 0.86
BARTpho seqeval 0.85 0.78 0.81
BARTpho SLUE 0.82 0.90 0.86
mBART-50 Mod. SLUE 0.83 0.87 0.85
mBART-50 seqeval 0.77 0.75 0.76
mBART-50 SLUE 0.82 0.86 0.84
PhoBERT_base Mod. SLUE 0.83 0.90 0.86
PhoBERT_base seqeval 0.76 0.74 0.75
PhoBERT_base SLUE 0.83 0.90 0.86
PhoBERT_base-v2 Mod. SLUE 0.84 0.89 0.86
PhoBERT_base-v2 seqeval 0.83 0.77 0.80
PhoBERT_base-v2 SLUE 0.84 0.89 0.86
PhoBERT_large Mod. SLUE 0.83 0.90 0.87
PhoBERT_large seqeval 0.87 0.79 0.83
PhoBERT_large SLUE 0.83 0.90 0.86
ViDeBERTa_base Mod. SLUE 0.00 0.00 0.00
ViDeBERTa_base seqeval 0.00 0.00 0.00
ViDeBERTa_base SLUE 0.00 0.00 0.00
ViT5_base Mod. SLUE 0.81 0.82 0.82
ViT5_base seqeval 0.67 0.71 0.69
ViT5_base SLUE 0.81 0.82 0.81
XLM-R_base Mod. SLUE 0.79 0.72 0.75
XLM-R_base seqeval 0.56 0.49 0.52
XLM-R_base SLUE 0.78 0.71 0.75
XLM-R_large Mod. SLUE 0.91 0.90 0.90
XLM-R_large seqeval 0.82 0.79 0.80
XLM-R_large SLUE 0.91 0.89 0.90
Table 11: NER results of GENDER entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
LOCATION BARTpho Mod. SLUE 0.77 0.88 0.82
BARTpho seqeval 0.75 0.68 0.71
BARTpho SLUE 0.76 0.87 0.81
mBART-50 Mod. SLUE 0.74 0.83 0.78
mBART-50 seqeval 0.73 0.68 0.71
mBART-50 SLUE 0.74 0.82 0.78
PhoBERT_base Mod. SLUE 0.75 0.92 0.82
PhoBERT_base seqeval 0.81 0.63 0.71
PhoBERT_base SLUE 0.74 0.91 0.81
PhoBERT_base-v2 Mod. SLUE 0.78 0.95 0.86
PhoBERT_base-v2 seqeval 0.86 0.69 0.77
PhoBERT_base-v2 SLUE 0.78 0.94 0.86
PhoBERT_large Mod. SLUE 0.80 0.91 0.85
PhoBERT_large seqeval 0.82 0.69 0.75
PhoBERT_large SLUE 0.80 0.90 0.84
ViDeBERTa_base Mod. SLUE 0.78 0.33 0.47
ViDeBERTa_base seqeval 0.32 0.60 0.42
ViDeBERTa_base SLUE 0.77 0.33 0.46
ViT5_base Mod. SLUE 0.74 0.92 0.82
ViT5_base seqeval 0.83 0.65 0.73
ViT5_base SLUE 0.73 0.92 0.81
XLM-R_base Mod. SLUE 0.75 0.84 0.79
XLM-R_base seqeval 0.75 0.59 0.66
XLM-R_base SLUE 0.74 0.82 0.78
XLM-R_large Mod. SLUE 0.85 0.93 0.89
XLM-R_large seqeval 0.87 0.75 0.81
XLM-R_large SLUE 0.85 0.93 0.88
Table 12: NER results of LOCATION entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
LOCATION BARTpho Mod. SLUE 0.54 0.21 0.30
BARTpho seqeval 0.16 0.39 0.23
BARTpho SLUE 0.48 0.18 0.26
mBART-50 Mod. SLUE 0.45 0.09 0.15
mBART-50 seqeval 0.08 0.35 0.13
mBART-50 SLUE 0.36 0.07 0.12
PhoBERT_base Mod. SLUE 0.55 0.38 0.45
PhoBERT_base seqeval 0.24 0.27 0.26
PhoBERT_base SLUE 0.49 0.34 0.40
PhoBERT_base-v2 Mod. SLUE 0.59 0.38 0.46
PhoBERT_base-v2 seqeval 0.27 0.35 0.31
PhoBERT_base-v2 SLUE 0.53 0.34 0.41
PhoBERT_large Mod. SLUE 0.61 0.30 0.40
PhoBERT_large seqeval 0.22 0.35 0.27
PhoBERT_large SLUE 0.55 0.27 0.36
ViDeBERTa_base Mod. SLUE 0.34 0.14 0.20
ViDeBERTa_base seqeval 0.06 0.09 0.07
ViDeBERTa_base SLUE 0.24 0.10 0.14
ViT5_base Mod. SLUE 0.52 0.21 0.30
ViT5_base seqeval 0.16 0.37 0.23
ViT5_base SLUE 0.47 0.19 0.27
XLM-R_base Mod. SLUE 0.48 0.27 0.34
XLM-R_base seqeval 0.12 0.18 0.14
XLM-R_base SLUE 0.41 0.23 0.29
XLM-R_large Mod. SLUE 0.58 0.28 0.38
XLM-R_large seqeval 0.20 0.32 0.25
XLM-R_large SLUE 0.51 0.25 0.33
Table 13: NER results of MEDDEVICETECHNIQUE entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
OCCUPATION BARTpho Mod. SLUE 0.91 0.93 0.92
BARTpho seqeval 0.87 0.87 0.87
BARTpho SLUE 0.91 0.92 0.92
mBART-50 Mod. SLUE 0.97 0.93 0.95
mBART-50 seqeval 0.91 0.95 0.93
mBART-50 SLUE 0.97 0.93 0.95
PhoBERT_base Mod. SLUE 0.95 0.96 0.95
PhoBERT_base seqeval 0.94 0.92 0.93
PhoBERT_base SLUE 0.95 0.96 0.95
PhoBERT_base-v2 Mod. SLUE 0.96 0.96 0.96
PhoBERT_base-v2 seqeval 0.95 0.93 0.94
PhoBERT_base-v2 SLUE 0.96 0.96 0.96
PhoBERT_large Mod. SLUE 0.97 0.95 0.96
PhoBERT_large seqeval 0.93 0.94 0.94
PhoBERT_large SLUE 0.97 0.95 0.96
ViDeBERTa_base Mod. SLUE 0.96 0.82 0.89
ViDeBERTa_base seqeval 0.81 0.91 0.86
ViDeBERTa_base SLUE 0.96 0.81 0.88
ViT5_base Mod. SLUE 0.95 0.93 0.94
ViT5_base seqeval 0.92 0.94 0.93
ViT5_base SLUE 0.95 0.93 0.94
XLM-R_base Mod. SLUE 0.88 0.96 0.92
XLM-R_base seqeval 0.93 0.83 0.88
XLM-R_base SLUE 0.88 0.96 0.92
XLM-R_large Mod. SLUE 0.97 0.97 0.97
XLM-R_large seqeval 0.95 0.95 0.95
XLM-R_large SLUE 0.97 0.97 0.97
Table 14: NER results of OCCUPATION entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
ORGAN BARTpho Mod. SLUE 0.72 0.80 0.76
BARTpho seqeval 0.63 0.56 0.60
BARTpho SLUE 0.70 0.77 0.73
mBART-50 Mod. SLUE 0.67 0.74 0.70
mBART-50 seqeval 0.61 0.51 0.56
mBART-50 SLUE 0.64 0.71 0.67
PhoBERT_base Mod. SLUE 0.74 0.86 0.79
PhoBERT_base seqeval 0.69 0.55 0.61
PhoBERT_base SLUE 0.71 0.83 0.77
PhoBERT_base-v2 Mod. SLUE 0.74 0.88 0.80
PhoBERT_base-v2 seqeval 0.70 0.55 0.62
PhoBERT_base-v2 SLUE 0.72 0.86 0.78
PhoBERT_large Mod. SLUE 0.73 0.87 0.80
PhoBERT_large seqeval 0.71 0.55 0.62
PhoBERT_large SLUE 0.71 0.85 0.77
ViDeBERTa_base Mod. SLUE 0.53 0.51 0.52
ViDeBERTa_base seqeval 0.32 0.27 0.30
ViDeBERTa_base SLUE 0.45 0.44 0.44
ViT5_base Mod. SLUE 0.71 0.85 0.78
ViT5_base seqeval 0.70 0.58 0.64
ViT5_base SLUE 0.70 0.84 0.76
XLM-R_base Mod. SLUE 0.71 0.87 0.78
XLM-R_base seqeval 0.70 0.55 0.61
XLM-R_base SLUE 0.69 0.85 0.76
XLM-R_large Mod. SLUE 0.77 0.87 0.82
XLM-R_large seqeval 0.75 0.60 0.66
XLM-R_large SLUE 0.75 0.85 0.80
Table 15: NER results of ORGAN entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
ORGANIZATION BARTpho Mod. SLUE 0.73 0.54 0.62
BARTpho seqeval 0.47 0.68 0.56
BARTpho SLUE 0.71 0.53 0.61
mBART-50 Mod. SLUE 0.77 0.31 0.44
mBART-50 seqeval 0.28 0.70 0.40
mBART-50 SLUE 0.77 0.31 0.44
PhoBERT_base Mod. SLUE 0.71 0.57 0.63
PhoBERT_base seqeval 0.51 0.56 0.53
PhoBERT_base SLUE 0.70 0.56 0.62
PhoBERT_base-v2 Mod. SLUE 0.76 0.71 0.73
PhoBERT_base-v2 seqeval 0.59 0.60 0.60
PhoBERT_base-v2 SLUE 0.74 0.70 0.72
PhoBERT_large Mod. SLUE 0.71 0.65 0.68
PhoBERT_large seqeval 0.59 0.50 0.54
PhoBERT_large SLUE 0.71 0.65 0.68
ViDeBERTa_base Mod. SLUE 0.00 0.00 0.00
ViDeBERTa_base seqeval 0.00 0.00 0.00
ViDeBERTa_base SLUE 0.00 0.00 0.00
ViT5_base Mod. SLUE 0.94 0.36 0.52
ViT5_base seqeval 0.34 0.91 0.50
ViT5_base SLUE 0.94 0.36 0.52
XLM-R_base Mod. SLUE 0.59 0.31 0.41
XLM-R_base seqeval 0.08 0.12 0.10
XLM-R_base SLUE 0.56 0.29 0.39
XLM-R_large Mod. SLUE 0.82 0.66 0.73
XLM-R_large seqeval 0.61 0.70 0.65
XLM-R_large SLUE 0.80 0.64 0.71
Table 16: NER results of ORGANIZATION entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
PERSONALCARE BARTpho Mod. SLUE 0.17 0.77 0.28
BARTpho seqeval 0.75 0.21 0.32
BARTpho SLUE 0.17 0.77 0.28
mBART-50 Mod. SLUE 0.23 0.80 0.36
mBART-50 seqeval 0.76 0.26 0.39
mBART-50 SLUE 0.23 0.80 0.36
PhoBERT_base Mod. SLUE 0.18 0.87 0.30
PhoBERT_base seqeval 0.85 0.20 0.33
PhoBERT_base SLUE 0.18 0.85 0.30
PhoBERT_base-v2 Mod. SLUE 0.16 0.85 0.27
PhoBERT_base-v2 seqeval 0.85 0.18 0.30
PhoBERT_base-v2 SLUE 0.16 0.84 0.27
PhoBERT_large Mod. SLUE 0.19 0.83 0.31
PhoBERT_large seqeval 0.80 0.21 0.33
PhoBERT_large SLUE 0.19 0.82 0.31
ViDeBERTa_base Mod. SLUE 0.17 0.76 0.28
ViDeBERTa_base seqeval 0.68 0.14 0.23
ViDeBERTa_base SLUE 0.17 0.75 0.28
ViT5_base Mod. SLUE 0.17 0.79 0.28
ViT5_base seqeval 0.75 0.18 0.29
ViT5_base SLUE 0.17 0.78 0.27
XLM-R_base Mod. SLUE 0.21 0.83 0.33
XLM-R_base seqeval 0.77 0.21 0.33
XLM-R_base SLUE 0.20 0.81 0.32
XLM-R_large Mod. SLUE 0.22 0.87 0.36
XLM-R_large seqeval 0.84 0.24 0.38
XLM-R_large SLUE 0.22 0.86 0.35
Table 17: NER results of PERSONALCARE entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
PREVENTIVEMED BARTpho Mod. SLUE 0.01 0.15 0.01
BARTpho seqeval 0.21 0.01 0.02
BARTpho SLUE 0.01 0.15 0.01
mBART-50 Mod. SLUE 0.02 0.45 0.04
mBART-50 seqeval 0.42 0.02 0.04
mBART-50 SLUE 0.02 0.45 0.04
PhoBERT_base Mod. SLUE 0.03 0.75 0.06
PhoBERT_base seqeval 0.56 0.02 0.05
PhoBERT_base SLUE 0.03 0.69 0.06
PhoBERT_base-v2 Mod. SLUE 0.02 0.44 0.04
PhoBERT_base-v2 seqeval 0.33 0.02 0.03
PhoBERT_base-v2 SLUE 0.02 0.39 0.03
PhoBERT_large Mod. SLUE 0.03 0.78 0.06
PhoBERT_large seqeval 0.72 0.03 0.06
PhoBERT_large SLUE 0.03 0.75 0.06
ViDeBERTa_base Mod. SLUE 0.00 0.06 0.01
ViDeBERTa_base seqeval 0.00 0.00 0.00
ViDeBERTa_base SLUE 0.00 0.00 0.00
ViT5_base Mod. SLUE 0.02 0.39 0.04
ViT5_base seqeval 0.39 0.03 0.05
ViT5_base SLUE 0.02 0.35 0.04
XLM-R_base Mod. SLUE 0.01 0.14 0.02
XLM-R_base seqeval 0.00 0.00 0.00
XLM-R_base SLUE 0.01 0.08 0.01
XLM-R_large Mod. SLUE 0.03 0.78 0.06
XLM-R_large seqeval 0.72 0.03 0.05
XLM-R_large SLUE 0.03 0.78 0.06
Table 18: NER results of PREVENTIVEMED entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
SURGERY BARTpho Mod. SLUE 0.57 0.57 0.57
BARTpho seqeval 0.48 0.55 0.51
BARTpho SLUE 0.57 0.56 0.57
mBART-50 Mod. SLUE 0.67 0.38 0.48
mBART-50 seqeval 0.29 0.55 0.38
mBART-50 SLUE 0.64 0.37 0.47
PhoBERT_base Mod. SLUE 0.70 0.48 0.56
PhoBERT_base seqeval 0.36 0.50 0.42
PhoBERT_base SLUE 0.68 0.46 0.55
PhoBERT_base-v2 Mod. SLUE 0.79 0.66 0.72
PhoBERT_base-v2 seqeval 0.62 0.59 0.61
PhoBERT_base-v2 SLUE 0.78 0.65 0.71
PhoBERT_large Mod. SLUE 0.85 0.46 0.59
PhoBERT_large seqeval 0.38 0.66 0.49
PhoBERT_large SLUE 0.85 0.45 0.59
ViDeBERTa_base Mod. SLUE 0.78 0.32 0.46
ViDeBERTa_base seqeval 0.24 0.45 0.32
ViDeBERTa_base SLUE 0.76 0.31 0.44
ViT5_base Mod. SLUE 0.61 0.69 0.65
ViT5_base seqeval 0.55 0.51 0.52
ViT5_base SLUE 0.60 0.69 0.64
XLM-R_base Mod. SLUE 0.66 0.67 0.66
XLM-R_base seqeval 0.59 0.46 0.52
XLM-R_base SLUE 0.66 0.66 0.66
XLM-R_large Mod. SLUE 0.80 0.46 0.59
XLM-R_large seqeval 0.39 0.65 0.49
XLM-R_large SLUE 0.80 0.46 0.58
Table 19: NER results of SURGERY entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
TRANSPORTATION BARTpho Mod. SLUE 0.98 0.96 0.97
BARTpho seqeval 0.93 0.93 0.93
BARTpho SLUE 0.98 0.96 0.97
mBART-50 Mod. SLUE 1,00 0.70 0.82
mBART-50 seqeval 0.67 0.95 0.78
mBART-50 SLUE 1,00 0.70 0.82
PhoBERT_base Mod. SLUE 1,00 0.90 0.95
PhoBERT_base seqeval 0.85 0.92 0.88
PhoBERT_base SLUE 1,00 0.90 0.95
PhoBERT_base-v2 Mod. SLUE 1,00 0.69 0.82
PhoBERT_base-v2 seqeval 0.59 0.80 0.68
PhoBERT_base-v2 SLUE 1,00 0.69 0.82
PhoBERT_large Mod. SLUE 0.95 0.91 0.93
PhoBERT_large seqeval 0.89 0.92 0.91
PhoBERT_large SLUE 0.95 0.91 0.93
ViDeBERTa_base Mod. SLUE 0.00 0.00 0.00
ViDeBERTa_base seqeval 0.00 0.00 0.00
ViDeBERTa_base SLUE 0.00 0.00 0.00
ViT5_base Mod. SLUE 0.00 0.00 0.00
ViT5_base seqeval 0.00 0.00 0.00
ViT5_base SLUE 0.00 0.00 0.00
XLM-R_base Mod. SLUE 0.00 0.00 0.00
XLM-R_base seqeval 0.00 0.00 0.00
XLM-R_base SLUE 0.00 0.00 0.00
XLM-R_large Mod. SLUE 1,00 0.72 0.84
XLM-R_large seqeval 0.63 0.81 0.71
XLM-R_large SLUE 1,00 0.72 0.84
Table 20: NER results of TRANSPORTATION entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
TREATMENT BARTpho Mod. SLUE 0.63 0.82 0.71
BARTpho seqeval 0.75 0.57 0.65
BARTpho SLUE 0.63 0.82 0.71
mBART-50 Mod. SLUE 0.66 0.83 0.74
mBART-50 seqeval 0.81 0.61 0.70
mBART-50 SLUE 0.66 0.83 0.74
PhoBERT_base Mod. SLUE 0.66 0.88 0.75
PhoBERT_base seqeval 0.86 0.60 0.71
PhoBERT_base SLUE 0.66 0.88 0.75
PhoBERT_base-v2 Mod. SLUE 0.62 0.88 0.73
PhoBERT_base-v2 seqeval 0.87 0.56 0.68
PhoBERT_base-v2 SLUE 0.62 0.88 0.72
PhoBERT_large Mod. SLUE 0.59 0.87 0.70
PhoBERT_large seqeval 0.86 0.53 0.65
PhoBERT_large SLUE 0.59 0.87 0.70
ViDeBERTa_base Mod. SLUE 0.62 0.86 0.72
ViDeBERTa_base seqeval 0.84 0.52 0.64
ViDeBERTa_base SLUE 0.62 0.86 0.72
ViT5_base Mod. SLUE 0.46 0.84 0.60
ViT5_base seqeval 0.81 0.41 0.55
ViT5_base SLUE 0.46 0.84 0.60
XLM-R_base Mod. SLUE 0.48 0.86 0.62
XLM-R_base seqeval 0.84 0.40 0.54
XLM-R_base SLUE 0.48 0.86 0.62
XLM-R_large Mod. SLUE 0.59 0.89 0.71
XLM-R_large seqeval 0.89 0.54 0.67
XLM-R_large SLUE 0.59 0.89 0.71
Table 21: NER results of TREATMENT entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
UNITCALIBRATOR BARTpho Mod. SLUE 0.42 0.65 0.51
BARTpho seqeval 0.52 0.31 0.39
BARTpho SLUE 0.41 0.63 0.50
mBART-50 Mod. SLUE 0.38 0.58 0.46
mBART-50 seqeval 0.45 0.26 0.33
mBART-50 SLUE 0.37 0.56 0.44
PhoBERT_base Mod. SLUE 0.44 0.73 0.55
PhoBERT_base seqeval 0.61 0.30 0.41
PhoBERT_base SLUE 0.43 0.71 0.53
PhoBERT_base-v2 Mod. SLUE 0.46 0.74 0.57
PhoBERT_base-v2 seqeval 0.61 0.33 0.43
PhoBERT_base-v2 SLUE 0.45 0.72 0.55
PhoBERT_large Mod. SLUE 0.48 0.73 0.58
PhoBERT_large seqeval 0.60 0.34 0.43
PhoBERT_large SLUE 0.47 0.71 0.56
ViDeBERTa_base Mod. SLUE 0.32 0.44 0.37
ViDeBERTa_base seqeval 0.20 0.10 0.14
ViDeBERTa_base SLUE 0.28 0.39 0.33
ViT5_base Mod. SLUE 0.34 0.63 0.44
ViT5_base seqeval 0.50 0.24 0.32
ViT5_base SLUE 0.33 0.62 0.43
XLM-R_base Mod. SLUE 0.44 0.72 0.54
XLM-R_base seqeval 0.54 0.27 0.36
XLM-R_base SLUE 0.42 0.69 0.52
XLM-R_large Mod. SLUE 0.50 0.75 0.60
XLM-R_large seqeval 0.65 0.37 0.47
XLM-R_large SLUE 0.49 0.74 0.59
Table 22: NER results of UNITCALIBRATOR entity type (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
Overall Macro BARTpho Mod. SLUE 0.64 0.72 0.66
BARTpho seqeval 0.64 0.58 0.59
BARTpho SLUE 0.63 0.71 0.65
mBART-50 Mod. SLUE 0.64 0.66 0.61
mBART-50 seqeval 0.59 0.57 0.55
mBART-50 SLUE 0.63 0.65 0.60
PhoBERT_base Mod. SLUE 0.67 0.79 0.69
PhoBERT_base seqeval 0.70 0.57 0.60
PhoBERT_base SLUE 0.66 0.78 0.68
PhoBERT_base-v2 Mod. SLUE 0.69 0.79 0.71
PhoBERT_base-v2 seqeval 0.71 0.59 0.62
PhoBERT_base-v2 SLUE 0.68 0.77 0.70
PhoBERT_large Mod. SLUE 0.69 0.79 0.70
PhoBERT_large seqeval 0.73 0.60 0.63
PhoBERT_large SLUE 0.68 0.78 0.69
ViDeBERTa_base Mod. SLUE 0.43 0.38 0.37
ViDeBERTa_base seqeval 0.31 0.28 0.27
ViDeBERTa_base SLUE 0.40 0.36 0.34
ViT5_base Mod. SLUE 0.59 0.69 0.60
ViT5_base seqeval 0.61 0.53 0.54
ViT5_base SLUE 0.58 0.68 0.59
XLM-R_base Mod. SLUE 0.57 0.67 0.59
XLM-R_base seqeval 0.57 0.44 0.48
XLM-R_base SLUE 0.56 0.65 0.58
XLM-R_large Mod. SLUE 0.72 0.78 0.71
XLM-R_large seqeval 0.72 0.62 0.63
XLM-R_large SLUE 0.71 0.77 0.70
Table 23: NER results of Overall Macro (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER Eval. Toolkit Prec. Rec. F1
Overall Micro BARTpho Mod. SLUE 0.66 0.74 0.70
BARTpho seqeval 0.64 0.58 0.61
BARTpho SLUE 0.64 0.73 0.68
mBART-50 Mod. SLUE 0.65 0.68 0.67
mBART-50 seqeval 0.60 0.57 0.59
mBART-50 SLUE 0.64 0.66 0.65
PhoBERT_base Mod. SLUE 0.69 0.79 0.74
PhoBERT_base seqeval 0.71 0.57 0.63
PhoBERT_base SLUE 0.67 0.78 0.72
PhoBERT_base-v2 Mod. SLUE 0.70 0.81 0.75
PhoBERT_base-v2 seqeval 0.74 0.59 0.66
PhoBERT_base-v2 SLUE 0.68 0.79 0.74
PhoBERT_large Mod. SLUE 0.70 0.79 0.74
PhoBERT_large seqeval 0.73 0.60 0.66
PhoBERT_large SLUE 0.69 0.77 0.73
ViDeBERTa_base Mod. SLUE 0.55 0.45 0.49
ViDeBERTa_base seqeval 0.33 0.34 0.34
ViDeBERTa_base SLUE 0.50 0.41 0.45
ViT5_base Mod. SLUE 0.65 0.76 0.70
ViT5_base seqeval 0.68 0.59 0.63
ViT5_base SLUE 0.64 0.74 0.69
XLM-R_base Mod. SLUE 0.66 0.75 0.70
XLM-R_base seqeval 0.66 0.53 0.59
XLM-R_base SLUE 0.64 0.73 0.69
XLM-R_large Mod. SLUE 0.72 0.78 0.75
XLM-R_large seqeval 0.72 0.63 0.67
XLM-R_large SLUE 0.71 0.77 0.74
Table 24: NER results of Overall Micro (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
AGE ViDeBERTa_base XLSR-53-Viet SLUE 0.24 0.03 0.05
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.32 0.04 0.07
ViDeBERTa_base w2v2-Viet SLUE 0.24 0.03 0.05
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.33 0.04 0.07
ViT5_base XLSR-53-Viet SLUE 0.63 0.49 0.55
ViT5_base XLSR-53-Viet Mod. SLUE 0.68 0.53 0.59
ViT5_base w2v2-Viet SLUE 0.63 0.52 0.57
ViT5_base w2v2-Viet Mod. SLUE 0.68 0.56 0.61
mBART-50 XLSR-53-Viet SLUE 0.44 0.06 0.11
mBART-50 XLSR-53-Viet Mod. SLUE 0.55 0.08 0.14
mBART-50 w2v2-Viet SLUE 0.44 0.08 0.13
mBART-50 w2v2-Viet Mod. SLUE 0.58 0.10 0.17
BARTpho XLSR-53-Viet SLUE 0.62 0.57 0.59
BARTpho XLSR-53-Viet Mod. SLUE 0.68 0.62 0.65
BARTpho w2v2-Viet SLUE 0.59 0.58 0.58
BARTpho w2v2-Viet Mod. SLUE 0.64 0.63 0.63
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.72 0.58 0.64
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.75 0.61 0.67
PhoBERT_base-v2 w2v2-Viet SLUE 0.70 0.58 0.64
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.74 0.61 0.67
PhoBERT_base XLSR-53-Viet SLUE 0.69 0.60 0.64
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.74 0.64 0.68
PhoBERT_base w2v2-Viet SLUE 0.67 0.61 0.64
PhoBERT_base w2v2-Viet Mod. SLUE 0.70 0.64 0.67
PhoBERT_large XLSR-53-Viet SLUE 0.71 0.59 0.65
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.75 0.62 0.68
PhoBERT_large w2v2-Viet SLUE 0.69 0.60 0.64
PhoBERT_large w2v2-Viet Mod. SLUE 0.72 0.63 0.67
XLM-R_base XLSR-53-Viet SLUE 0.61 0.58 0.60
XLM-R_base XLSR-53-Viet Mod. SLUE 0.65 0.62 0.63
XLM-R_base w2v2-Viet SLUE 0.59 0.59 0.59
XLM-R_base w2v2-Viet Mod. SLUE 0.63 0.63 0.63
XLM-R_large XLSR-53-Viet SLUE 0.72 0.61 0.66
XLM-R_large XLSR-53-Viet Mod. SLUE 0.76 0.64 0.69
XLM-R_large w2v2-Viet SLUE 0.70 0.62 0.66
XLM-R_large w2v2-Viet Mod. SLUE 0.73 0.65 0.69
Table 25: NER results of AGE entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
DATETIME ViDeBERTa_base XLSR-53-Viet SLUE 0.52 0.56 0.54
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.56 0.60 0.58
ViDeBERTa_base w2v2-Viet SLUE 0.53 0.57 0.55
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.57 0.60 0.59
ViT5_base XLSR-53-Viet SLUE 0.57 0.57 0.57
ViT5_base XLSR-53-Viet Mod. SLUE 0.63 0.62 0.62
ViT5_base w2v2-Viet SLUE 0.60 0.56 0.58
ViT5_base w2v2-Viet Mod. SLUE 0.63 0.59 0.61
mBART-50 XLSR-53-Viet SLUE 0.42 0.06 0.11
mBART-50 XLSR-53-Viet Mod. SLUE 0.57 0.09 0.15
mBART-50 w2v2-Viet SLUE 0.40 0.06 0.10
mBART-50 w2v2-Viet Mod. SLUE 0.61 0.09 0.15
BARTpho XLSR-53-Viet SLUE 0.64 0.62 0.63
BARTpho XLSR-53-Viet Mod. SLUE 0.68 0.66 0.67
BARTpho w2v2-Viet SLUE 0.65 0.60 0.62
BARTpho w2v2-Viet Mod. SLUE 0.69 0.64 0.66
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.67 0.69 0.68
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.70 0.72 0.71
PhoBERT_base-v2 w2v2-Viet SLUE 0.69 0.69 0.69
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.71 0.71 0.71
PhoBERT_base XLSR-53-Viet SLUE 0.66 0.68 0.67
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.69 0.72 0.70
PhoBERT_base w2v2-Viet SLUE 0.68 0.67 0.68
PhoBERT_base w2v2-Viet Mod. SLUE 0.71 0.70 0.70
PhoBERT_large XLSR-53-Viet SLUE 0.67 0.68 0.68
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.70 0.71 0.71
PhoBERT_large w2v2-Viet SLUE 0.70 0.66 0.68
PhoBERT_large w2v2-Viet Mod. SLUE 0.72 0.69 0.70
XLM-R_base XLSR-53-Viet SLUE 0.64 0.65 0.64
XLM-R_base XLSR-53-Viet Mod. SLUE 0.67 0.68 0.68
XLM-R_base w2v2-Viet SLUE 0.66 0.64 0.65
XLM-R_base w2v2-Viet Mod. SLUE 0.69 0.66 0.67
XLM-R_large XLSR-53-Viet SLUE 0.69 0.67 0.68
XLM-R_large XLSR-53-Viet Mod. SLUE 0.72 0.70 0.71
XLM-R_large w2v2-Viet SLUE 0.72 0.66 0.69
XLM-R_large w2v2-Viet Mod. SLUE 0.74 0.69 0.71
Table 26: NER results of DATETIME entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
DIAGNOSTICS ViDeBERTa_base XLSR-53-Viet SLUE 0.53 0.47 0.50
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.55 0.49 0.52
ViDeBERTa_base w2v2-Viet SLUE 0.51 0.44 0.47
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.53 0.46 0.49
ViT5_base XLSR-53-Viet SLUE 0.52 0.47 0.50
ViT5_base XLSR-53-Viet Mod. SLUE 0.54 0.49 0.51
ViT5_base w2v2-Viet SLUE 0.53 0.46 0.49
ViT5_base w2v2-Viet Mod. SLUE 0.56 0.49 0.52
mBART-50 XLSR-53-Viet SLUE 0.30 0.08 0.13
mBART-50 XLSR-53-Viet Mod. SLUE 0.34 0.10 0.15
mBART-50 w2v2-Viet SLUE 0.41 0.06 0.11
mBART-50 w2v2-Viet Mod. SLUE 0.48 0.08 0.13
BARTpho XLSR-53-Viet SLUE 0.59 0.56 0.58
BARTpho XLSR-53-Viet Mod. SLUE 0.62 0.59 0.60
BARTpho w2v2-Viet SLUE 0.60 0.53 0.56
BARTpho w2v2-Viet Mod. SLUE 0.63 0.56 0.59
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.56 0.57 0.57
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.58 0.59 0.58
PhoBERT_base-v2 w2v2-Viet SLUE 0.56 0.56 0.56
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.59 0.59 0.59
PhoBERT_base XLSR-53-Viet SLUE 0.57 0.57 0.57
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.59 0.59 0.59
PhoBERT_base w2v2-Viet SLUE 0.55 0.56 0.55
PhoBERT_base w2v2-Viet Mod. SLUE 0.57 0.58 0.57
PhoBERT_large XLSR-53-Viet SLUE 0.61 0.58 0.59
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.62 0.59 0.61
PhoBERT_large w2v2-Viet SLUE 0.58 0.57 0.57
PhoBERT_large w2v2-Viet Mod. SLUE 0.61 0.59 0.60
XLM-R_base XLSR-53-Viet SLUE 0.51 0.58 0.54
XLM-R_base XLSR-53-Viet Mod. SLUE 0.52 0.59 0.55
XLM-R_base w2v2-Viet SLUE 0.50 0.57 0.53
XLM-R_base w2v2-Viet Mod. SLUE 0.52 0.59 0.55
XLM-R_large XLSR-53-Viet SLUE 0.59 0.58 0.59
XLM-R_large XLSR-53-Viet Mod. SLUE 0.61 0.59 0.60
XLM-R_large w2v2-Viet SLUE 0.58 0.57 0.57
XLM-R_large w2v2-Viet Mod. SLUE 0.60 0.59 0.60
Table 27: NER results of DIAGNOSTICS entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
DISEASESYMTOM ViDeBERTa_base XLSR-53-Viet SLUE 0.53 0.39 0.45
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.65 0.48 0.55
ViDeBERTa_base w2v2-Viet SLUE 0.53 0.39 0.45
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.64 0.48 0.55
ViT5_base XLSR-53-Viet SLUE 0.59 0.49 0.54
ViT5_base XLSR-53-Viet Mod. SLUE 0.69 0.58 0.63
ViT5_base w2v2-Viet SLUE 0.59 0.49 0.54
ViT5_base w2v2-Viet Mod. SLUE 0.70 0.58 0.64
mBART-50 XLSR-53-Viet SLUE 0.51 0.09 0.15
mBART-50 XLSR-53-Viet Mod. SLUE 0.64 0.11 0.19
mBART-50 w2v2-Viet SLUE 0.48 0.07 0.12
mBART-50 w2v2-Viet Mod. SLUE 0.65 0.09 0.16
BARTpho XLSR-53-Viet SLUE 0.61 0.53 0.57
BARTpho XLSR-53-Viet Mod. SLUE 0.71 0.62 0.66
BARTpho w2v2-Viet SLUE 0.58 0.53 0.55
BARTpho w2v2-Viet Mod. SLUE 0.69 0.62 0.65
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.63 0.57 0.60
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.71 0.65 0.68
PhoBERT_base-v2 w2v2-Viet SLUE 0.61 0.57 0.59
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.71 0.66 0.68
PhoBERT_base XLSR-53-Viet SLUE 0.62 0.56 0.59
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.71 0.64 0.67
PhoBERT_base w2v2-Viet SLUE 0.61 0.55 0.58
PhoBERT_base w2v2-Viet Mod. SLUE 0.71 0.64 0.67
PhoBERT_large XLSR-53-Viet SLUE 0.63 0.56 0.59
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.71 0.63 0.67
PhoBERT_large w2v2-Viet SLUE 0.62 0.54 0.58
PhoBERT_large w2v2-Viet Mod. SLUE 0.71 0.62 0.66
XLM-R_base XLSR-53-Viet SLUE 0.58 0.55 0.57
XLM-R_base XLSR-53-Viet Mod. SLUE 0.67 0.64 0.65
XLM-R_base w2v2-Viet SLUE 0.57 0.54 0.56
XLM-R_base w2v2-Viet Mod. SLUE 0.67 0.64 0.65
XLM-R_large XLSR-53-Viet SLUE 0.64 0.57 0.60
XLM-R_large XLSR-53-Viet Mod. SLUE 0.72 0.64 0.68
XLM-R_large w2v2-Viet SLUE 0.64 0.56 0.60
XLM-R_large w2v2-Viet Mod. SLUE 0.72 0.63 0.67
Table 28: NER results of DISEASESYMTOM entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
DRUGCHEMICAL ViDeBERTa_base XLSR-53-Viet SLUE 0.51 0.34 0.41
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.59 0.40 0.47
ViDeBERTa_base w2v2-Viet SLUE 0.49 0.34 0.40
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.56 0.39 0.46
ViT5_base XLSR-53-Viet SLUE 0.58 0.52 0.55
ViT5_base XLSR-53-Viet Mod. SLUE 0.66 0.59 0.62
ViT5_base w2v2-Viet SLUE 0.60 0.51 0.55
ViT5_base w2v2-Viet Mod. SLUE 0.68 0.58 0.62
mBART-50 XLSR-53-Viet SLUE 0.32 0.06 0.11
mBART-50 XLSR-53-Viet Mod. SLUE 0.46 0.09 0.15
mBART-50 w2v2-Viet SLUE 0.35 0.07 0.11
mBART-50 w2v2-Viet Mod. SLUE 0.54 0.10 0.17
BARTpho XLSR-53-Viet SLUE 0.60 0.55 0.57
BARTpho XLSR-53-Viet Mod. SLUE 0.67 0.62 0.64
BARTpho w2v2-Viet SLUE 0.60 0.55 0.57
BARTpho w2v2-Viet Mod. SLUE 0.66 0.61 0.64
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.67 0.66 0.67
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.73 0.72 0.72
PhoBERT_base-v2 w2v2-Viet SLUE 0.66 0.66 0.66
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.72 0.72 0.72
PhoBERT_base XLSR-53-Viet SLUE 0.66 0.66 0.66
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.72 0.72 0.72
PhoBERT_base w2v2-Viet SLUE 0.64 0.66 0.65
PhoBERT_base w2v2-Viet Mod. SLUE 0.71 0.72 0.72
PhoBERT_large XLSR-53-Viet SLUE 0.63 0.67 0.65
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.69 0.72 0.71
PhoBERT_large w2v2-Viet SLUE 0.64 0.66 0.65
PhoBERT_large w2v2-Viet Mod. SLUE 0.70 0.71 0.70
XLM-R_base XLSR-53-Viet SLUE 0.63 0.52 0.57
XLM-R_base XLSR-53-Viet Mod. SLUE 0.71 0.57 0.63
XLM-R_base w2v2-Viet SLUE 0.64 0.52 0.57
XLM-R_base w2v2-Viet Mod. SLUE 0.71 0.57 0.64
XLM-R_large XLSR-53-Viet SLUE 0.70 0.68 0.69
XLM-R_large XLSR-53-Viet Mod. SLUE 0.76 0.73 0.74
XLM-R_large w2v2-Viet SLUE 0.69 0.66 0.67
XLM-R_large w2v2-Viet Mod. SLUE 0.75 0.71 0.73
Table 29: NER results of DRUGCHEMICAL entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
FOODDRINK ViDeBERTa_base XLSR-53-Viet SLUE 0.25 0.29 0.27
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.28 0.33 0.30
ViDeBERTa_base w2v2-Viet SLUE 0.22 0.27 0.24
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.25 0.31 0.28
ViT5_base XLSR-53-Viet SLUE 0.54 0.53 0.54
ViT5_base XLSR-53-Viet Mod. SLUE 0.58 0.56 0.57
ViT5_base w2v2-Viet SLUE 0.55 0.53 0.54
ViT5_base w2v2-Viet Mod. SLUE 0.59 0.58 0.59
mBART-50 XLSR-53-Viet SLUE 0.59 0.06 0.12
mBART-50 XLSR-53-Viet Mod. SLUE 0.61 0.07 0.12
mBART-50 w2v2-Viet SLUE 0.54 0.05 0.09
mBART-50 w2v2-Viet Mod. SLUE 0.63 0.06 0.11
BARTpho XLSR-53-Viet SLUE 0.65 0.49 0.56
BARTpho XLSR-53-Viet Mod. SLUE 0.68 0.51 0.58
BARTpho w2v2-Viet SLUE 0.68 0.52 0.59
BARTpho w2v2-Viet Mod. SLUE 0.71 0.54 0.61
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.68 0.68 0.68
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.69 0.68 0.68
PhoBERT_base-v2 w2v2-Viet SLUE 0.69 0.66 0.67
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.69 0.67 0.68
PhoBERT_base XLSR-53-Viet SLUE 0.63 0.65 0.64
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.64 0.66 0.65
PhoBERT_base w2v2-Viet SLUE 0.64 0.62 0.63
PhoBERT_base w2v2-Viet Mod. SLUE 0.66 0.64 0.65
PhoBERT_large XLSR-53-Viet SLUE 0.68 0.69 0.68
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.69 0.70 0.69
PhoBERT_large w2v2-Viet SLUE 0.67 0.67 0.67
PhoBERT_large w2v2-Viet Mod. SLUE 0.68 0.68 0.68
XLM-R_base XLSR-53-Viet SLUE 0.52 0.63 0.57
XLM-R_base XLSR-53-Viet Mod. SLUE 0.55 0.67 0.61
XLM-R_base w2v2-Viet SLUE 0.53 0.63 0.57
XLM-R_base w2v2-Viet Mod. SLUE 0.56 0.67 0.61
XLM-R_large XLSR-53-Viet SLUE 0.75 0.69 0.72
XLM-R_large XLSR-53-Viet Mod. SLUE 0.76 0.70 0.73
XLM-R_large w2v2-Viet SLUE 0.76 0.68 0.72
XLM-R_large w2v2-Viet Mod. SLUE 0.77 0.69 0.73
Table 30: NER results of FOODDRINK entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
GENDER ViDeBERTa_base XLSR-53-Viet SLUE 0.00 0.00 0.00
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.00 0.00 0.00
ViDeBERTa_base w2v2-Viet SLUE 1,00 0.00 0.00
ViDeBERTa_base w2v2-Viet Mod. SLUE 1,00 0.00 0.00
ViT5_base XLSR-53-Viet SLUE 0.74 0.51 0.60
ViT5_base XLSR-53-Viet Mod. SLUE 0.79 0.54 0.65
ViT5_base w2v2-Viet SLUE 0.74 0.51 0.60
ViT5_base w2v2-Viet Mod. SLUE 0.76 0.53 0.62
mBART-50 XLSR-53-Viet SLUE 0.39 0.03 0.06
mBART-50 XLSR-53-Viet Mod. SLUE 0.50 0.04 0.08
mBART-50 w2v2-Viet SLUE 0.40 0.05 0.09
mBART-50 w2v2-Viet Mod. SLUE 0.53 0.07 0.12
BARTpho XLSR-53-Viet SLUE 0.78 0.59 0.67
BARTpho XLSR-53-Viet Mod. SLUE 0.82 0.61 0.70
BARTpho w2v2-Viet SLUE 0.78 0.58 0.66
BARTpho w2v2-Viet Mod. SLUE 0.80 0.59 0.68
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.78 0.60 0.68
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.82 0.63 0.71
PhoBERT_base-v2 w2v2-Viet SLUE 0.78 0.60 0.68
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.80 0.62 0.70
PhoBERT_base XLSR-53-Viet SLUE 0.78 0.61 0.68
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.81 0.63 0.71
PhoBERT_base w2v2-Viet SLUE 0.77 0.61 0.68
PhoBERT_base w2v2-Viet Mod. SLUE 0.80 0.62 0.70
PhoBERT_large XLSR-53-Viet SLUE 0.75 0.60 0.67
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.78 0.63 0.70
PhoBERT_large w2v2-Viet SLUE 0.76 0.60 0.67
PhoBERT_large w2v2-Viet Mod. SLUE 0.78 0.62 0.69
XLM-R_base XLSR-53-Viet SLUE 0.72 0.42 0.53
XLM-R_base XLSR-53-Viet Mod. SLUE 0.78 0.45 0.57
XLM-R_base w2v2-Viet SLUE 0.70 0.44 0.54
XLM-R_base w2v2-Viet Mod. SLUE 0.74 0.47 0.57
XLM-R_large XLSR-53-Viet SLUE 0.79 0.63 0.70
XLM-R_large XLSR-53-Viet Mod. SLUE 0.82 0.66 0.73
XLM-R_large w2v2-Viet SLUE 0.79 0.64 0.71
XLM-R_large w2v2-Viet Mod. SLUE 0.81 0.65 0.72
Table 31: NER results of GENDER entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
LOCATION ViDeBERTa_base XLSR-53-Viet SLUE 0.63 0.28 0.39
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.67 0.29 0.41
ViDeBERTa_base w2v2-Viet SLUE 0.59 0.26 0.36
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.60 0.26 0.37
ViT5_base XLSR-53-Viet SLUE 0.62 0.57 0.59
ViT5_base XLSR-53-Viet Mod. SLUE 0.65 0.60 0.63
ViT5_base w2v2-Viet SLUE 0.61 0.53 0.57
ViT5_base w2v2-Viet Mod. SLUE 0.64 0.56 0.60
mBART-50 XLSR-53-Viet SLUE 0.18 0.01 0.02
mBART-50 XLSR-53-Viet Mod. SLUE 0.41 0.02 0.04
mBART-50 w2v2-Viet SLUE 0.26 0.01 0.03
mBART-50 w2v2-Viet Mod. SLUE 0.33 0.02 0.04
BARTpho XLSR-53-Viet SLUE 0.65 0.66 0.65
BARTpho XLSR-53-Viet Mod. SLUE 0.68 0.69 0.69
BARTpho w2v2-Viet SLUE 0.65 0.65 0.65
BARTpho w2v2-Viet Mod. SLUE 0.69 0.68 0.68
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.66 0.71 0.69
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.69 0.74 0.71
PhoBERT_base-v2 w2v2-Viet SLUE 0.67 0.69 0.68
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.70 0.72 0.71
PhoBERT_base XLSR-53-Viet SLUE 0.62 0.68 0.65
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.66 0.72 0.69
PhoBERT_base w2v2-Viet SLUE 0.64 0.67 0.65
PhoBERT_base w2v2-Viet Mod. SLUE 0.67 0.71 0.69
PhoBERT_large XLSR-53-Viet SLUE 0.67 0.68 0.67
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.69 0.70 0.70
PhoBERT_large w2v2-Viet SLUE 0.70 0.66 0.68
PhoBERT_large w2v2-Viet Mod. SLUE 0.72 0.68 0.70
XLM-R_base XLSR-53-Viet SLUE 0.64 0.61 0.62
XLM-R_base XLSR-53-Viet Mod. SLUE 0.67 0.64 0.65
XLM-R_base w2v2-Viet SLUE 0.65 0.59 0.62
XLM-R_base w2v2-Viet Mod. SLUE 0.68 0.62 0.65
XLM-R_large XLSR-53-Viet SLUE 0.71 0.69 0.70
XLM-R_large XLSR-53-Viet Mod. SLUE 0.74 0.72 0.73
XLM-R_large w2v2-Viet SLUE 0.72 0.67 0.69
XLM-R_large w2v2-Viet Mod. SLUE 0.74 0.70 0.72
Table 32: NER results of LOCATION entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
MEDDEVICETECHNIQUE ViDeBERTa_base XLSR-53-Viet SLUE 0.22 0.07 0.10
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.33 0.10 0.16
ViDeBERTa_base w2v2-Viet SLUE 0.20 0.06 0.10
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.33 0.11 0.16
ViT5_base XLSR-53-Viet SLUE 0.39 0.09 0.14
ViT5_base XLSR-53-Viet Mod. SLUE 0.48 0.11 0.18
ViT5_base w2v2-Viet SLUE 0.44 0.11 0.17
ViT5_base w2v2-Viet Mod. SLUE 0.53 0.13 0.21
mBART-50 XLSR-53-Viet SLUE 0.31 0.01 0.01
mBART-50 XLSR-53-Viet Mod. SLUE 0.38 0.01 0.01
mBART-50 w2v2-Viet SLUE 0.26 0.01 0.01
mBART-50 w2v2-Viet Mod. SLUE 0.32 0.01 0.01
BARTpho XLSR-53-Viet SLUE 0.42 0.09 0.15
BARTpho XLSR-53-Viet Mod. SLUE 0.51 0.11 0.18
BARTpho w2v2-Viet SLUE 0.38 0.09 0.14
BARTpho w2v2-Viet Mod. SLUE 0.49 0.11 0.18
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.46 0.20 0.28
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.56 0.24 0.33
PhoBERT_base-v2 w2v2-Viet SLUE 0.43 0.18 0.25
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.52 0.22 0.30
PhoBERT_base XLSR-53-Viet SLUE 0.48 0.22 0.30
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.56 0.26 0.35
PhoBERT_base w2v2-Viet SLUE 0.44 0.20 0.28
PhoBERT_base w2v2-Viet Mod. SLUE 0.52 0.24 0.33
PhoBERT_large XLSR-53-Viet SLUE 0.46 0.15 0.23
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.55 0.18 0.27
PhoBERT_large w2v2-Viet SLUE 0.45 0.15 0.22
PhoBERT_large w2v2-Viet Mod. SLUE 0.55 0.18 0.27
XLM-R_base XLSR-53-Viet SLUE 0.38 0.13 0.19
XLM-R_base XLSR-53-Viet Mod. SLUE 0.47 0.16 0.24
XLM-R_base w2v2-Viet SLUE 0.34 0.12 0.17
XLM-R_base w2v2-Viet Mod. SLUE 0.44 0.15 0.22
XLM-R_large XLSR-53-Viet SLUE 0.47 0.15 0.23
XLM-R_large XLSR-53-Viet Mod. SLUE 0.55 0.18 0.27
XLM-R_large w2v2-Viet SLUE 0.43 0.13 0.20
XLM-R_large w2v2-Viet Mod. SLUE 0.53 0.16 0.25
Table 33: NER results of MEDDEVICETECHNIQUE entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
OCCUPATION ViDeBERTa_base XLSR-53-Viet SLUE 0.76 0.75 0.76
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.78 0.76 0.77
ViDeBERTa_base w2v2-Viet SLUE 0.77 0.74 0.76
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.79 0.75 0.77
ViT5_base XLSR-53-Viet SLUE 0.78 0.68 0.73
ViT5_base XLSR-53-Viet Mod. SLUE 0.79 0.70 0.74
ViT5_base w2v2-Viet SLUE 0.79 0.69 0.74
ViT5_base w2v2-Viet Mod. SLUE 0.80 0.70 0.75
mBART-50 XLSR-53-Viet SLUE 0.52 0.02 0.04
mBART-50 XLSR-53-Viet Mod. SLUE 0.54 0.02 0.05
mBART-50 w2v2-Viet SLUE 0.57 0.03 0.05
mBART-50 w2v2-Viet Mod. SLUE 0.60 0.03 0.06
BARTpho XLSR-53-Viet SLUE 0.79 0.78 0.79
BARTpho XLSR-53-Viet Mod. SLUE 0.80 0.79 0.80
BARTpho w2v2-Viet SLUE 0.79 0.77 0.78
BARTpho w2v2-Viet Mod. SLUE 0.81 0.79 0.80
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.79 0.84 0.81
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.80 0.85 0.82
PhoBERT_base-v2 w2v2-Viet SLUE 0.81 0.84 0.82
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.82 0.84 0.83
PhoBERT_base XLSR-53-Viet SLUE 0.79 0.84 0.81
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.79 0.85 0.82
PhoBERT_base w2v2-Viet SLUE 0.80 0.84 0.82
PhoBERT_base w2v2-Viet Mod. SLUE 0.81 0.85 0.83
PhoBERT_large XLSR-53-Viet SLUE 0.80 0.84 0.82
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.81 0.85 0.83
PhoBERT_large w2v2-Viet SLUE 0.82 0.84 0.83
PhoBERT_large w2v2-Viet Mod. SLUE 0.82 0.85 0.83
XLM-R_base XLSR-53-Viet SLUE 0.74 0.83 0.79
XLM-R_base XLSR-53-Viet Mod. SLUE 0.75 0.85 0.80
XLM-R_base w2v2-Viet SLUE 0.76 0.83 0.79
XLM-R_base w2v2-Viet Mod. SLUE 0.77 0.83 0.80
XLM-R_large XLSR-53-Viet SLUE 0.81 0.84 0.82
XLM-R_large XLSR-53-Viet Mod. SLUE 0.81 0.85 0.83
XLM-R_large w2v2-Viet SLUE 0.82 0.84 0.83
XLM-R_large w2v2-Viet Mod. SLUE 0.82 0.84 0.83
Table 34: NER results of OCCUPATION entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
ORGAN ViDeBERTa_base XLSR-53-Viet SLUE 0.36 0.31 0.33
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.49 0.41 0.45
ViDeBERTa_base w2v2-Viet SLUE 0.37 0.31 0.34
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.50 0.42 0.46
ViT5_base XLSR-53-Viet SLUE 0.50 0.39 0.44
ViT5_base XLSR-53-Viet Mod. SLUE 0.59 0.46 0.52
ViT5_base w2v2-Viet SLUE 0.53 0.40 0.46
ViT5_base w2v2-Viet Mod. SLUE 0.64 0.48 0.55
mBART-50 XLSR-53-Viet SLUE 0.29 0.06 0.09
mBART-50 XLSR-53-Viet Mod. SLUE 0.46 0.09 0.15
mBART-50 w2v2-Viet SLUE 0.30 0.06 0.10
mBART-50 w2v2-Viet Mod. SLUE 0.42 0.08 0.14
BARTpho XLSR-53-Viet SLUE 0.52 0.42 0.46
BARTpho XLSR-53-Viet Mod. SLUE 0.62 0.50 0.55
BARTpho w2v2-Viet SLUE 0.53 0.42 0.47
BARTpho w2v2-Viet Mod. SLUE 0.63 0.50 0.56
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.53 0.48 0.50
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.62 0.56 0.59
PhoBERT_base-v2 w2v2-Viet SLUE 0.55 0.49 0.52
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.64 0.57 0.60
PhoBERT_base XLSR-53-Viet SLUE 0.53 0.47 0.50
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.63 0.56 0.59
PhoBERT_base w2v2-Viet SLUE 0.54 0.48 0.51
PhoBERT_base w2v2-Viet Mod. SLUE 0.63 0.56 0.59
PhoBERT_large XLSR-53-Viet SLUE 0.52 0.46 0.49
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.62 0.55 0.58
PhoBERT_large w2v2-Viet SLUE 0.53 0.48 0.50
PhoBERT_large w2v2-Viet Mod. SLUE 0.63 0.57 0.59
XLM-R_base XLSR-53-Viet SLUE 0.52 0.48 0.50
XLM-R_base XLSR-53-Viet Mod. SLUE 0.61 0.56 0.59
XLM-R_base w2v2-Viet SLUE 0.53 0.50 0.51
XLM-R_base w2v2-Viet Mod. SLUE 0.62 0.58 0.60
XLM-R_large XLSR-53-Viet SLUE 0.55 0.47 0.50
XLM-R_large XLSR-53-Viet Mod. SLUE 0.64 0.55 0.59
XLM-R_large w2v2-Viet SLUE 0.56 0.48 0.52
XLM-R_large w2v2-Viet Mod. SLUE 0.65 0.56 0.60
Table 35: NER results of ORGAN entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
ORGANIZATION ViDeBERTa_base XLSR-53-Viet SLUE 0.00 0.00 0.00
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.00 0.00 0.00
ViDeBERTa_base w2v2-Viet SLUE 0.00 0.00 0.00
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.00 0.00 0.00
ViT5_base XLSR-53-Viet SLUE 0.88 0.30 0.45
ViT5_base XLSR-53-Viet Mod. SLUE 0.88 0.30 0.45
ViT5_base w2v2-Viet SLUE 0.81 0.27 0.41
ViT5_base w2v2-Viet Mod. SLUE 0.81 0.27 0.41
mBART-50 XLSR-53-Viet SLUE 0.00 0.00 0.00
mBART-50 XLSR-53-Viet Mod. SLUE 0.00 0.00 0.00
mBART-50 w2v2-Viet SLUE 0.00 0.00 0.00
mBART-50 w2v2-Viet Mod. SLUE 0.00 0.00 0.00
BARTpho XLSR-53-Viet SLUE 0.78 0.34 0.47
BARTpho XLSR-53-Viet Mod. SLUE 0.82 0.36 0.50
BARTpho w2v2-Viet SLUE 0.76 0.36 0.48
BARTpho w2v2-Viet Mod. SLUE 0.82 0.39 0.53
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.75 0.56 0.64
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.78 0.59 0.67
PhoBERT_base-v2 w2v2-Viet SLUE 0.76 0.55 0.64
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.79 0.57 0.66
PhoBERT_base XLSR-53-Viet SLUE 0.65 0.49 0.56
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.66 0.49 0.56
PhoBERT_base w2v2-Viet SLUE 0.62 0.45 0.52
PhoBERT_base w2v2-Viet Mod. SLUE 0.63 0.46 0.53
PhoBERT_large XLSR-53-Viet SLUE 0.73 0.50 0.59
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.73 0.50 0.60
PhoBERT_large w2v2-Viet SLUE 0.69 0.49 0.57
PhoBERT_large w2v2-Viet Mod. SLUE 0.71 0.50 0.59
XLM-R_base XLSR-53-Viet SLUE 0.62 0.26 0.37
XLM-R_base XLSR-53-Viet Mod. SLUE 0.64 0.27 0.38
XLM-R_base w2v2-Viet SLUE 0.57 0.26 0.35
XLM-R_base w2v2-Viet Mod. SLUE 0.59 0.27 0.37
XLM-R_large XLSR-53-Viet SLUE 0.66 0.56 0.61
XLM-R_large XLSR-53-Viet Mod. SLUE 0.68 0.58 0.63
XLM-R_large w2v2-Viet SLUE 0.65 0.55 0.60
XLM-R_large w2v2-Viet Mod. SLUE 0.69 0.58 0.63
Table 36: NER results of ORGANIZATION entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
PERSONALCARE ViDeBERTa_base XLSR-53-Viet SLUE 0.20 0.67 0.30
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.20 0.70 0.32
ViDeBERTa_base w2v2-Viet SLUE 0.20 0.67 0.31
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.20 0.69 0.31
ViT5_base XLSR-53-Viet SLUE 0.17 0.54 0.26
ViT5_base XLSR-53-Viet Mod. SLUE 0.17 0.57 0.27
ViT5_base w2v2-Viet SLUE 0.17 0.53 0.26
ViT5_base w2v2-Viet Mod. SLUE 0.18 0.57 0.27
mBART-50 XLSR-53-Viet SLUE 0.03 0.02 0.02
mBART-50 XLSR-53-Viet Mod. SLUE 0.06 0.05 0.06
mBART-50 w2v2-Viet SLUE 0.05 0.04 0.04
mBART-50 w2v2-Viet Mod. SLUE 0.05 0.05 0.05
BARTpho XLSR-53-Viet SLUE 0.19 0.63 0.29
BARTpho XLSR-53-Viet Mod. SLUE 0.19 0.65 0.30
BARTpho w2v2-Viet SLUE 0.18 0.63 0.28
BARTpho w2v2-Viet Mod. SLUE 0.19 0.64 0.29
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.18 0.71 0.29
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.19 0.73 0.30
PhoBERT_base-v2 w2v2-Viet SLUE 0.18 0.70 0.29
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.18 0.71 0.29
PhoBERT_base XLSR-53-Viet SLUE 0.19 0.69 0.30
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.20 0.71 0.31
PhoBERT_base w2v2-Viet SLUE 0.20 0.69 0.31
PhoBERT_base w2v2-Viet Mod. SLUE 0.20 0.70 0.32
PhoBERT_large XLSR-53-Viet SLUE 0.20 0.70 0.32
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.21 0.71 0.32
PhoBERT_large w2v2-Viet SLUE 0.20 0.68 0.31
PhoBERT_large w2v2-Viet Mod. SLUE 0.21 0.69 0.32
XLM-R_base XLSR-53-Viet SLUE 0.22 0.70 0.34
XLM-R_base XLSR-53-Viet Mod. SLUE 0.23 0.73 0.35
XLM-R_base w2v2-Viet SLUE 0.23 0.70 0.34
XLM-R_base w2v2-Viet Mod. SLUE 0.23 0.72 0.35
XLM-R_large XLSR-53-Viet SLUE 0.24 0.70 0.36
XLM-R_large XLSR-53-Viet Mod. SLUE 0.25 0.71 0.37
XLM-R_large w2v2-Viet SLUE 0.25 0.69 0.37
XLM-R_large w2v2-Viet Mod. SLUE 0.26 0.70 0.38
Table 37: NER results of PERSONALCARE entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
PREVENTIVEMED ViDeBERTa_base XLSR-53-Viet SLUE 0.00 0.00 0.00
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.00 0.06 0.01
ViDeBERTa_base w2v2-Viet SLUE 0.00 0.00 0.00
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.00 0.06 0.01
ViT5_base XLSR-53-Viet SLUE 0.01 0.11 0.01
ViT5_base XLSR-53-Viet Mod. SLUE 0.01 0.14 0.02
ViT5_base w2v2-Viet SLUE 0.01 0.19 0.03
ViT5_base w2v2-Viet Mod. SLUE 0.02 0.22 0.03
mBART-50 XLSR-53-Viet SLUE 0.00 0.00 0.00
mBART-50 XLSR-53-Viet Mod. SLUE 0.00 0.00 0.00
mBART-50 w2v2-Viet SLUE 0.00 0.00 0.00
mBART-50 w2v2-Viet Mod. SLUE 0.00 0.00 0.00
BARTpho XLSR-53-Viet SLUE 0.00 0.06 0.01
BARTpho XLSR-53-Viet Mod. SLUE 0.01 0.19 0.02
BARTpho w2v2-Viet SLUE 0.00 0.06 0.01
BARTpho w2v2-Viet Mod. SLUE 0.01 0.19 0.02
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.01 0.14 0.01
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.01 0.22 0.02
PhoBERT_base-v2 w2v2-Viet SLUE 0.01 0.19 0.02
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.01 0.28 0.03
PhoBERT_base XLSR-53-Viet SLUE 0.01 0.22 0.02
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.01 0.31 0.03
PhoBERT_base w2v2-Viet SLUE 0.01 0.28 0.03
PhoBERT_base w2v2-Viet Mod. SLUE 0.02 0.39 0.04
PhoBERT_large XLSR-53-Viet SLUE 0.01 0.22 0.02
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.02 0.31 0.03
PhoBERT_large w2v2-Viet SLUE 0.01 0.28 0.03
PhoBERT_large w2v2-Viet Mod. SLUE 0.02 0.36 0.04
XLM-R_base XLSR-53-Viet SLUE 0.00 0.00 0.00
XLM-R_base XLSR-53-Viet Mod. SLUE 0.00 0.06 0.01
XLM-R_base w2v2-Viet SLUE 0.00 0.00 0.00
XLM-R_base w2v2-Viet Mod. SLUE 0.00 0.06 0.01
XLM-R_large XLSR-53-Viet SLUE 0.01 0.11 0.01
XLM-R_large XLSR-53-Viet Mod. SLUE 0.01 0.19 0.02
XLM-R_large w2v2-Viet SLUE 0.01 0.14 0.01
XLM-R_large w2v2-Viet Mod. SLUE 0.01 0.22 0.02
Table 38: NER results of PREVENTIVEMED entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
SURGERY ViDeBERTa_base XLSR-53-Viet SLUE 0.56 0.24 0.33
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.61 0.26 0.36
ViDeBERTa_base w2v2-Viet SLUE 0.57 0.24 0.34
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.62 0.26 0.36
ViT5_base XLSR-53-Viet SLUE 0.46 0.29 0.36
ViT5_base XLSR-53-Viet Mod. SLUE 0.51 0.32 0.39
ViT5_base w2v2-Viet SLUE 0.51 0.28 0.36
ViT5_base w2v2-Viet Mod. SLUE 0.57 0.32 0.41
mBART-50 XLSR-53-Viet SLUE 0.22 0.05 0.08
mBART-50 XLSR-53-Viet Mod. SLUE 0.32 0.07 0.11
mBART-50 w2v2-Viet SLUE 0.22 0.06 0.09
mBART-50 w2v2-Viet Mod. SLUE 0.31 0.08 0.13
BARTpho XLSR-53-Viet SLUE 0.52 0.29 0.37
BARTpho XLSR-53-Viet Mod. SLUE 0.59 0.33 0.42
BARTpho w2v2-Viet SLUE 0.54 0.29 0.38
BARTpho w2v2-Viet Mod. SLUE 0.61 0.33 0.43
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.57 0.29 0.39
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.62 0.32 0.42
PhoBERT_base-v2 w2v2-Viet SLUE 0.60 0.32 0.42
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.65 0.35 0.46
PhoBERT_base XLSR-53-Viet SLUE 0.53 0.26 0.35
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.56 0.27 0.37
PhoBERT_base w2v2-Viet SLUE 0.56 0.27 0.37
PhoBERT_base w2v2-Viet Mod. SLUE 0.61 0.30 0.40
PhoBERT_large XLSR-53-Viet SLUE 0.57 0.28 0.38
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.61 0.31 0.41
PhoBERT_large w2v2-Viet SLUE 0.61 0.28 0.39
PhoBERT_large w2v2-Viet Mod. SLUE 0.66 0.31 0.42
XLM-R_base XLSR-53-Viet SLUE 0.50 0.32 0.39
XLM-R_base XLSR-53-Viet Mod. SLUE 0.54 0.35 0.42
XLM-R_base w2v2-Viet SLUE 0.54 0.33 0.41
XLM-R_base w2v2-Viet Mod. SLUE 0.58 0.35 0.44
XLM-R_large XLSR-53-Viet SLUE 0.58 0.28 0.38
XLM-R_large XLSR-53-Viet Mod. SLUE 0.62 0.30 0.41
XLM-R_large w2v2-Viet SLUE 0.62 0.30 0.40
XLM-R_large w2v2-Viet Mod. SLUE 0.65 0.32 0.43
Table 39: NER results of SURGERY entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
TRANSPORTATION ViDeBERTa_base XLSR-53-Viet SLUE 0.00 0.00 0.00
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.00 0.00 0.00
ViDeBERTa_base w2v2-Viet SLUE 0.00 0.00 0.00
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.00 0.00 0.00
ViT5_base XLSR-53-Viet SLUE 0.00 0.00 0.00
ViT5_base XLSR-53-Viet Mod. SLUE 0.00 0.00 0.00
ViT5_base w2v2-Viet SLUE 0.00 0.00 0.00
ViT5_base w2v2-Viet Mod. SLUE 0.00 0.00 0.00
mBART-50 XLSR-53-Viet SLUE 0.00 0.00 0.00
mBART-50 XLSR-53-Viet Mod. SLUE 0.00 0.00 0.00
mBART-50 w2v2-Viet SLUE 0.00 0.00 0.00
mBART-50 w2v2-Viet Mod. SLUE 0.00 0.00 0.00
BARTpho XLSR-53-Viet SLUE 0.17 0.50 0.25
BARTpho XLSR-53-Viet Mod. SLUE 0.17 0.50 0.25
BARTpho w2v2-Viet SLUE 0.00 0.00 0.00
BARTpho w2v2-Viet Mod. SLUE 0.00 0.00 0.00
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.50 1,00 0.67
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.50 1,00 0.67
PhoBERT_base-v2 w2v2-Viet SLUE 0.00 0.00 0.00
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.00 0.00 0.00
PhoBERT_base XLSR-53-Viet SLUE 0.40 1,00 0.57
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.40 1,00 0.57
PhoBERT_base w2v2-Viet SLUE 0.00 0.00 0.00
PhoBERT_base w2v2-Viet Mod. SLUE 0.00 0.00 0.00
PhoBERT_large XLSR-53-Viet SLUE 0.50 1,00 0.67
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.50 1,00 0.67
PhoBERT_large w2v2-Viet SLUE 0.00 0.00 0.00
PhoBERT_large w2v2-Viet Mod. SLUE 0.00 0.00 0.00
XLM-R_base XLSR-53-Viet SLUE 0.00 0.00 0.00
XLM-R_base XLSR-53-Viet Mod. SLUE 0.00 0.00 0.00
XLM-R_base w2v2-Viet SLUE 0.00 0.00 0.00
XLM-R_base w2v2-Viet Mod. SLUE 0.00 0.00 0.00
XLM-R_large XLSR-53-Viet SLUE 0.50 1,00 0.67
XLM-R_large XLSR-53-Viet Mod. SLUE 0.50 1,00 0.67
XLM-R_large w2v2-Viet SLUE 0.00 0.00 0.00
XLM-R_large w2v2-Viet Mod. SLUE 0.00 0.00 0.00
Table 40: NER results of TRANSPORTATION entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
TREATMENT ViDeBERTa_base XLSR-53-Viet SLUE 0.56 0.77 0.65
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.57 0.78 0.66
ViDeBERTa_base w2v2-Viet SLUE 0.58 0.75 0.65
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.58 0.76 0.66
ViT5_base XLSR-53-Viet SLUE 0.45 0.62 0.52
ViT5_base XLSR-53-Viet Mod. SLUE 0.46 0.63 0.53
ViT5_base w2v2-Viet SLUE 0.44 0.62 0.52
ViT5_base w2v2-Viet Mod. SLUE 0.45 0.64 0.53
mBART-50 XLSR-53-Viet SLUE 0.19 0.08 0.11
mBART-50 XLSR-53-Viet Mod. SLUE 0.23 0.09 0.13
mBART-50 w2v2-Viet SLUE 0.30 0.12 0.17
mBART-50 w2v2-Viet Mod. SLUE 0.34 0.13 0.19
BARTpho XLSR-53-Viet SLUE 0.56 0.71 0.62
BARTpho XLSR-53-Viet Mod. SLUE 0.57 0.72 0.64
BARTpho w2v2-Viet SLUE 0.58 0.71 0.64
BARTpho w2v2-Viet Mod. SLUE 0.60 0.73 0.66
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.54 0.77 0.64
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.55 0.78 0.65
PhoBERT_base-v2 w2v2-Viet SLUE 0.55 0.76 0.64
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.56 0.77 0.65
PhoBERT_base XLSR-53-Viet SLUE 0.56 0.77 0.65
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.58 0.79 0.67
PhoBERT_base w2v2-Viet SLUE 0.57 0.77 0.65
PhoBERT_base w2v2-Viet Mod. SLUE 0.58 0.78 0.66
PhoBERT_large XLSR-53-Viet SLUE 0.54 0.77 0.64
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.55 0.79 0.65
PhoBERT_large w2v2-Viet SLUE 0.54 0.76 0.63
PhoBERT_large w2v2-Viet Mod. SLUE 0.55 0.77 0.64
XLM-R_base XLSR-53-Viet SLUE 0.46 0.76 0.57
XLM-R_base XLSR-53-Viet Mod. SLUE 0.47 0.78 0.58
XLM-R_base w2v2-Viet SLUE 0.45 0.76 0.56
XLM-R_base w2v2-Viet Mod. SLUE 0.46 0.77 0.57
XLM-R_large XLSR-53-Viet SLUE 0.60 0.77 0.68
XLM-R_large XLSR-53-Viet Mod. SLUE 0.61 0.78 0.69
XLM-R_large w2v2-Viet SLUE 0.60 0.76 0.67
XLM-R_large w2v2-Viet Mod. SLUE 0.61 0.77 0.68
Table 41: NER results of TREATMENT entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
UNITCALIBRATOR ViDeBERTa_base XLSR-53-Viet SLUE 0.24 0.34 0.28
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.29 0.42 0.34
ViDeBERTa_base w2v2-Viet SLUE 0.25 0.34 0.29
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.29 0.40 0.34
ViT5_base XLSR-53-Viet SLUE 0.25 0.38 0.30
ViT5_base XLSR-53-Viet Mod. SLUE 0.28 0.42 0.33
ViT5_base w2v2-Viet SLUE 0.26 0.38 0.31
ViT5_base w2v2-Viet Mod. SLUE 0.28 0.42 0.34
mBART-50 XLSR-53-Viet SLUE 0.26 0.05 0.09
mBART-50 XLSR-53-Viet Mod. SLUE 0.41 0.08 0.14
mBART-50 w2v2-Viet SLUE 0.28 0.06 0.10
mBART-50 w2v2-Viet Mod. SLUE 0.36 0.08 0.13
BARTpho XLSR-53-Viet SLUE 0.34 0.41 0.37
BARTpho XLSR-53-Viet Mod. SLUE 0.37 0.44 0.40
BARTpho w2v2-Viet SLUE 0.35 0.42 0.39
BARTpho w2v2-Viet Mod. SLUE 0.38 0.45 0.41
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.35 0.55 0.43
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.40 0.63 0.49
PhoBERT_base-v2 w2v2-Viet SLUE 0.39 0.55 0.45
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.42 0.60 0.50
PhoBERT_base XLSR-53-Viet SLUE 0.34 0.52 0.41
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.39 0.59 0.47
PhoBERT_base w2v2-Viet SLUE 0.36 0.53 0.43
PhoBERT_base w2v2-Viet Mod. SLUE 0.40 0.59 0.47
PhoBERT_large XLSR-53-Viet SLUE 0.36 0.54 0.43
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.40 0.60 0.48
PhoBERT_large w2v2-Viet SLUE 0.38 0.55 0.45
PhoBERT_large w2v2-Viet Mod. SLUE 0.41 0.59 0.49
XLM-R_base XLSR-53-Viet SLUE 0.32 0.51 0.40
XLM-R_base XLSR-53-Viet Mod. SLUE 0.36 0.57 0.44
XLM-R_base w2v2-Viet SLUE 0.34 0.51 0.41
XLM-R_base w2v2-Viet Mod. SLUE 0.37 0.55 0.44
XLM-R_large XLSR-53-Viet SLUE 0.36 0.52 0.42
XLM-R_large XLSR-53-Viet Mod. SLUE 0.40 0.58 0.47
XLM-R_large w2v2-Viet SLUE 0.40 0.55 0.46
XLM-R_large w2v2-Viet Mod. SLUE 0.42 0.58 0.49
Table 42: NER results of UNITCALIBRATOR entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
Overall Macro ViDeBERTa_base XLSR-53-Viet SLUE 0.34 0.31 0.30
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.38 0.34 0.33
ViDeBERTa_base w2v2-Viet SLUE 0.39 0.30 0.29
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.43 0.33 0.33
ViT5_base XLSR-53-Viet SLUE 0.48 0.42 0.42
ViT5_base XLSR-53-Viet Mod. SLUE 0.52 0.45 0.46
ViT5_base w2v2-Viet SLUE 0.49 0.42 0.43
ViT5_base w2v2-Viet Mod. SLUE 0.53 0.46 0.46
mBART-50 XLSR-53-Viet SLUE 0.28 0.04 0.07
mBART-50 XLSR-53-Viet Mod. SLUE 0.36 0.06 0.09
mBART-50 w2v2-Viet SLUE 0.29 0.05 0.08
mBART-50 w2v2-Viet Mod. SLUE 0.38 0.06 0.10
BARTpho XLSR-53-Viet SLUE 0.52 0.49 0.48
BARTpho XLSR-53-Viet Mod. SLUE 0.57 0.53 0.51
BARTpho w2v2-Viet SLUE 0.51 0.46 0.47
BARTpho w2v2-Viet Mod. SLUE 0.56 0.50 0.50
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.56 0.59 0.55
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.59 0.63 0.58
PhoBERT_base-v2 w2v2-Viet SLUE 0.54 0.53 0.51
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.57 0.57 0.54
PhoBERT_base XLSR-53-Viet SLUE 0.54 0.58 0.53
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.57 0.62 0.56
PhoBERT_base w2v2-Viet SLUE 0.52 0.52 0.50
PhoBERT_base w2v2-Viet Mod. SLUE 0.55 0.56 0.53
PhoBERT_large XLSR-53-Viet SLUE 0.56 0.58 0.54
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.59 0.62 0.57
PhoBERT_large w2v2-Viet SLUE 0.53 0.53 0.50
PhoBERT_large w2v2-Viet Mod. SLUE 0.57 0.56 0.53
XLM-R_base XLSR-53-Viet SLUE 0.48 0.47 0.45
XLM-R_base XLSR-53-Viet Mod. SLUE 0.52 0.51 0.49
XLM-R_base w2v2-Viet SLUE 0.48 0.47 0.45
XLM-R_base w2v2-Viet Mod. SLUE 0.51 0.51 0.49
XLM-R_large XLSR-53-Viet SLUE 0.58 0.59 0.56
XLM-R_large XLSR-53-Viet Mod. SLUE 0.61 0.62 0.59
XLM-R_large w2v2-Viet SLUE 0.55 0.53 0.52
XLM-R_large w2v2-Viet Mod. SLUE 0.58 0.56 0.55
Table 43: NER results of Overall Macro entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.
Entity Type NER ASR Eval. Toolkit Prec. Rec. F1
Overall Micro ViDeBERTa_base XLSR-53-Viet SLUE 0.45 0.34 0.39
ViDeBERTa_base XLSR-53-Viet Mod. SLUE 0.51 0.39 0.45
ViDeBERTa_base w2v2-Viet SLUE 0.45 0.34 0.39
ViDeBERTa_base w2v2-Viet Mod. SLUE 0.51 0.39 0.44
ViT5_base XLSR-53-Viet SLUE 0.52 0.46 0.48
ViT5_base XLSR-53-Viet Mod. SLUE 0.58 0.50 0.54
ViT5_base w2v2-Viet SLUE 0.53 0.46 0.49
ViT5_base w2v2-Viet Mod. SLUE 0.59 0.51 0.54
mBART-50 XLSR-53-Viet SLUE 0.35 0.05 0.09
mBART-50 XLSR-53-Viet Mod. SLUE 0.46 0.07 0.12
mBART-50 w2v2-Viet SLUE 0.35 0.05 0.09
mBART-50 w2v2-Viet Mod. SLUE 0.47 0.07 0.12
BARTpho XLSR-53-Viet SLUE 0.56 0.50 0.53
BARTpho XLSR-53-Viet Mod. SLUE 0.61 0.55 0.58
BARTpho w2v2-Viet SLUE 0.55 0.50 0.52
BARTpho w2v2-Viet Mod. SLUE 0.61 0.55 0.58
PhoBERT_base-v2 XLSR-53-Viet SLUE 0.57 0.57 0.57
PhoBERT_base-v2 XLSR-53-Viet Mod. SLUE 0.62 0.61 0.62
PhoBERT_base-v2 w2v2-Viet SLUE 0.58 0.56 0.57
PhoBERT_base-v2 w2v2-Viet Mod. SLUE 0.62 0.61 0.62
PhoBERT_base XLSR-53-Viet SLUE 0.56 0.56 0.56
PhoBERT_base XLSR-53-Viet Mod. SLUE 0.61 0.61 0.61
PhoBERT_base w2v2-Viet SLUE 0.56 0.56 0.56
PhoBERT_base w2v2-Viet Mod. SLUE 0.61 0.60 0.61
PhoBERT_large XLSR-53-Viet SLUE 0.57 0.55 0.56
PhoBERT_large XLSR-53-Viet Mod. SLUE 0.62 0.60 0.61
PhoBERT_large w2v2-Viet SLUE 0.58 0.55 0.56
PhoBERT_large w2v2-Viet Mod. SLUE 0.62 0.59 0.61
XLM-R_base XLSR-53-Viet SLUE 0.54 0.52 0.53
XLM-R_base XLSR-53-Viet Mod. SLUE 0.59 0.57 0.58
XLM-R_base w2v2-Viet SLUE 0.54 0.52 0.53
XLM-R_base w2v2-Viet Mod. SLUE 0.59 0.57 0.58
XLM-R_large XLSR-53-Viet SLUE 0.60 0.56 0.58
XLM-R_large XLSR-53-Viet Mod. SLUE 0.64 0.60 0.62
XLM-R_large w2v2-Viet SLUE 0.60 0.56 0.58
XLM-R_large w2v2-Viet Mod. SLUE 0.64 0.60 0.62
Table 44: NER results of Overall Micro entity type (in percent) on ASR output of test set using various pre-trained language models and ASR models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: SLUE and our modified SLUE.